### v1

# Deep Learning-based Intrusion Detection Systems: A Survey

Abstract.
Intrusion Detection Systems (IDS) have long been a hot topic in the cybersecurity community. In recent years, with the introduction of deep learning (DL) techniques, IDS have made great progress due to their increasing generalizability. The rationale behind this is that by learning the underlying patterns of known system behaviors, IDS detection can be generalized to intrusions that exploit zero-day vulnerabilities. In this survey, we refer to this type of IDS as DL-based IDS (DL-IDS). From the perspective of DL, this survey systematically reviews all the stages of DL-IDS, including data collection, log storage, log parsing, graph summarization, attack detection, and attack investigation. To accommodate current researchers, a section describing the publicly available benchmark datasets is included. This survey further discusses current challenges and potential future research directions, aiming to help researchers understand the basic ideas and visions of DL-IDS research, as well as to motivate their research interests.
Intrusion detection systems, deep learning, survey
copyright: nonejournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 1journalnumber: 1article: 1publicationmonth: 4ccs: Security and privacy Intrusion detection systemsccs: Computing methodologies Machine learningccs: General and reference Surveys and overviews
1. Introduction
The promising Internet of Everything connects people, processes, data, and things through the Internet (Evans, 2012), bringing convenience and efficiency to the world. Yet its inevitable security vulnerabilities could be exploited by deliberate attackers. With increasingly sophisticated attack methods such as Advanced Persistent Threat (APT), the attackers are in a threatening position to sabotage network systems or steal sensitive data. The detection of intrusions, particularly based on DL, has consequently been a prominent topic in the cybersecurity community.
Generally, the intrusion detection process is automated through software or hardware systems, which are known as IDS (Rebecca and Mell, 2001; Liao et al., 2013; Nassar et al., 2019; Xue et al., 2021; Nasir et al., 2022; Ahmad et al., 2023; Awad et al., 2022; Goyal et al., 2023). Intrusion detection is the process of monitoring and analyzing the events occurring in a computer or a network for signs of intrusion. Intrusions occur when attackers access the systems via the Internet, as well as authorized users misuse their privileges or attempt to gain unauthorized privileges.
IDS play a critical role in protecting computer and network systems. The limitations of IDS may result in terrible damage to enterprises. One example is the recent Colonial Pipeline Ransomware Attack (Beerman et al., 2023). In April 2021, the hacking group DarkSide launched a ransomware attack on Colonial Pipeline, the biggest oil pipeline company in the United States, using an unused VPN account. Due to this attack, 5,500 miles of transportation pipelines were forced to shut down, affecting nearly 45% of the fuel supply on the Eastern Coast. The Colonial Pipeline paid \$4.4 million ransom money, in addition to the theft of over 100 GB of data. If the malware intrusion can be detected in time, the influence of this attack is allowed to be greatly mitigated or even eliminated.
1.1. Tough but Bright Intrusion Detection System
(a) Trend of CVE records and IDS papers.
(b) Category of CNNVD vulnerabilities.
Figure 1. Recent situation of IDS.
IDS have been increasingly challenged to effectively deal with intrusions for decades. It is noted in Figure 1(a) that the number of CVE [^1] Common Vulnerabilities and Exposures (CVE) is a security project for security information sharing and vulnerability management. CVE is a publicly accessible vulnerability recording database where each vulnerability has a common name and a unique identifier. records has presented an accelerating uptrend, especially in 2016, which suffered a sharp rise. After 2016, the number of CVE records stays growing at a high speed, reaching around 30,000 in 2024. Besides, according to the CNNVD [^2] Chinese National Vulnerability Database (CNNVD) is a Chinese national database that catalogs security vulnerabilities in software and hardware products. CNNVD also provides unique identifiers and descriptions similar to CVE. report shown in Figure 1(b), we can observe that almost all (i.e., 97.2%) vulnerabilities are medium risk or above, with high and critical risk accounting for 40% of them. The growing number of vulnerabilities and the large percentage of high-risk vulnerabilities both reveal the tough situation faced by IDS.
Nevertheless, an interesting observation from Figure 1(a) is that, against the number of CVE records, DL-IDS papers also started to emerge in 2016 and their amount grew year by year subsequently. We can notably find that the growth trend of DL-IDS papers is nearly the same as that of CVE records. The potential reason can be speculated as DL is an effective way for IDS to cope with their tough situation. Borrowing the strong generalizability from DL techniques, DL-IDS detection can be extended to zero-day intrusions that are almost impossible to detect with the traditional DL-IDS. Some studies (Yang et al., 2023b; Wang et al., 2022a; Zengy et al., 2022) demonstrate this speculation. In their experiments, DL-IDS are all reported with an achievement of over 90% detection accuracy while the traditional DL-IDS sometimes only have around 50% detection accuracy.
The IDS future is not only tough but also bright with the aid of DL - it is evident that the growth in the number of IDS papers primarily comes from those based on DL techniques. The proportion of DL-IDS papers rises from about 0% in 2016 to a very high 65.7% in 2024. This phenomenon reflects the great interests and visions of the cybersecurity community in DL-IDS. To date, the DL-IDS development has almost reached a decade, and thus, it is time, and also essential, to revisit how DL and IDS interact, identify emerging trends, and guide future research directions.
1.2. Related Surveys and Our Scope
Unfortunately, none of the related surveys in the last decade have systematically investigated DL-IDS. On one hand, some related surveys may only focus on a few parts of DL-IDS, such as log parsers (S and Nasreen, 2021; Lupton et al., 2021; Zhang et al., 2023), datasets (Stojanović et al., 2020) and attack modeling (Stojanović et al., 2020; Alshamrani et al., 2019). On the other hand, while several surveys (He et al., 2020; Inam et al., 2023; Lv et al., 2022; Li et al., 2021; Zipperle et al., 2022; Nassar et al., 2019; Martins et al., 2022; Bridges et al., 2019; Nasir et al., 2022; Khraisat et al., 2019; Liao et al., 2013) involve some DL-based approaches, they did not review DL-IDS from the perspective of DL particularly.
Partial Investigation for DL-IDS
The surveys (Zhang et al., 2023; S and Nasreen, 2021; Lupton et al., 2021; Stojanović et al., 2020; Alshamrani et al., 2019) are the typical example papers describing only a few parts of DL-IDS. Among them, Adel (Alshamrani et al., 2019) mainly studied various techniques and solutions that were tailored to APT attacks, as well as discussed where to make the APT detection framework smart. Scott et al. (Lupton et al., 2021) and Tejaswini et al. (S and Nasreen, 2021) detailedly discussed online log parsers and their applications for anomaly detection. Branka et al. (Stojanović et al., 2020) review APT datasets and their creation, along with feature engineering in attack modeling. Zhang et al. (Zhang et al., 2023) created an exhaustive taxonomy of system log parsers and empirically analyzed the critical performance and operational features of 17 open-source log parsers. For DL-IDS, all the above surveys are obviously insufficient to advance research understanding and provide theoretical suggestions.
Different Perspectives from DL-IDS
Another type of existing surveys engaged DL-IDS but studied them from the other perspectives (He et al., 2020; Inam et al., 2023; Lv et al., 2022; Li et al., 2021; Zipperle et al., 2022; Nassar et al., 2019; Martins et al., 2022; Bridges et al., 2019; Nasir et al., 2022; Khraisat et al., 2019; Liao et al., 2013; Ahmad et al., 2021). Specifically, the surveys (Khraisat et al., 2019; Liao et al., 2013) aim to give an elaborate image of IDS and comprehensively explain methods from signature checking to anomaly detection algorithms. Originating from log data, the survey (He et al., 2020) presented a detailed overview of automated log analysis for reliability engineering and introduced three tasks including anomaly detection, failure prediction, and failure diagnosis. In survey (Nasir et al., 2022), Nasir et al. explored the efficacy of swarm intelligence on IDS and highlighted the corresponding challenges in multi-objective IDS problems.
Additionally, data types inspire and contribute significantly to the related surveys, whose categories include host-based IDS (HIDS) (Martins et al., 2022; Bridges et al., 2019; Lv et al., 2022; Li et al., 2021; Zipperle et al., 2022) and network-based IDS (NIDS) (Nassar et al., 2019; Ahmad et al., 2021). Bridges et al. (Bridges et al., 2019) focused on IDS leveraging host data for the enterprise network. Martins et al. (Martins et al., 2022) brought the HIDS concept to the Internet of Things. As a representative form of data in HIDS, the provenance graph (Lv et al., 2022; Li et al., 2021; Zipperle et al., 2022) and its reduction techniques (Inam et al., 2023) were also extensively studied in survey literature. In NIDS, Nassar et al. (Nassar et al., 2019) studied the techniques of network intrusion detection, especially those with machine learning (ML). Ahmad et al. (Ahmad et al., 2021) further incorporated ML and DL into their NIDS survey and studied the downstream learning methods detailedly.
The above surveys, however, lack investigation and discussion about DL-IDS. DL techniques are only what they cover or involve, rather than the primary focus of their research.
Scope of Our Survey
Our work distinguishes from the related surveys by providing a comprehensive literature review of DL-IDS. From the perspective of DL, our survey elaborates a common workflow of DL-IDS and introduces the corresponding taxonomies of all modules within this workflow. Moreover, our survey discusses the possible challenges and research visions for DL-IDS, which include many DL-related issues that have not yet been studied by the existing surveys.
1.3. Contributions and Organization
In summary, this survey makes the following contributions:
  • Realizing that IDS has made significant progress with the aid of DL, we present a thorough survey for DL-IDS, formalizing its definition and clarifying its type location.
  • We outline the common workflow for DL-IDS, consisting of the data management stage and intrusion detection stage. Further, we systematically illustrate the research advances in all modules of this workflow and innovatively taxonomize the papers based on DL techniques.
  • From the perspective of DL, we discuss the potential challenges and future directions for DL-IDS, especially highlighting the ones unique to DL-IDS for accommodating current researchers.
Survey Structure
Section 2 introduces the survey methodology of this work. Section 3 describes the background knowledge about DL-IDS. Section 4 and Section 5 elaborates the recent research trends on data management and intrusion detection, respectively. Section 6 illustrates the benchmark datasets and their feature dimensions. Section 7 discusses the visions and challenges for future research. Lastly, the conclusion is presented in Section 8.
2. Survey Methodology
Figure 2. Source distribution of references.
To start our literature review, we selected several popular literature databases, including Web of Science (Analytics, 1997), IEEE Xplore (IEEE, 2000), and Scopus (Elsevier, 2021), as the search engine. For search keywords, we determined from generalized terms associated with DL-IDS, such as intrusion detection system, attack investigation, anomaly detection, threat detection, Advanced Persistent Threats, data provenance analysis, forensic analysis, causality analysis, log collection, log compression, log parsing, log storage, and log summarization. Then, we employed Connected Papers (Papers, 2020), a visual tool that assists researchers in finding relevant academic papers, to ensure that we did not overlook the typical related literature. Since the found literature is numerous and rather generalized for the DL-IDS scope, we carefully checked their topics and prioritized only academic papers that are significantly related to DL. Finally, all these papers were filtered based on the impact factors of their published journals or academic conferences, leaving us a total of 42 papers.
We identified a few venues that have published many significant papers in the field of DL-IDS. They can be broadly divided into three categories: network security, software and system, and database. We report the distribution of these papers with their published years in Figure 2.
3. Background
3.1. Intrusion Detection System
3.1.1. Definition of IDS
IDS have long been a central issue in the cybersecurity community, whose research can be traced back to the 1990s (Rebecca and Mell, 2001) or even earlier. According to the existing literature (Rebecca and Mell, 2001; Liao et al., 2013; Nassar et al., 2019; Xue et al., 2021; Nasir et al., 2022; Goyal et al., 2023), IDS can be defined progressively as follows:
Definition 3.1.
(Intrusion Detection System). Intrusion detection system is a software or hardware system to automate the process of intrusion detection.
Definition 3.2.
(Intrusion Detection). Intrusion detection is the process of monitoring and analyzing the events occurring in a computer or a network for signs of intrusions.
Definition 3.3.
(Intrusion). Intrusion is the attempt to undermine the confidentiality, integrity, and availability of a computer or a network, or to circumvent its security facilities.
3.1.2. Types of IDS
Figure 3. Types of IDS.
Generally, IDS can be further categorized into various types based on their data source. Well-known types include NIDS (Meng et al., 2020; Min et al., 2021; Trivedi et al., 2023; Qazi et al., 2023), HIDS (Xie et al., 2018; Wang et al., 2018; Han et al., 2020; Wang et al., 2020; Shin and Kim, 2020; Wei et al., 2021a; Fält et al., 2021; Anjum et al., 2022; Wu et al., 2022; Zengy et al., 2022; Wang et al., 2022a; Yang et al., 2023b; Jia et al., 2024; Rehman et al., 2024; Kapoor et al., 2021; Liao et al., 2023), and Provenance-based IDS (PIDS) (Xie et al., 2018; Han et al., 2020; Wang et al., 2020; Wei et al., 2021a; Anjum et al., 2022; Wu et al., 2022; Zengy et al., 2022; Wang et al., 2022a; Yang et al., 2023b; Jia et al., 2024; Rehman et al., 2024; Kapoor et al., 2021; Liao et al., 2023). In Figure 3, we present the taxonomy of these IDS types and their corresponding data sources, along with the location of DL-IDS within the taxonomy.
Definition 3.4.
(NIDS). NIDS are IDS whose data sources are network traffic between hosts.
NIDS takes network traffic between hosts as its input. It is usually deployed at the edge or key node of the network, allowing it to secure the whole computer system with limited data. Benefiting from the global perception of the whole computer system, NIDS does well in large-scale multi-host intrusions such as Distributed Denial-of-Service (DDoS) attacks. However, NIDS performs poorly in intra-host intrusions and is unable to analyze intrusions in the form of encrypted network traffic.
Definition 3.5.
(HIDS). HIDS are IDS whose data sources are system events within hosts.
HIDS, in contrast, uncovers intrusions through system events of individual hosts. Its data sources include file system changes, system calls, process activities, etc. HIDS can conduct comprehensive detection for a host, and is not affected by encrypted data since the decryption is also performed in the host. Nevertheless, the deployment and maintenance of HIDS is relatively difficult. HIDS should be adapted to hosts of different operating systems and runtime environments. This simultaneously introduces computation overhead for the hosts.
Definition 3.6.
(PIDS). PIDS are HIDS whose data sources are data provenance.
Definition 3.7.
(Data Provenance). Data provenance refers to the origin and the processes that an event has undergone from its creation to its current state.
PIDS is a subtype of HIDS, particularly referring to HIDS that utilizes data provenance as its data source. Due to analysis in the intact trail of events, PIDS is proven to be effective in coping with advanced attacks (Zipperle et al., 2022). By performing causality analysis on data provenance, PIDS can significantly reduce false alarms compared with traditional IDS. However, data provenance can be very expensive to obtain, requiring complicated technical tools for monitoring operating systems, network protocols, and applications.
Definition 3.8.
(DL-IDS.) DL-IDS are IDS that utilize DL techniques to detect intrusions, whose data sources can be network traffic between hosts, system events within hosts, or their combination.
Unlike the other types of IDS such as NIDS and HIDS are categorized by their data sources, DL-IDS is defined by the techniques used in intrusion detection. As shown in Figure 3, the data source of DL-IDS can be network traffic, system events, or both. Taking advantage of the generalizability of DL techniques, DL-IDS is allowed to handle zero-day attacks precisely and thus become extremely interested in the cybersecurity community recently. In this work, we comprehensively review DL-IDS, detailing its entire workflow and describing its visions and challenges.
Figure 4. Common workflow of DL-IDS.
3.2. Common Workflow
Figure 4 illustrates the common workflow of DL-IDS. It usually consists of 7 steps: raw data, collection, storage, parsing, summarization, detection, and investigation. The brief introductions of these steps are as follows:
  • Raw Data is unprocessed data for uncovering attack details or benign system behaviors. The raw data analyzed by cyber experts commonly include network traffic and audit logs.
  • Collection indicates data collection tools for different systems, such as cloud and cross-platforms, which gather valuable raw data to describe important system behavior scenarios.
  • Storage involves managing large amounts of collected log data, with storage engines and blockchain methods offering solutions for reliable and efficient performance.
  • Parsing is the act of analyzing the stored logs and other useful data. It extracts and organizes the underlying information within the data for subsequent processing.
  • Summarization refers to the operation of summarizing large volumes of parsed data based on its semantics. This reduces storage costs while preserving critical events.
  • Detection is the process of using detection tools such as models and algorithms to detect anomalies in analyzed data to determine whether the data contains intrusions.
  • Investigation is the further process of Detection. It reconstructs the entire attack scenarios from the detected malicious data by analyzing the causal relationship between them.
Note that DL-IDS can also be performed in other step orders by skipping some of the steps. For example, log data can be first parsed before storage (Liu et al., 2019). Attack investigation can be directly conducted without detection of intrusions (Alsaheel et al., 2021). Basically, the above steps of DL-IDS can be divided into two stages: data management stage and intrusion detection stage.
In the data preprocessing stage, security analysts collect crucial data to describe the states of computer systems via various data collection tools. This data is then efficiently and reliably stored and managed with the aid of ELK and blockchain methods. Lastly, log parsers are applied to organize the data into structured ones in which only key information are reserved.
In the intrusion detection stage, data is utilized at the semantic level. Based on its semantics, redundant and missing values of data are merged or removed to significantly reduce the computing costs. Subsequently, DL-based detection methods are developed to detect anomalies within data to identify the presence of intrusions. In the end, by conducting a causality analysis of the anomalies, intrusion scenarios can be recovered as detection reports that are provided to security analysts.
Our survey will be structured according to these two stages, illustrating the data management stage in Section 4 and the intrusion forensics stage in Section 5.
4. Data Management
This section elaborates on the data management stage of DL-IDS, including data collection (Section 4.1), log storage (Section 4.2), and log parsing (Section 4.3).
4.1. Data Collection
The first step of DL-IDS is to collect useful data from raw data. Raw data indicates records that document events, activities, and operations that occur within a system, application, or network (a.k.a., logs), represented by network traffic or audit logs. Generally, network traffic refers to data transmitted over a network, governed by standardized network protocols. And audit logs mainly refer to chronological record data generated by systems or applications. By collecting useful logs, DL-IDS is allowed to monitor the health condition and operational status of information systems (Zhang et al., 2023). Common attributes of logs include timestamp, event type, subject, object, description, etc.
However, on different platforms, logs possess different formats and organizational structures (Zhang et al., 2023; Zipperle et al., 2022; Bridges et al., 2019; Li et al., 2021). To counter this, researchers have created various log collection tools specialized for various systems. For example, in Windows systems, Event Viewer is employed to manage system logs. Yet in Linux systems, log files are usually saved in the /var/log/ directory. The classification of data collection tools is shown in Table 1, including Windows, Linux, Cloud, and Cross platforms.
Table 1. Log collection tools on different platforms.
Platform Type Tool                         Description
Windows platform ETW (Microsoft, [n. d.]) Providing developers comprehensive event tracing ability
Panorama (Yin et al., 2007) Hardware-level and OS-aware dynamic taint tracking
Linux platform auditd (Grubb, 2008) Native tools supported by the Linux kernel
sysdig (Kili, [n. d.]) Focusing on runtime monitoring and fault troubleshooting
CamFlow (Pasquier et al., 2017) Self-contained, easily maintainable implementation
Tracee (Tracee, 2022) Exposing system information as events based on eBPF
DataTracker (Stamatogiannakis et al., 2015) Monitoring unmodified binaries without their source codes
Inspector (Thalheim et al., 2016) Parallel provenance library that is POSIX-compliant
AutoLog (Huo et al., 2023) Analyzing programs so no need to run them
eAudit (Sekar et al., 2024) Fast, scalable and easily deployable data collection tools
Cloud platform K8S tools (Chen et al., 2023; Horalek et al., 2022) Adapting to cloud scenarios to meet enterprise needs
saBPF (Lim et al., 2021) An extension tool of eBPF for containers in cloud computing
Cross platform DTrace (Gregg and Mauro, 2011) Real-time tracing framework that supports many platforms
SPADE (Gehani and Tariq, 2012) Novel provenance kernel for cross-platform logging
4.1.1. Windows Platform Tools
Event Tracing for Windows (ETW) (Microsoft, [n. d.]) is a powerful event tracing mechanism provided by Microsoft. It consists of three components: providers, controllers, and consumers. ETW instrument applications to provide kernel event logging and allows developers to start and stop event tracing sessions momentarily.
Panorama (Yin et al., 2007) exploits hardware-level and OS-aware dynamic taint tracking to collect logs. Moreover, it develops a series of automated tests to detect malware based on several kinds of anomalous behaviors.
4.1.2. Linux Platform Tools
auditd (Grubb, 2008) is a native collection tools supported by the Linux kernel, which is responsible for writing audit logs to disk and monitoring a variety of auditable events such as system calls, file accesses, and modifications.
sysdig (Kili, [n. d.]) relies on the kernel module to achieve monitoring and data collection of the system. sysdig focuses on system runtime monitoring and fault troubleshooting, which is also widely used in containers and cloud-native environments.
CamFlow (Pasquier et al., 2017) designs a self-contained, easily maintainable implementation of whole-system provenance based on Linux Security Module, NetFilter, and other kernel facilities. Furthermore, it provides a mechanism to adapt the captured data provenance to applications and can be integrated across distributed systems.
Tracee (Tracee, 2022) takes advantage of the extended Berkeley Packet Filter (eBPF) framework to observe systems efficiently. It uses eBPF to tap into systems and expose that information as events.
DataTracker (Stamatogiannakis et al., 2015) is an open-source data provenance collection tool using dynamic instrumentation. It is able to identify data provenance relations of unmodified binaries without access to or knowledge of the source codes.
Inspector (Thalheim et al., 2016) is a Portable Operating System Interface (POSIX)-compliant data provenance library for shared-memory multi-threaded applications. It is implemented as a parallel provenance algorithm on a concurrent provenance graph.
AutoLog (Huo et al., 2023) generates runtime log sequences by analyzing source codes and does not need to execute any programs. It can efficiently produce log datasets (e.g., over 10,000 messages/min on Java projects) and has the flexibility to adapt to several scenarios.
eAudit (Sekar et al., 2024) is a scalable and easily deployable data collection tools. eAudit relies on the eBPF framework built into recent Linux versions, making it work out of the box on most of the Linux distributions.
4.1.3. Cloud Platform Tools
Cloud-native scenarios introduce different challenges compared with Windows or Linux platforms. For example, there are many different types of components such as containers, microservices, and Kubernetes (K8S) clusters, each of which generates its own logs with widely varying formats and contents. Additionally, components are basically characterized by dynamic expansion and contraction, making it hard to capture the whole data provenance.
To address the above challenges, Chen et al. (Chen et al., 2023) design a cloud log collection architecture on the basis of K8S, which is a central platform based on cloud-native technology. Josef et al. (Horalek et al., 2022) propose a log collection and analysis tool operated as Software as a Service (SaaS) in the cloud environment in Kubernetes technology, aiming to provide comprehensive logs across all microservices.
saBPF (Lim et al., 2021) is an extension tool of eBPF, aiming to deploy fully-configurable, high-fidelity, system-level audit mechanisms at the granularity of containers. saBPF is further developed with proof-of-concept intrusion detection system and access control mechanism to demonstrate its practicability.
Note that some collection tools in Windows and Linux platforms such as auditd (Grubb, 2008), sysdig (Kili, [n. d.]), and Tracee (Tracee, 2022) can also be applied in cloud computing environment. To distinguish, we only introduce collection tools tailored to the cloud platform here.
Cross-platform Tools
To effectively detect intrusions, an intuitive idea is to incorporate log data from various platforms to obtain a global view of the running system. However, it is impossible to merge these data directly due to the difference in their log formats.
DTrace (Gregg and Mauro, 2011) is a real-time dynamic tracing framework for troubleshooting kernel and application problems on production systems. It supports many platforms, including Linux, Windows, Solaris, macOS, FreeBSD, NetBSD, etc.
Support for Provenance Auditing in Distributed Environments (SPADE) (Gehani and Tariq, 2012) develops a novel provenance kernel that mediates between the producers and consumers of provenance information, and handles the persistent storage of records. Generally, it supports heterogeneous aggregating for system-level data provenance for data analysis across multiple platforms.
4.2. Log Storage
After collecting logs from raw data, the subsequent step is to store these logs (Zhang et al., 2023; Ding et al., 2023a). Data storage involves integration, management, maintenance, and retrieval of the data, mainly focusing on metrics such as reliability and efficiency. The storage cannot tamper with the completeness of data, enabling developers to precisely reconstruct history scenarios. Additionally, when utilizing the collected data, developers should pursue low latency and high throughput for efficient data queries.
Here, we will introduce two essential components for the data storage process: log storage systems and compression algorithms for these systems.
4.2.1. Log Storage Systems
The two most commonly used log storage systems are ELK (Ahmed et al., 2020) and Loki (Bautista et al., 2022). ELK is a powerful log management solution consisting of three open-source software components: Elasticsearch (Elastic, 2010), Logstash (Elastic, 2009), and Kibana (Elastic, 2013). Elasticsearch is the leading distributed, RESTful search and analytics data engine designed with speed and scalability. In Elasticsearch, everything is indexed with finite state transducers and thus it supports rapid searches across vast repositories. Meanwhile, Elasticsearch implements a powerful Domain Specific Language, making it convenient and flexible to retrieve data. Logstash is a server-side data preprocessing pipeline to collect and integrate data from multiple sources. It performs filtration and transformation on its data and then sends the data to a favorable “stash” (usually Elasticsearch for ELK). Kibana is a data analytics and visualization platform at both speed and scale. It provides straightforward user interfaces to create a variety of charts and dashboards to visualize the trends, distribution, or other information of the data stored in Elasticsearch. ELK is powerful enough to be applied in enterprise scenarios, however, its performance comes at a price. ELK sacrifices ease of configuration and installation, and may simultaneously introduce severe runtime overhead for its hosts.
In contrast, Loki is a lightweight logging system with low resource overhead developed by Grafana Labs. It is designed with simple operations and efficient storage. Instead of indexing everything of data like ELK does, Loki mainly creates indices grounded in log labels. Moreover, Loki is well suited for open-source monitoring and visualization tools such as Prometheus (Prometheus, 2014) and Grafana (Labs, 2014). Integrating these two tools enables Loki to construct a complete monitoring and log analysis platform for information systems.
With the development of cloud computing, log storage systems are usually deployed on the cloud and reinforced by blockchain methods. Cao et al. (Cao et al., 2021) propose a cloud-native log database named LogStore. It uses highly-scalable, cost-effective cloud object storage to overcome the bandwidth limitations that arise when writing large amounts of logs. In their work, an optimized column index structure called LogBlock has been designed to support full-text search queries. Huang (Huang, 2019) proposes a log storage framework that uses the Ethereum blockchain to store the hash of log files and employs smart contracts to create indices for log files. Xu et al. (Xu et al., 2023b) propose a log data storage scheme that combines blockchain networks and IPFS technology. IPFS is used to store a large amount of file log data and the blockchain system is combined to achieve decentralized and secure log storage. Pourmajidi et al. (Pourmajidi and Miranskyy, 2018) propose Logchain, a blockchain-based logging system. It encrypts and seals logs and adds them to a hierarchical ledger to prevent log tampering, providing an immutable platform for log storage.
4.2.2. Log Compression Algorithms
Logs are generated quickly and require significant memory usage. For example, it is measured that a browser can produce about 10 GB of log data each day (Ding et al., 2023a). Such oversize data thus should be compressed before storage. Typically, log compression algorithms can be categorized into two types: general-purpose algorithms and those specifically adapted to log data.
General Compression Algorithms
Table 2. Well-acknowledged general compression algorithms for log data.
Type Well-acknoledged compression algorithm
Dictionary-based LZ77 in gzip (Foundation, 1992), LZMA in 7zip_lzma (Pavlov, 2001), LZSS in quickLZ (QuickLZ, 2006)
Sorting-based BWT in bzip2 (Seward, 1996), ST in szip (Schindler, 1999)
Statistical-based PPMD in 7zip_ppmd, DMC in ocamyd (Schwellinger, 2008)
General compression algorithms refer to algorithms to reduce the size of data while preserving the information of the data. It is worth noting that general compression algorithms are also capable of compressing data other than logs, namely, log data is only one type of data they can compress. General compression algorithms can be classified into three categories based on their principles (Yao et al., 2020):
  • Dictionary-based Compression: It records repeated data as keys and replaces these data with their corresponding keys.
  • Sorting-based Compression: It sorts data to enable strategies that require ordering features.
  • Statistical-based Compression: It exploits statistical techniques to learn and predict the possible next token for existing tokens. The data is thus compressed as a statistical model.
Table 2 presents representative algorithms of the above three types. Due to the indeterminacy of statistical techniques, this type of compression algorithm may introduce losses in compression. Yet the other two types of algorithms are generally lossless. By evaluating the compression performance on 9 log files and 2 natural language files, a study shows that some general compression algorithms can achieve high compression ratios for log data, and log data is even easier to compress than natural language data (Yao et al., 2020).
Tailored Compression Algorithms
Different from natural language data, log data usually has specific structures and formal expressions. Therefore, compression algorithms should be adjusted for log data to improve the compression performance further. Yao et al. (Yao et al., 2021) propose LogBlock, which obtains small log blocks before compression and then uses a generic compressor to compress logs. Liu et al. (Liu et al., 2019) propose Logzip, which employs clustering algorithms to iteratively extract templates from raw logs and then obtain coherent intermediate representations for compressing logs. Facing the massive data streams generated by cloud services, Lin et al. (Lin et al., 2015) propose Cowic, which utilizes the characteristics of logs for column compression and can independently analyze one or several columns in log entries. Wei et al. (Wei et al., 2021b) propose a parser-based log compression tool, LogReducer, which shows a high compression ratio and fast compression speed in large-scale log scenarios such as the Alibaba Cloud environment. They further propose LogGrep (Wei et al., 2023) for fine-grained organization of log data using static and runtime patterns. Rodrigues et al. (Rodrigues et al., 2021) propose the lossless compression tool CLP, aiming to quickly retrieve log data while meeting compression requirements. CLP proposes to combine domain-specific compression and search with a generic lightweight compression algorithm. Li et al. (Li et al., 2024d) conduct empirical research on log data and propose LogShrink to overcome their observed limitations by leveraging the commonality and variability of log data. LogBlock (Yao et al., 2021) is designed to help existing jobs perform better. It reduces duplicate logs by preprocessing log headers and rearranging log contents, thereby improving the compression ratio of log files. LogReducer (Yu et al., 2023) is a framework that combines log hotspot identification and online dynamic log filtering. Its non-intrusive design significantly reduces log storage and runtime overhead. $\mu$Slope (Wang et al., 2024) is a compression and search method for semi-structured log data. It achieves efficient storage and query performance through data segmentation, pattern extraction, and index-free design. Denum (Yu et al., 2024) significantly improves log compression rates by optimizing the compression of digital tokens in logs. It is an efficient log compression tool suitable for scenarios where you need to save storage space or transmission bandwidth.
4.3. Log Parsing
To analyze log data, DL-IDS should first parse the data into structured and standardized formats for unified processing. This is because log data often originates from multiple different devices such as terminals, sensors, and network devices, which will easily lead to differences in their expressions.
The data parsing focuses on application logs and is usually executed by data classification and template extraction. Data classification is to classify log data into several groups through data structures such as vector representations of the data. Each group constitutes a template, which is used to extract features from log data and construct the structured logs. Basically, the existing log parsers mainly have three categories: clustering-based parsers, pattern-based parsers, and heuristic-based parsers. Figure 5 depicts the taxonomy of log parsers.
[…]
Figure 5. Taxonomy of data parsing.
4.3.1. Clustering-based Parsing.
Clustering-based parsers classify data using clustering algorithms. Xiao et al. (Xiao et al., 2023) propose LPV, which employs a hierarchical clustering algorithm to incrementally group logs based on Euclidean distance. Pokharel et al. (Pokharel et al., 2023) convert log messages into bi-grams and directly cluster them using the DBSCAN algorithm. Hamooni et al. (Hamooni et al., 2016) present a rapid log pattern recognition approach named LogMine. It is implemented in the map-reduce framework for distributed platforms to process millions of log messages in seconds. LogCluster (Lin et al., 2016) reduces the number of logs that need to be manually checked and improves the accuracy of problem identification through log clustering and the use of knowledge bases. LogOHC (Yang et al., 2019) achieves efficient log template extraction by combining text vectorization and online hierarchical clustering technology. It has high accuracy and real-time performance and is particularly suitable for processing multi-source heterogeneous log data. METING (Coustié et al., 2020) provides a robust and efficient log parsing method through frequent n-gram mining and flexible log grouping strategy, which can effectively process various types of log data. Kimura et al. (Kimura et al., 2019) propose an active fault detection system based on log generation patterns. Active fault detection of network logs was achieved by automatically extracting log templates, constructing log feature vectors, aggregating features, and using machine learning classifiers.
4.3.2. Frequency-based Parsing.
Frequency-based parsers discover patterns that exceed the frequency threshold and employ the mined patterns to parse logs. Sedki et al. (Sedki et al., 2022) propose the log parsing tool ULP, which combines string matching and local frequency analysis to efficiently parse large log files. This method first uses text processing methods to group log events, and then conducts frequency analysis on the log instances in each group to identify fixed and variable content in the logs. Dai et al. (Dai et al., 2020) propose Logram, which utilizes an n-gram dictionary for log parsing. For n-grams with a frequency below the threshold, Logram recursively converts to (n-1)-grams until a list of uncommon 2-grams is obtained. To mitigate the parameter sensitivity issue in log parsers, Dai et al. (Dai et al., 2023) further proposed an entropy-based log parser PILAR, which balances parsing accuracy and efficiency. Zhou et al. (Zhou and Su, 2023) propose the log parser Polo, which extracts log templates from logs using a prefix forest composed of a ternary search tree. Xu et al. (Xu et al., 2023a) propose a hybrid log parsing model called Hue, which performs parsing through user-adaptive methods. Prefix-Graph (Chu et al., 2021) is an efficient, adaptive, and universal log parsing method that can stably extract log templates without relying on domain knowledge and manual parameter tuning. FT-tree (Zhang et al., 2017) dynamically builds a template tree by identifying frequently appearing word combinations in system logs, supporting incremental learning. It is used to extract templates from switch system logs in data center networks.
4.3.3. Heuristic-based Parsing.
Heuristic-based parsers rely on empirical knowledge to classify log data. He et al. (He et al., 2017) propose the online log parsing method Drain, which employs a depth-fixed parsing tree to group the original logs and encodes them using specially designed parsing rules. Each new log is matched layer by layer and assigned to the group with the highest similarity at the leaf node. Le et al. (Le and Zhang, 2023) propose to use a hint-based few-sample learning algorithm, LogPPT, to capture log template patterns. Utilizing new prompt tuning methods and an adaptive random sampling algorithm, LogPPT performs well on multiple public datasets. Liu et al. (Liu et al., 2022b) propose the UniParser parser to address the issue of difficult processing of heterogeneous logs, using the Token Encoder and Context Encoder modules to learn log context features. Spell (Du and Li, 2016) is an efficient streaming log parsing method that can dynamically extract log patterns in online processing and significantly improve processing efficiency through pre-filtering steps. Logan (Agrawal et al., 2019) achieves efficient and scalable log parsing through distributed processing, LCS matching, dynamic matching tolerance, and periodic merging. USTEP (Vervaet et al., 2021) is an online log parsing method based on an evolutionary tree structure that can discover and encode new parsing rules. It achieves constant parsing time and can efficiently parse raw log messages in a streaming manner. MoLFI (Messaoudi et al., 2018) transforms the log message format recognition problem into a multi-objective optimization problem and uses the NSGA-II algorithm to automatically generate high-quality log message templates.
5. Intrusion Detection
The intrusion detection stage uncovers intrusions using the semantic-level information. This section classifies and summarizes the mainstream graph summarization (Section 5.1), attack detection (Section 5.2), and attack investigation (Section 5.3).
5.1. Graph Summarization
It is illustrated that stealthy malware will inevitably interact with the underlying OS and be captured by provenance monitoring systems (Wang et al., 2020), which is the reason why PIDS (a form of DL-IDS) work and flourish recently. Log data generated from provenance monitoring system is referred to as data provenance as mentioned. Offering advantages in high precision, however, DL-IDS using data provenance sacrifices memory performance. They are required to process a vast volume of log data, since data provenance records all trails of events from their creations to their current states at a very fine-grained level, even some of which is trivial.
Therefore, unlike network traffic and application logs used by the other forms of DL-IDS, data provenance is fine-grained, detailed, and rich in semantics. The aforementioned log storage systems (Section 4.2) and log compression algorithms (Section 4.2.2) are insufficient in improving the space efficiency of data provenance due to the absence of semantic information. They compress log data at the token level or the byte level, aiming to find duplicates that can be merged or re-represented.
To this end, graph summarization is widely investigated to incorporate semantic-level information to further reduce the size of log data. In graph summarization, data provenance is transformed into a provenance graph, of which the causal relations are utilized to build the semantic understanding of system activities. Referring to the definition of data provenance (Definition 3.7), the definition of provenance graph can be defined as follows:
Definition 5.1.
(Provenance Graph). Provenance graph is a representation of a collection of data provenance with causal relations. It is a directed acyclic graph $G=<V,E>$ where nodes $V$ are system entities and edges $E$ are system events.
Based on provenance graphs, graph summarization approaches are allowed to reduce the size of log data by directly removing irrelevant events, aggregating similar events, gathering similar execution entities, etc. Due to the exploitation of specific causal relations (or semantic information), graph summarization approaches are capable of performing erasion confidently and straightforwardly. This categorizes them as a type of lossy reduction, whereas the aforementioned log storage systems and log compression algorithms are usually lossless. We note that some surveys (e.g., (Zipperle et al., 2022; Inam et al., 2023)) may interchangeably use graph summarization and log compression to identify the approaches that reduce the size of log data. In this work, we explicitly distinguish them and refer to the lossless reduction as compression and the opposite one as summarization, respectively.
Table 3. Overview of graph summarization approaches.
Mode Approach Release Baseline Requirement
Offline ProvCompress (Xie et al., 2011) 2011 No Summarization None
BEEP (Lee et al., 2013a) 2013 No Summarization Instrumentation
LogGC (Lee et al., 2013b) 2013 BEEP + No Summarization Instrumentation
CPR + PCAR (Xu et al., 2016) 2016 No Summarization None
FD + SD (Hossain et al., 2018) 2018 CPR + PCAR None
LogApprox (Michael et al., 2020) 2020 GC + CPR + DPR None
Online ProTracer (Ma et al., 2016) 2016 BEEP + No Summarization Instrumentation
NodeMerge (Tang et al., 2018) 2018 No Summarization None
Winnower (Hassan et al., 2018) 2018 No Summarization None
KCAL (Ma et al., 2018) 2018 No Summarization Instrumentation
GS + SS (Zhu et al., 2021) 2021 FD + SD None
SEAL (Fei et al., 2021) 2021 FD None
FAuST (Inam et al., 2022) 2022 CPR + DPR None
AudiTrim (Sun et al., 2024) 2024 CPR + GS + F-DPR None
Table 3 presents representative data summarization approaches. In terms of their execution modes, we classify them into two categories: offline graph summarization and online graph summarization.
5.1.1. Offline Graph Summarization
Offline graph summarization requires historical log data to provide global knowledge, which extracts log data from persistent storage, summarizes the data, and pushes back the summarized data to the persistent storage.
In 2011, Xie et al. (Xie et al., 2011) take inspiration from web graphs to summarize provenance graphs. They argue that provenance graphs have similar organizational structure and characteristics to web graphs, such as locality, similarity, and consecutiveness. On this basis, existing web graph compression algorithms, enhanced with name-identified reference list and node-crossing gaps, are applied in provenance graphs. BEEP (Lee et al., 2013a) is developed based on the fact that a long-running execution can be partitioned into individual units. BEEP reverse engineers application binaries and instructions to perform selective logging for unit boundaries and unit dependencies. LogGC (Lee et al., 2013b) is an audit log system equipped with Garbage Collecting capabilities. During the system’s execution, LogGC can be invoked at any time to obtain the current audit logs and then generate audit logs that have been summarized. Xu et al. (Xu et al., 2016) initially put forward an aggregation algorithm PCR that preserves event dependencies during the data reduction process, ensuring high-quality forensic analysis. Subsequently, they propose a summarization algorithm PCAR and utilized domain knowledge to conduct further data summarization. Hossain et al. (Hossain et al., 2018) propose two dependency-preserving summarization approaches, FD and SD. FD is allowed to keep backward and forward forensic analysis results. SD preserves the results of common forensic analysis, which runs backward to find the entry points of intrusions and then runs forward from these points to unveil their impacts. LogApprox (Michael et al., 2020) targets the most space-intensive events found in logs, namely file I/O activity, which can account for up to 90% of the log content. Once a regular expression for a given process is learned, LogApprox matches and eliminates new events that match the regular expression.
5.1.2. Online Graph Summarization
Online graph summarization performs real-time summarization for continually coming provenance graphs, rather than dealing with a static graph.
ProTracer (Ma et al., 2016) alternates between system event logging and unit-level taint propagation. It has a lightweight kernel module and user space daemon for concurrent, out-of-order event processing. By combining logging and tainting, and building/evaluating a prototype, it boosts provenance tracing. NodeMerge (Tang et al., 2018) is a template-based graph summarization system for online event storage. It can directly work on the system-dependent provenance streams and compress data provenance via read-only file access patterns. Winnower (Hassan et al., 2018) is an extensible audit-based cluster monitoring system. For tasks replicated across nodes in distributed applications, it can define a model over audit logs to concisely summarize the behaviors of multiple nodes, thus eliminating the necessity of transmitting redundant audit records to the central monitoring node. Ma et al. (Ma et al., 2018) propose an online audit log compression system based on kernel cache to achieve high-performance audit logging. It features a multi-layer cache scheme, which is distributed among various kernel data structures and utilizes the cache to detect and compress redundant events. The approach proposed by Zhu et al. (Zhu et al., 2021) includes two real-time graph summarization strategies. The first strategy maintains global semantics, which identifies and removes redundant events that do not affect global dependencies. The second strategy is based on suspicious semantics. SEAL (Fei et al., 2021) is a novel graph summarization approach for causal analysis. Based on information-theoretic observations of system event data, it achieves lossless compression and supports real-time historical event retrieval. FAuST (Inam et al., 2022) is a logging daemon that performs transparent and modular graph summarization directly on system endpoints. FAuST consists of modular parsers that parse different audit log formats to create a unified in-memory provenance graph representation. AudiTrim (Sun et al., 2024) is an online data summarization approach based on event dependency analysis, which can significantly reduce the amount of log data while ensuring the accuracy of intrusion detection.
5.2. Attack Detection
Attack detection is located at the central position of DL-IDS. The objective of attack detection is to accurately identify malicious system events in log data while minimizing false alarms of normal system behaviors. Based on the types of log data used in DL-IDS, we categorize the attack detection approaches into audit log-based detectors, application log-based detectors, network traffic-based detectors, and cross-log-based detectors.
Table 4. Overview of attack detection approaches.
Data Type Approach Release Time Base Model Use of Graph Detection Style Detection Granularity
Audit Pagoda (Xie et al., 2018) 2018 - Online Edge
Wang et al. (Wang et al., 2018) 2018 LSTM - - Log Sequence
Unicorn (Han et al., 2020) 2020 - Online Node
Shin et al. (Shin and Kim, 2020) 2020 SVM, LR, KNN - - Sequence
DeepHunter (Wei et al., 2021a) 2021 GNN - Node
Fält et al. (Fält et al., 2021) 2021 Transformer - Offline Log Sequence
ANUBIS (Anjum et al., 2022) 2022 BNN - Event Traces
Paradise (Wu et al., 2022) 2022 - Online Node
ShadeWatcher (Zengy et al., 2022) 2022 GNN Offline Node
threaTrace (Wang et al., 2022a) 2022 GraphSAGE Online Node
ProGrapher (Yang et al., 2023b) 2023 - - Subgraph
MAGIC (Jia et al., 2024) 2024 GAT On/Offline Node, Subgraph
Flash (Rehman et al., 2024) 2024 GNN Online Node
Application DeepLog (Du et al., 2017) 2017 LSTM - Online Log Sequence
LogC (Yin et al., 2020) 2020 LSTM - Online Log Sequence
Yu et al. (Yu et al., 2021a) 2021 Bi-LSTM - - Log Sequence
LogST (Zhang et al., 2022) 2022 Sentence-BERT - - Log Sequence
LogBD (Liu et al., 2023) 2023 BERT - - Log Sequence
MDFULog (Li et al., 2023a) 2023 BERT - - Log Sequence
Traffic ACO-SVM (Meng et al., 2020) 2020 SVM - - Sequence
Min et al. (Min et al., 2021) 2021 MemAE - - Network Traffic
SigML++ (Trivedi et al., 2023) 2023 ANN - - Single Log
HDLNIDS (Qazi et al., 2023) 2023 CRNN - - Network Traffic
Audit and Application LogAnomaly (Meng et al., 2019) 2019 LSTM - Online Log Sequence
PROV-GEM (Kapoor et al., 2021) 2021 GCN - Node
LogBERT (Guo et al., 2021) 2021 BERT - - Log Sequence
NeuralLog (Le and Zhang, 2021) 2021 BERT - Online Log Sequence
GAN-EDC (Duan et al., 2021) 2021 GAN - - Single Log
LogLS (Chen et al., 2022) 2022 Dual LSTM - Offline Log Sequence
LogUAD (Wang et al., 2022b) 2022 Word2Vec - Offline Log Sequence
Adanomaly (Qi et al., 2022) 2022 BiGAN - - Single Log
LogBASA (Liao et al., 2023) 2023 BERT - - Log Sequence
LAnoBERT (Lee et al., 2023) 2023 BERT - Online Log Sequence
LogDAPT (Zheng et al., 2023) 2023 BERT - - Log Sequence
ASGNet (Yang et al., 2023a) 2023 RoBERTa - - Log Sequence
TCN-Log2Vec (Jiang et al., 2023) 2023 BERT - - Log Sequence
The overview of attack detection approaches is presented in Table 4 and its corresponding taxonomy is depicted in Figure 6. We note that recent years have also published many research papers for attack detection (Li et al., 2024a, 2023c; Xu et al., 2022; Hassan et al., 2020; Milajerdi et al., 2019; Yu et al., 2021b; Chen et al., 2024; El Khairi et al., 2024; Wang et al., 2023). Yet these papers do not belong to DL-IDS, which are thus excluded hereafter in our survey.
5.2.1. Audit Log-based Detectors
Audit logs are collected from hosts and provide atomic information to construct data provenance. In terms of learning techniques, detectors based on audit logs can be classified as traditional learning, graph neural network, and sequence neural network.
Traditional Learning
Traditional learning-based detectors refer to those that utilize fundamental learning techniques. Shin et al. (Shin and Kim, 2020) propose to employ diverse machine learning algorithms to categorize benign and malicious logs. Pagoda (Xie et al., 2018) takes into account the degree of abnormality of individual provenance paths and the entire provenance graph. It can quickly uncover intrusions if a serious compromise has been identified on paths. The detection rate of Pagoda is further improved by considering the behavior representation in the whole provenance graph. Unicorn (Han et al., 2020) is a real-time intrusion detector that efficiently constructs a streaming histogram to represent the history of system executions. The counting results of the histogram are updated immediately if new edges are generated. By exploring the neighborhood relationships of provenance graphs, Unicorn discovers the causal relationships of system entities in the context. Similarly, ANUBIS (Anjum et al., 2022) also employs provenance graphs to capture causal relationships to achieve high detection performance. The core predictive capability of ANUBIS is a Bayesian Neural Network, which can inform users about how confident it is in its predictions. Paradise (Wu et al., 2022) introduces a novel extraction strategy at the system log level to prune and extract process feature vectors, which are then stored in an efficient in-memory database. Paradise is independent of types of operating systems or provenance collection frameworks, and is capable of independently calculating provenance-based dependencies at the detection stage.
[…]
Figure 6. Taxonomy of attack detection approaches.
Graph Neural Network
Graph Neural Network (GNN) is demonstrated to do well in graph-structured DL-IDS (Wei et al., 2021a; Yang et al., 2023b; Jia et al., 2024; Rehman et al., 2024; Zengy et al., 2022). DeepHunter (Wei et al., 2021a) is a GNN-based graph matching approach that is designed with two novel networks: attribute embedding networks that could incorporate Indicators of Compromise (IoC) information and graph embedding networks that could capture the relationships between IOCs. threaTrace (Wang et al., 2022a) emerges as an online approach dedicated to detecting and tracing host-level threats. Its model is custom-tailored on the basis of GraphSAGE (Hamilton et al., 2017), an inductive graph neural network framework, aiming to learn the role of every benign entity within the data source graph. MAGIC (Jia et al., 2024) leverages Graph Attention Network (GAT) (Velickovic et al., 2017) as a crucial constituent of its graph representation module. It employs masked graph representation learning to mimic benign system events and executes multi-granularity detection under diverse supervision levels. MAGIC can be applied in both online detection and offline detection scenarios. MAGIC is able to adapt to concept drift and has minimal computational overhead, which makes it applicable to real-world online APT detection. ProGrapher (Yang et al., 2023b) is a graph embedding-based anomaly detection system that extracts features and detects anomalies by utilizing information in relational networks. ShadeWatcher (Zengy et al., 2022) is a recommendation-guided detector. It models high-order connectivity via item-side information, borrowing the recommendation concepts of user-item interactions and applying them to the security concepts of system entity interactions.
Sequence Neural Network
All logs can be formatted as sequential data. Hence, sequence neural networks such as Recurrent Neural Network (RNN) and Transformer are employed in some work (Wang et al., 2018; Fält et al., 2021). The approach of Wang et al. (Wang et al., 2018) employs Long Short-Term Memory (LSTM), a variant of RNN, for anomaly detection. In their work, LSTM is compared with other machine learning algorithms such as Gradient Boosting Decision Tree and Naive Bayes. The results indicate that LSTM can attain the optimal anomaly detection performance when utilizing two feature extraction methods. Fält et al. (Fält et al., 2021) present a log anomaly detection method that is entirely trained based on auxiliary system data and is unsupervised with respect to the target system. It utilizes transformers as its base model, whose performance results reveal great potential of the self-attention mechanism in the field of DL-IDS.
5.2.2. Application Log-based Detectors
Application logs are generated from the installed binaries. Generally, application logs are in the form of natural language text, as the generated templates of them are pre-defined by their developers. Hence, many Natural Language Processing (NLP) techniques are introduced in this type of DL-IDS.
Recurrent Neural Network
As one of the most classic methods, RNN is used commonly in DL-IDS. DeepLog (Du et al., 2017) employs LSTM to model system logs as natural language sequences. As a result, it is able to automatically learn log patterns from normal operations and detect anomalies when there is a deviation between log patterns and the trained model. LogC (Yin et al., 2020) comprises two stages: transforming log messages into log template sequences and component sequences. These two sequences are then input into a combined LSTM model for the detection of abnormal logs. LogC only requires normal log sequences to train the combined model. Both DeepLog and LogC are capable of conducting online detection. Yu et al. (Yu et al., 2021a) propose a novel framework for the automatic detection of log anomalies, which leverages a bidirectional LSTM (Bi-LSTM) model based on an attention mechanism. In contrast to existing log anomaly detection methods, this approach not only pays attention to the sequential and quantitative information of log sequences but also considers the hidden semantic information in logs, thus enhancing the efficiency of anomaly detection.
Transformer Architecture
Transformer (Vaswani et al., 2017) is an effective model for sequence data, which exploits a multi-head self-attention mechanism to realize a large model that can be parallelly computed. The most common Transformer-based model in DL-IDS is BERT (Devlin et al., 2019), which takes advantage of encoding sequence data and pre-training model parameters. LogBD (Liu et al., 2023) utilizes BERT to acquire the semantic information of logs and constructs distance-based anomaly detection via a domain adaptation approach. It employs a Temporal Convolutional Network (TCN) to extract the common characteristics of different system logs and map them into the same hypersphere space. MDFULog (Li et al., 2023a) introduces a BERT-based semantic feature extraction model to preserve the semantics of log messages and map them to log vectors. This effectively eliminates the work randomness and noise injection caused by log template updates and enhances the robustness of anomaly detection. LogST (Zhang et al., 2022) obtains the semantic representation of log events through the SBERT model, taking into account the semantics of each word and the word order relationship, thus facilitating a better understanding of the context of log sequences.
5.2.3. Network Traffic-based Detectors
Network traffic comes from communications between hosts across a computer network. It is ruled by network protocols such as Transmission Control Protocol (TCP) and the User Datagram Protocol (UDP).
Traditional Learning
ACO-SVM (Meng et al., 2020) is a network intrusion detection model that optimizes the parameters of SVM based on the ant colony algorithm. In this approach, the parameters of the SVM are regarded as the paths traversed by ants. Through collecting network status information, performing feature extraction and processing, and utilizing the pheromone update and node transfer of the ant colony algorithm to achieve path crawling, the optimal parameter combination of ACO-SVM is found. Min et al. (Min et al., 2021) propose an approach for network anomaly detection employing a memory-augmented deep autoencoder. This model learns the prototype patterns of normal inputs via the memory module and reconstructs the inputs of abnormal samples that are close to normal samples during training. Thus, even the latent vectors of abnormal inputs are transformed through the memory module and aggregated and reconstructed from similar normal samples learned from the memory that is learned only with normal samples, thereby resolving the issue of over-generalization. SigML++ (Trivedi et al., 2023) is a supervised anomaly detection approach featuring probabilistic polynomial approximation. It employs Fully Homomorphic Encryption and Artificial Neural Networks to approximate the sigmoid activation function, enabling the execution of anomaly detection without decrypting logs.
Recurrent Neural Network
HDLNIDS (Qazi et al., 2023) employs a Convolutional Recurrent Neural Network (CRNN) for the detection and classification of malicious network traffic, with the objective of enhancing the detection efficiency and accuracy of existing intrusion detection systems. It conducts local feature extraction via CNN and then utilizes the RNN layer to capture the temporal dependency of features, thereby enabling a more comprehensive identification of network traffic patterns.
5.2.4. Cross Log-based Detectors
Audit logs contain detailed system-level underlying information, network traffic logs depict interactions between systems, and application logs possess high-level semantic information. Employing various kinds of logs in DL-IDS can enhance the interpretability of logs and obtain a more comprehensive understanding of intrusions.
Traditional Learning
LogUAD (Wang et al., 2022b) is a method for offline detection. It does not require a log parsing procedure. LogUAD directly takes the original log message as input to avert the noise brought in by parsing. It uses Word2Vec to generate word vectors and combines Term Frequency-Inverse Document Frequency (TF-IDF) to generate weighted log sequence feature vectors for handling the evolution of log statements. Ultimately, it makes use of a computationally efficient unsupervised clustering algorithm to detect anomalies.
Recurrent Neural Network
LogAnomaly (Meng et al., 2019) constitutes a system log anomaly detection framework. The central notion of this framework is to model the log stream as a natural language sequence. LogAnomaly employs a pattern of offline learning and online detection. In the offline learning segment, LogAnomaly utilizes the FT-Tree to extract templates from historical logs and matches the historical logs with these templates. Subsequently, each template is transformed into a template vector sequence, and sequential and quantitative features are extracted via the LSTM model. In the online detection portion, LogAnomaly initially determines whether real-time logs can be matched with existing templates. If affirmative, it is converted into a template vector; otherwise, based on the similarity among template vectors, it will be approximated to an existing template vector. LogLS (Chen et al., 2022) is a system log anomaly detection approach based on symmetrically structured double-long short-term memory. LogLS is an optimization of the DeepLog method referred to in Section 5.2.2. It is designed to tackle the gradient issue of LSTM in long sequence prediction and accomplishes the prediction of unobserved logs by providing a feedback mechanism.
Transformer Architecture
Both NeuralLog (Le and Zhang, 2021) and LAnoBERT (Lee et al., 2023) are free from the need for a log parser. They are able to learn directly from the raw log data, thereby avoiding the issue of potentially losing crucial information during the standardization process. They both make use of BERT to encode the semantic significance of log messages and are capable of conducting online detection. The work presented in the paper LogDAPT (Zheng et al., 2023) is predominantly founded on NeuralLog. It is noted that NeuralLog necessitates a substantial amount of labeled data to yield satisfactory outcomes. In contrast, LogDAPT tackles the issue of insufficient labeled data by applying the second stage of in-domain pre-training, namely DAPT, on the BERT base model. Not only LogBERT (Guo et al., 2021) but also LogBASA (Liao et al., 2023) employ the log parser Drain to parse log events. Subsequently, they make use of the pre-trained BERT model to extract the semantic information of the events. In the TCN-Log2Vec (Jiang et al., 2023) framework, the log parsing component is optimized on the basis of Drain. Meanwhile, the BERT model is employed to precisely extract semantics from log templates. Additionally, TF-IDF is utilized to capture the context of log templates, and window-based techniques are adopted to obtain log sequence information. The ASGNet (Yang et al., 2023a) model is composed of three components: log statistics information representation (V-Net), log deep semantic representation (S-Net), and an adaptive semantic threshold mechanism (G-Net). V-Net employs an unsupervised variational autoencoder (VAE) to learn the global representation of each statistical feature vector. S-Net extracts the latent semantic representation of text input via a pre-trained RoBERTa model. G-Net aligns the information from the two sources and then adjusts the information flow.
Generative Adversarial Network
GAN-EDC (Duan et al., 2021) and Adanomaly (Qi et al., 2022) utilize technologies related to Generative Adversarial Network (GAN). GAN-EDC employs an LSTM-based encoder-decoder framework as the generator. Log keywords are taken as the input of the encoder, and the decoder outputs the generated log template. The discriminator utilizes CNN to identify the disparities between the generated log template and the real log template. Adanomaly is a log anomaly detection method based on Bidirectional GAN (BiGAN) and the stacking approach. It uses a CNN-based BiGAN model to extract features. Subsequently, it trains multiple classifiers by partitioning the extracted features into multiple subsets and integrates the classifiers using the stacking method.
Graph Neural Network
As data provenance graphs are complex datasets, typically represented as heterogeneous graphs and multiplex networks, PROV-GEM (Kapoor et al., 2021) employs a GCN-based approach. It is specially designed for attribute heterogeneous multiplex networks and is able to handle large networks and the heterogeneity of data provenance graphs.
5.3. Attack Investigation
Table 5. Overview of attack investigation approaches.
Approach Release Time Audit Log Applica- tion Log Use of Graph Base Model Starting Node Investigation Granularity
ProvDetector (Wang et al., 2020) 2020 - Doc2Vec - Path
ATLAS (Alsaheel et al., 2021) 2021 - LSTM Sequence
Liu et al. (Liu et al., 2022a) 2022 - Struc2Vec Provenance Graph
LogTracer (Niu et al., 2022) 2022 - DeepLog Path
ConLBS (Li et al., 2023b) 2023 - Transformer Provenance Graph
AirTag (Ding et al., 2023b) 2023 - BERT - Sequence
Karios (Cheng et al., 2024) 2023 GNN - Summary Graph
CAPTAIN (Ramaki et al., 2023) 2023 - - Provenance Graph
Except for identifying individual compromised nodes, IDS are supposed to unveil the full story of intrusions (a.k.a., attack scenarios). This process is referred to as attack investigation, which can be done by directly detecting attack scenarios (Wang et al., 2020), or analyzing the causal relations between compromised nodes (Alsaheel et al., 2021; Ding et al., 2023b). The attack scenarios are defined with scenario graphs as follows:
Definition 5.2.
(Scenario Graph). Scenario graph is a subgraph of its given provenance graph, which is constructed by the nodes and edges causally dependent on nodes of interest.
Definition 5.3.
(Attack Scenario). Attack scenario is a form of scenario graph, where its nodes of interest are compromised nodes.
In the past, attack investigation is conducted by forward analysis and backward analysis. Forward analysis discovers the influence that nodes of interest will cause and backward analysis traces back how nodes of interest are generated. Benefiting from DL techniques, both forward and backward analysis can be achieved by learning patterns and predicting possible attack scenarios. Table 5 summarizes the overview of attack investigation approaches.
Traditional Learning
ProvDetector (Wang et al., 2020) utilizes doc2vec to learn the embedding representation of paths in the provenance graph. This method enables the mapping of different components in the path (e.g., processes, files, and network connections) to a multidimensional numerical vector space. CAPTAIN (Ramaki et al., 2023) is a community-based APT analyzer in IT networks to detect Intrusion Kill Chain (IKC). CAPTAIN consists of a total of 12 different activities in two stages to detect possible IKCs of the APT attacks. CAPTAIN uses graph-based algorithms to model relationships among logged events. It detects event communities and summarizes them with centrality measures. By considering the nature of included events, it identifies malicious communities related to IKC stages. This ultimately leads to the reconstruction of the IKC of APT attacks.
Recurrent Neural Network
ATLAS (Alsaheel et al., 2021) is a framework that constructs end-to-end attack stories from readily available audit logs. It employs a novel combination of causal analysis, natural language processing, and machine learning techniques to build a sequence-based model. In ATLAS, LSTM enables the model to automatically learn to distinguish patterns in attack and non-attack sequences. This model establishes key patterns of attack and non-attack behaviors from causal graphs. LogTracer (Niu et al., 2022) is an efficient log-based anomaly tracing method. In attack path extraction, it weights edges of the initial provenance graph with an outlier algorithm, filters edges by threshold, and restructures the scenario graph. It finds the longest path as an attack path by assigning anomaly degrees to edges. It also uses an incremental graph processing strategy and the DeepLog model to process coarse-grained log information to assist the tracing process.
Transformer Architecture
AirTag (Ding et al., 2023b) employs unsupervised learning to train DL models directly from log texts rather than relying on labeled causal graphs. It uses the BERT model for unsupervised learning and retrains the BERT model based on unlabeled log data. ConLBS (Li et al., 2023b) combines a contrastive learning framework and multilayer Transformer network for behavior sequence classification. It constructs behavior sequences from audit logs, uses a lemmatization strategy to map semantics to the attack pattern layer, and can perform unsupervised representation learning on unlabeled sequences.
Graph Neural Network
Liu et al. (Liu et al., 2022a) propose an automated attack detection and investigation method via learning the context semantics of the provenance graph. The provenance graph analyzes temporal and causal dependencies of system events. The pre-trained Struc2Vec model is used to learn representations of unique nodes in a general set for node embedding. Kairos (Cheng et al., 2024) is a practical intrusion detection and investigation tool based on whole-system provenance. Kairos utilizes GNN to analyze system execution history and detect and reconstruct complex APTs. It employs a graph neural network-based encoder-decoder architecture to learn the temporal evolution of provenance graph structure changes and quantify the abnormal degree of each system event.
[…]
Figure 7. Taxonomy of attack investigation.
6. Benchmark Datasets
DL-IDS relies on high-quality data to train an effective model. This section introduces the dimensions of datasets (Section 6.1) and some public datasets widely used in DL-IDS (Section 6.2).
6.1. Dimensions of Datasets
To illustrate the quality of datasets, it is general to use the following dimensions:
  • Benign Scenarios: Benign data should cover benign behaviors and system activities to the greatest extent, enabling DL-IDS to learn patterns of benign behaviors to differentiate malicious behaviors.
  • Malicious Scenarios: Malicious data ought to incorporate typical attack scenarios while taking into account the diversity of attacks, including short-term and long-term attacks, as well as simple attacks and multi-stage attacks.
  • Ground-truth Labels: Data should be labeled as benign or malicious. For multi-stage attacks, it is useful to indicate the attack type or the attack stage it belongs to.
  • Data Granularities: Datasets can be in the form of different granularities. The most accepted one is to provide raw log data. Due to copyright concerns, some replicates (Jia et al., 2024; Ding et al., 2023b) merely provide post-processed log data without their processing source codes.
6.2. Public Datasets
Publicly available datasets bring a lot of convenience to research on DL-IDS. However, some researchers use self-made datasets that are not publicly available, making it difficult for other researchers to reuse their datasets (El Khairi et al., 2024). To address this issue, we collect and organize some open-source datasets for further studies, which are listed in Table 6.
Table 6. Overview of public datasets. W, L, F, A, M, and S represent the operating system of Windows, Linux, FreeBSD, Android, Mac, and supercomputer, respectively.
Dataset Release Size Scenarios Label Format System
LANL Dataset (Kent, 2015) 2015 12 GB - Yes .txt W
StreamSpot (Manzoor et al., 2016) 2016 2 GB 1 Yes .tsv L
AWSCTD (Čeponis and Goranin, 2018) 2018 39 GB - No SQLite W
DARPA TC E3 (DARPA, 2022) 2018 366 GB (Griffith et al., 2020) 6 No CDM W, L, F, A
DARPA TC E5 (DARPA, 2022) 2019 2,699 GB (Griffith et al., 2020) 8 No CDM W, L, F, A
DARPA OpTC (DARPA, 2019) 2020 1,100 GB (Anjum et al., 2021) - No eCAR W
Unicorn SC (Han et al., 2020) 2020 147 GB 2 Yes CDM L
CERT Dataset (Glasser and Lindauer, 2013; Lindauer, 2020) 2020 87 GB - Yes .csv W
Loghub (Zhu et al., 2023) 2020 77 GB - - .txt W, L, M, S
ProvSec (Shrestha et al., 2023) 2023 - 11 Yes .json L
LANL Dataset (Kent, 2015) is collected within the internal computer network of Los Alamos National Laboratory’s corporate. The dataset consists of 58 consecutive days of de-identified data, covering about 165 million events from 12 thousand users. To obtain, its data sources include Windows-based authentication events, process start and stop events, DNS lookups, network flows, and a set of well-defined red teaming events.
StreamSpot dataset (Manzoor et al., 2016) is made up of 1 attack and 5 benign scenarios. The attack scenario exploits a Flash vulnerability and gains root access to the visiting host by visiting a malicious drive-by download URL. The benign scenarios are relevant to normal browsing activity, specifically watching YouTube, browsing news pages, checking Gmail, downloading files, and playing a video game. All the scenarios are simulated through 100 automated tasks with the Selenium RC (ThoughtWorks, 2004).
DARPA TC datasets (DARPA, 2022) are sourced from the DARPA Transparent Computing (TC) program, identified by the number of engagements from E1 to E5. Among them, DARPA TC E3 is the most widely used. The TC program aims to make current computing systems transparent by providing high-fidelity visibility during system operations across all layers of software abstraction. Unfortunately, DARPA TC datasets are released without labels, and DARPA makes no warranties as to the correctness, accuracy, or usefulness of the datasets.
DARPA Operationally Transparent Cyber (OpTC) (DARPA, 2019) is a technology transition pilot study funded under Boston Fusion Corporate. The OpTC system architecture is based on the one used in TC program evaluation. In OpTC, every Windows 10 endpoint is equipped with an endpoint sensor that monitors post events, packs them into JSON records, and sends them to Kafka. A translation server aggregates the data into eCAR format and pushes them back to Kafka. OpTC scales TC components from 2 to 1,000 hosts. The dataset consists of approximately 1 TB of compressed JSON data in a highly instrumented environment over two weeks.
Unicorn SC (Han et al., 2020) is a dataset specifically designed for APT detection, proposed by Han et al., authors of the Unicorn model. The dataset includes two supply chain scenarios, wget and shell shock, where each scenario lasts for 3 days to simulate the long-term feature of APT attacks, resulting in provenance data containing 125 benign behaviors and 25 malicious behaviors. The data is saved in the form of provenance graphs, describing the causal relationships during the system execution process.
CERT Dataset (Lindauer, 2020) is a collection of synthetic insider threat test datasets that provide both background and malicious actor synthetic data. It is developed by the CERT Division, in collaboration with ExactData, LLC, and under sponsorship from DARPA I2O. CERT dataset learned important lessons about the benefits and limitations of synthetic data in the cybersecurity domain and carefully discussed models of realism for synthetic data.
Loghub dataset (Zhu et al., 2023) is a large collection of system log datasets, providing 19 real-world log data from various software systems, including distributed systems, supercomputers, operating systems, mobile systems, server applications, and standalone software. The objective of Loghub is to fill the significant gap between intelligent automated log analysis techniques and successful deployments in the industry. For the usage scenarios of Loghub, about 35% are anomaly detection, 13% are log analysis, and 8% are security.
ProvSec dataset (Shrestha et al., 2023) is created for system provenance forensic analysis. To fulfill data provenance requirements, ProvSec includes the full details of system calls including system parameters. In ProvSec, 11 realistic attack scenarios with real software vulnerabilities and exploits are used and an algorithm to improve the data quality in the system provenance forensics analysis is presented.
7. Challenges and Future Directions
After the detailed introduction to the data management stage and the intrusion detection stage, as well as the widely-used benchmark datasets, this section further discusses challenges encountered in existing DL-IDS and summarizes the corresponding visions. These include fundamental resources (Section 7.1), pre-trained large models (Section 7.2), and comprehensive applications (Section 7.3).
7.1. Fundamental Resources
Effective DL-IDS heavily depends on core fundamental resources such as datasets and computing facilities to develop (Khraisat et al., 2019). Here, we will discuss their challenges one after the other.
7.1.1. Poor Data Quality
Existing datasets for DL-IDS may contain errors, inaccuracies, or missing values. This leads to unreliable descriptions of system behaviors that may mislead DL-IDS. For example, in some cases of the DARPA TC dataset, the PROCESS object and its source fail to properly resolve conflicts, resulting in possible incorrect entries in the actor_id field. And the acuity_level value of the FLOW object is 0, while the value range for this field in other objects is from 1 to 5. Another example could be the LogChunks (Brandt et al., 2020) dataset. In this dataset, the content describing the failure reasons is possibly incomplete. This is because a chunk in LogChunks only contains a continuous substring of the log text and a failure reason may be described across multiple sections of the log. Moreover, LogChunks neglects the classification of failure reasons like test, compilation, and code inspection errors, which hinders further research from analyzing failure reasons.
Meanwhile, high-quality ground-truth labels are hard to acquire, which is impeded by the contradiction between fine-grained manual labeling and automated label generation. On one hand, for unknown intrusions such as zero-day attacks, security analysts usually have no trouble recovering attack scenarios, while it is very labor-intensive to correspond each attack scenario to certain log entries. The DAPRA TC dataset (DARPA, 2022) is a typical example of this phenomenon. It only provides a ground truth report for attack scenarios, which does not correspond to any specific log entries. Although a few researchers (Wang et al., 2022a) provide the third-party ground-truth labels that are manually identified by themselves, we empirically find some ambiguities between their ground-truth labels and the official attack scenario report. These ambiguities have an obviously negative effect on DL-IDS, and to some extent, they may even cause the accumulation of errors. On the other hand, the development of automated labeling tools is in an awkward position. The automation requires prior knowledge of intrusions to design auto-labeling techniques such as instrumentation. In other words, the automation can only handle intrusions that it can understand and detect, making it relatively useless to develop such auto-labeling tools.
In addition, there are no unified evaluation metrics for DL-IDS on these datasets (Cheng et al., 2024). For example, precision, recall, F1 score are usually exploited in most of work (Wang et al., 2020; Alsaheel et al., 2021; Rehman et al., 2024; Jia et al., 2024). Yet some papers (Ding et al., 2023b) propose to use True Positive Rate (TPR) and False Positive Rate (FPR) as evaluation metrics. All the comparison experiments between DL-IDS are thus hard to guarantee that the obtained performance is significant and practical. The potential of datasets will be weakened under this non-unified performance evaluation.
7.1.2. Insufficient Amount of Data
Although log data is generated very quickly (e.g., eBay generates 1.2 PB log data per day by 2018 (Samuel, 2018)), DL-IDS is still facing challenges in insufficient amounts of data. The reasons can be three-fold:
First, log data has an extremely large number of trivial events, which are proven to be not effective and usually removed by graph summarization (Zengy et al., 2022; Yang et al., 2023b). For example, data provenance provides fine-grained information about memory-related events, such as data-to-memory mapping and protection of certain memory addresses. These memory-related events basically do not involve attacks, and unfortunately, are always orthogonal to the existing DL-IDS. However, to ensure the completeness of data provenance and to capture very infrequent but inevitable memory attacks, these memory-related events are still recorded in benchmark datasets. Therefore, the usable part of each dataset is rather smalle for DL-IDS, which can be reflected by the high summarization ratio achieved by graph summarization approaches (e.g., 70% (Xu et al., 2016)).
The second reason for an insufficient amount of data is the limited dataset representativeness. As observed in Table 6, most of the datasets have no more than 10 attack scenarios, not to mention that each of these attack scenarios has been carefully chosen by their authors. This limited number of attack scenarios suggests that existing datasets are almost impossible to represent the diversified attack methods, as the number of CVE records has already been over 280,000 (Corporation, 2025). Furthermore, the existing datasets such as DAPRA TC (DARPA, 2022) are collected in a specific experimental environment and may not cover other types of normal system behaviors. Unicorn SC (Han et al., 2020) is generated by an idealized simulation of supply chain scenarios, which means many real-world features are prone to be ignored in this dataset. Hence, training DL-IDS on these non-representative datasets could be a disaster for the computer systems that they protect.
Table 7. Summarization of computating resources.
Type Approach Device Quantity
GPU DeepHunter (Wei et al., 2021a) Tesla P4 2
SHADEWATCHER (Zengy et al., 2022) Tesla V100 1
PROGRAPHER (Yang et al., 2023b) GeForce GTX TITAN X 1
FLASH (Rehman et al., 2024) GeForce RTX 2080 1
LogBD (Liu et al., 2023) GeForce RTX 2080 1
HDLNIDS (Qazi et al., 2023) GeForce GTX 2080Ti 1
NeuralLog (Le and Zhang, 2021) Tesla K40c 1
LogUAD (Wang et al., 2022b) Tesla V100 1
TCN-Log2Vec (Jiang et al., 2023) GeForce GTX 1660s 1
LogTracer (Niu et al., 2022) GeForce GTX 3080Ti 1
AIRTAG (Ding et al., 2023b) GeForce RTX 6000 1
MAGIC (Jia et al., 2024) GPU Unknown
CPU Only UNICORN (Han et al., 2020) vCPU 8
THREATRACE (Wang et al., 2022a) vCPU 16
Pagoda (Xie et al., 2018) Physical CPU 2
PROVDETECTOR (Wang et al., 2020) Physical CPU 1
SigML++ (Trivedi et al., 2023) Physical CPU 1
LogAnomaly (Meng et al., 2019) Physical CPU 1
GAN-EDC (Duan et al., 2021) Physical CPU 1
LogBASA (Liao et al., 2023) Physical CPU 1
KAIROS (Cheng et al., 2024) Physical CPU 1
CAPTAIN (Ramaki et al., 2023) Physical CPU 1
Paradise (Wu et al., 2022) Physical CPU Unknown
Finally, the accessibility of datasets further exacerbates the insufficient data problem. Due to privacy and copyright issues, some datasets may be proprietary or difficult to obtain (Wang et al., 2020, 2023). For example, ProvDetector (Wang et al., 2020) conducted a three-month system evaluation in an enterprise environment with 306 hosts and collected benign provenance data of 23 target programs. Yet this dataset has not been made public, rendering it unavailable to improve other DL-IDS and almost all the assessment settings related to ProvDetector are susceptible to inequity.
7.1.3. Potential Heavy Computation Requirements
Similar to DL techniques, DL-IDS also requires a large amount of computing resources to improve their performance. According to (Roberts et al., 2024), the generalizability of neural models is proportional to the investment of computing resources. To illustrate, we report the computing resources of some well-known DL-IDS in Table 7. It shows that existing DL-IDS are all relatively lightweight, as manifested by only one NVIDIA Tesla V100 GPU or only one NVIDIA GeForce RTX 6000 GPU. However, the requirements for computing resources are related to the number of log data. When the challenge of insufficient data is mitigated and more data is available, more computing resources are inevitably required. Meanwhile, we will illustrate in Section 7.2 that there are plenty of powerful techniques that have not been introduced in DL-IDS, which similarly will bring in computation requirements.
Additionally, several DL-IDS (Han et al., 2020; Cheng et al., 2024; Wu et al., 2022; Trivedi et al., 2023; Duan et al., 2021; Meng et al., 2019) have been specifically engineered for real-time intrusion detection. For example, the experiments in Unicorn (Han et al., 2020) demonstrate that the processing time of Unicorn is linear to its workloads with only one core. Yet the computational units, either CPU or GPU, all have their performance upper bound. Unicorn should be implemented in parallel if its workload is too high.
7.1.4. Future Directions
To conclude, the challenges for DL-IDS in fundamental resources consist of data quality, data volume, and computational overhead. Apart from unintentional errors and non-technical issues in fundamental resources, the research questions that urgently need to be addressed include the contradiction between unaffordable manual labeling and non-generalizable auto-labeling techniques, non-unified benchmark datasets and performance metrics, as well as potential heavy computational overheads. Therefore, we summarize the future directions as follows:
Future Directions Developing efficient man-machine interactive log labeling mechanisms and organizing open-source data-sharing platforms accordingly to provide large amounts of high-quality datasets. Maintaining effective and comprehensive benchmark datasets, accompanied by a unified performance metric framework for a fair comparison. Investigating parallel or simplified strategies for DL-IDS, and studying their integration with log storage systems to achieve end-to-end acceleration.
7.2. Pre-training Theories and Techniques
In recent years, significant progress has been made by Large Language Models (LLMs) in the field of DL. Their capacity to understand and generate dialogue has been greatly enhanced as the model parameters of LLMs keep rising. T5 (Raffel et al., 2020), BERT (Devlin et al., 2019), GPT (Radford, 2018), GPT-4 (Achiam et al., 2023), LaMDA (Thoppilan et al., 2022), and LLaMA (Touvron et al., 2023) are notable examples.
With the development of pre-training techniques, LLMs have been adopted in many fields such as finance (Zhao et al., 2024), education (Neumann et al., 2024), medicine (Peng et al., 2023), and even other domains of cybersecurity (Cummins et al., 2025; Hu et al., 2024; Gu, 2023). Nevertheless, the adoption of LLMs in DL-IDS is stagnant, as shown in Figure 8. We can observe that LLMs developed at full speed beginning in 2019. Their prosperity, however, has not extended to DL-IDS. Until now, the only two DL-IDS that incorporate pre-training techniques, AirTag (Ding et al., 2023b) and MAGIC (Jia et al., 2024), still do not make full use of the potential of LLMs. AirTag pre-trains a BERT model on application logs and detects intrusions in terms of embeddings generated by BERT. MAGIC introduces GraphMAE (Hou et al., 2022), a model architecture derived from Graph Autoencoder (Kipf and Welling, 2016) in 2016 but integrated with the famous masked self-supervised learning method (He et al., 2022) in 2022, to conduct self-supervised learning on provenance graphs. MAGIC further designs an adapter to apply the pre-trained model in different detection scenarios. Nevertheless, both AirTag and MAGIC can be regarded as preliminary explorations of pre-training techniques. According to the scaling law (Kaplan et al., 2020), the performance of LLMs will steadily improve, as the parameters, data, and computation increase. And the reasoning ability of LLMs will suddenly emerge (Wei et al., 2022), allowing them to chat with humans smoothly. Such advantageous abilities obviously have not been incorporated into DL-IDS.
Nowadays, some researchers (Ali, 2024; G. Lira et al., 2024; Li et al., 2024b) have started to explore the applications of LLMs on DL-IDS. Yet the theories and techniques of such combination remain challenges. In the following, we will illustrate the identified issues and then point out the future directions.
Figure 8. Interactions between DL models and DL-IDS. While DL models proposed before 2019 have already leveraged in DL-IDS, the emerging LLMs since 2020 remains underdeveloped in this domain.
7.2.1. Trade-off between Reliability and Generalizability
The governing concern for the employment of LLMs in DL-IDS is reliability (or explainability). Although offering generalizability, LLMs have long been denounced to have issues with hallucinations (Martino et al., 2023; Yao et al., 2023), privacy (Yao et al., 2024) and overreliance (Kim et al., 2024). These unexplainable and uncontrollable features are an absolute disaster for DL-IDS. For example, when feeding log data to LLMs, they sometimes are prone to hallucinate and provide wrong detection results. Attacks thus successfully bypass the detection facilities and can exfiltrate sensitive data in the victim computer systems. Another example for this is that sensitive information may leak from LLMs. Hui et al. (Hui et al., 2024) present a prompt leakage attack for LLMs, which is demonstrated to be effective in both offline settings and real-world LLM applications.
7.2.2. Short of Statistical Log Modeling
LLMs are developed on the basis of statistical language modeling (Rosenfeld, 2000; Jozefowicz et al., 2016), which is not insufficiently studied for log data. The statistical modeling of natural language can be traced back to the early 1950s, when Shannon pioneered the technique of predicting the next element of natural language text (Shannon, 1948) and discussed the n-gram model for English (Shannon, 1951). After that, as machine learning came into view of the NLP research communities, language modeling flourished, and many models such as TreeBank (Marcus et al., 1993), word2vec (Mikolov et al., 2013a, b) and LSTM (Hochreiter and Schmidhuber, 1997) were proposed. Over decades, researchers in NLP have gained solid knowledge of language modeling, whose interests gradually shifted to efficiency. An epoch-making model, Transformer (Vaswani et al., 2017), was presented using the multi-head self-attention mechanism to fulfill parallel computing, which was widely exploited in popular pre-trained models such as BERT (Devlin et al., 2019) and GPT (Achiam et al., 2023) afterward. It is evident that the success of LLMs comes from the prolonged studies on statistical language modeling.
Table 8. Comparison of research advances in statistical modeling of various data. “NL”, “PL” and “FL” represent Natural Language, Programming Language, and Formal Language, respectively. Note that PL is a type of FL.
Data Form Content Generation Rules Statistical Modeling Studies Pre-training
Text NL Grammar, pragmatics, semantics, etc (Rosenfeld, 2000; Jozefowicz et al., 2016; Shannon, 1951; Marcus et al., 1993) well-done
Speech NL Text rules (see above) and phonetics (Nussbaum et al., 2025; Kersta et al., 1960) well-done
Source code PL Lexical and syntactic definitions (Hindle et al., 2016; Ray et al., 2016; Allamanis et al., 2018) well-done
Log NL + FL Log template defined by developers future work underdeveloped
Unfortunately, there are almost no research efforts on statistical modeling of log data, resulting in pre-training techniques of DL-IDS remaining underdeveloped. By contrast, as illustrated in Table 8, other data types whose statistical modeling studies started earlier have already done well in pre-training techniques. For example, Hindle et al. (Hindle et al., 2016) demonstrate that the source code is very repetitive and predictable, and, in fact, even more so than natural language. Driven by such statistical modeling conclusion, DL-based source code applications (Xu et al., 2023c, 2024; Li et al., 2024c, 2022; Svyatkovskiy et al., 2020; Feng et al., 2020; Gu et al., 2025) such as code generation and code clone detection make great progress, many of which are proven to be greatly strengthened by pre-training techniques. Similar cases can be found for speech data, whose applications are like text to speech (Ren et al., 2019; Guo et al., 2024; Park et al., 2025) and speech recognition (Baevski et al., 2021).
Log data is also created by humans, similar to text, speech, and source code. It is generated according to developer-defined log templates, with a form of natural language (e.g., application logs) and formal language (e.g., data provenance in CDM format). Given the fact that natural language (e.g., text and speech) and formal language (e.g., source code) both exhibit positive features in statistical modeling, log data urgently demands similar achievements to facilitate its pre-training research. Although several works (Michael et al., 2020; Inam et al., 2023) have discussed the features of log data, they are orthogonal to the explainable combination of DL and IDS. Compared with the other data types, challenges in statistical log modeling, for instance, may lie in that logs are extremely long and detailed for reliable purposes. It is very common that the length of one single log entry is the same as that of one paragraph in natural language texts. This happens to be the shortcomings of LLMs - not being trustworthy in generated contents and not being able to handle long text.
7.2.3. Future Directions
According to the scaling laws (Kaplan et al., 2020) and emergent abilities theory (Wei et al., 2022), as the model size continues to grow, the performance of DL-IDS will increase simultaneously. Thus, increasing the amount of model parameters will be an inevitable trend for DL-IDS. The underlying research questions include the strategies for incorporating existing LLMs in intrusion detection, since it is infeasible to directly leverage unreliable LLMs to detect intrusions, and the theories and techniques for modeling long and detailed log data. We summarize the future directions as follows:
Future Directions Investigating how and where to introduce LLMs into DL-IDS, with the objective of balancing the generalizability provided by LLMs and the reliability required by DL-IDS. Exploring fundamental statistical modeling theories for log data. On this basis, designing pre-training frameworks for log data and its downstream tasks such as steps within the workflow of DL-IDS (see Section 3.2).
7.3. Comprehensive Applications and Scenarios
DL-IDS possess abilities that the traditional IDS lack, or are difficult to realize, such as generalizability for zero-day attacks and modeling ability for complicated downstream tasks. We will elaborate the possible new-style applications and discuss the challenges in and introduced by them.
7.3.1. Limited Forward and Backward Tracing Scope
Forward tracing and backward tracing are employed in attack investigation, as illustrated in Section 5.3. Under traditional settings, the forward tracing analyzes the influence a symptom node would have on the victim computer system, and the backward tracing discovers the starting node where the vulnerabilities exist (Zipperle et al., 2022).
We argue that the existing tracing scope is too limited to handle increasingly complicated intrusions and DL-IDS can be defined more broadly. In addition to investigating scenario graphs of intrusions, DL-IDS are supposed to further investigate why these intrusions occur and how to hold back them. The broader definition introduces more downstream tasks that would be difficult to accomplish without the assistance of DL techniques. Based on Definition 3.3, we reformulate the definition of intrusion in a broad sense for DL-IDS as follows:
Definition 7.1.
(Generalized Intrusion). Generalized intrusion is the malicious attempts against a computer, a network, or the corresponding security facilities, whose attributes encompass not only itself but also its underlying root causes and the relevant control measures.
In this way, the detection of DL-IDS has been extended to the broadly defined intrusions, including their attributes such as root causes and control measures. When executing backward tracing analysis, DL-IDS are not only required to detect the starting symptom nodes, but also required to find the root causes of these nodes (or vulnerabilities in source codes). In the forward tracing analysis, except for detecting the symptom nodes affected by intrusions, DL-IDS should perform an in-depth analysis to discover the potential symptom nodes and provide control measures for the intrusions.
Thankfully, several pioneering works have studied similar problems (Chen et al., 2024; Manso et al., 2019). In AiVl (Chen et al., 2024), algorithms to bridge log entries and program models are developed using dynamic-static program analysis. Root causes of code vulnerabilities are thus allowed to derive from intrusions. Pedro et al. (Manso et al., 2019) investigate detection and mitigation for DDoS attacks, which can be regarded as control measures against specific intrusions. We note that these research attempts are all based on heuristics and can only achieve limited efficacy, either in generating root causes using pre-defined rules, or in developing control measures for specific intrusions. The main obstacle is the lack of robust modeling and reasoning capabilities, which gives rise to a substantial need for the introduction of advanced techniques like DL.
7.3.2. Concerns about Data-driven Adversarial Attacks
To clarify the detection performance, DL-IDS commonly idealize the parts other than detection models in their threat model, of which the most commonly idealized part is data. Such idealization, however, leaves DL-IDS with weaknesses that could be exploited by invaders.
One usual assumption for data is related to data poisoning attacks, as (Rehman et al., 2024; Hassan et al., 2019; Jia et al., 2024; Han et al., 2020) did. In this assumption, no attacks are considered to compromise the security of the log collection systems, that is, log data utilized in DL-IDS is absolutely harmless. But apparently, as attacks become more stealthy and complicated, it is impossible to satisfy such an assumption to some extent. Besides, the harmful data could emerge through non-attacks. Security analysts may fail to label log entries correctly, resulting in an effect similar to the data poison attacks. Our manual checking empirically identifies that the labeling in threaTrace (Wang et al., 2022a) is not fully consistent with the official ground truth documentation. In many cases, by emulating benign system behaviors, the attackers are able to plant backdoors in the training data for DL-IDS, putting the computer systems at invading risk.
The robustness of DL-IDS is also challenged by evasion attacks. To evade the detection, the malicious behaviors usually mimic the benign ones (a.k.a., mimicry attack), making them hard to be detected. By early 2002, David et al. (Wagner and Soto, 2002) had indicated the danger of mimicry attacks on HIDS. Recently, researchers have started to investigate mimicry attacks on DL-IDS (Mukherjee et al., 2023; Goyal et al., 2023) and their studies all present effective evasion of detection. While recognized to be powerful, DL-IDS are plagued by the issue that even a trivial perturbation in the data can result in a successful evasion attack (Chakraborty et al., 2018). Aware of this issue, R-caid (Goyal et al., 2024) proposes to embed root causes into the detection model for countering adversarial attacks. However, as noted in recent work (Mukherjee et al., 2023; Goyal et al., 2023, 2024), data-driven attacks still remain a major challenge for DL-IDS.
7.3.3. Underexplored Promising Scenarios
DL-IDS show excellent performance in the protection of computer and network systems recently. Unfortunately, there are still many promising scenarios for DL-IDS that have not been explored sufficiently yet.
Mobile edge computing (MEC) (Mao et al., 2017; Abbas et al., 2017) is a typical scenario. In the MEC environment, mobile computing, network control, and storage are pushed at the network edges so as to enable computation-intensive tasks at the resource-limited devices. At the network edges, devices such as Unmanned Aerial Vehicles (UAVs) and New Energy Vehicles (NEVs) usually lack computing power and security facilities, making it difficult to prevent them from intrusions (Shrestha et al., 2021). In the meantime, containerized deployment has become one of the dominant ways to deploy microservices. Detecting intrusions on containers is thus of great importance, for which ReplicaWatcher (El Khairi et al., 2024) is a representative work with a special design for microservices. Additionally, industrial networks are characterized by high fidelity, stability, and real-time responsiveness (Knapp, 2024), leading to challenges in adapting DL-IDS to their infrastructures.
7.3.4. Future Directions
Although there has been plenty of research on DL-IDS, many applications and scenarios remain underdeveloped. DL-IDS are sought to be more broadly defined and applied. Based on the above discussion, we briefly summarize the future directions as follows:
Future Directions Extending the scope of forward tracing and backward tracing to intrusions in a broad sense, so that generating root causes and control measures for the broadly defined intrusions. Understanding data-driven adversarial attacks such as data poisoning attacks and mimicry attacks for devising more robust DL-IDS. Applying DL-IDS widely in more underexplored promising scenarios, and if possible, implementing unified frameworks for them.
8. Conclusion
The DL techniques bring reform to IDS, whose generalizability enables them to detect intrusions that have never been encountered before. Recognizing that the IDS development over the past decade primarily comes from DL-IDS, this survey revisits the common workflow for DL-IDS, elaborates each module in the workflow, and taxonomizes the research papers innovatively based on their DL techniques. Publicly available datasets for stimulating future research are introduced subsequently. In addition, from the perspective of DL, this survey digs deep into the potential challenges, emerging trends, and future directions for DL-IDS. The discussions suggest to us that DL-IDS are, fascinatingly, in an underdeveloped state. We hope that this survey can somewhat inspire current researchers and facilitate future investigations on DL-IDS.
Acknowledgements.
This research is sponsored in part by the NSFC program (No. 6212780016 and No. 62021002).

### v2

@@
 The detection of intrusions, particularly based on DL, has consequently been a prominent topic in the cybersecurity community.
-Generally, the intrusion detection process is automated through software or hardware systems, which are known as IDS (Rebecca and Mell, 2001; Liao et al., 2013; Nassar et al., 2019; Xue et al., 2021; Nasir et al., 2022; Ahmad et al., 2023; Awad et al., 2022; Goyal et al., 2023).
-Intrusion detection is the process of monitoring and analyzing the events occurring in a computer or a network for signs of intrusion.
-Intrusions occur when attackers access the systems via the Internet, as well as authorized users misuse their privileges or attempt to gain unauthorized privileges.
-IDS play a critical role in protecting computer and network systems.
+The automated system for detecting intrusions is known as IDS.
 The limitations of IDS may result in terrible damage to enterprises.
@@
 The Colonial Pipeline paid \$4.4 million ransom money, in addition to the theft of over 100 GB of data.
-If the malware intrusion can be detected in time, the influence of this attack is allowed to be greatly mitigated or even eliminated.
+If the malware intrusion can be detected in time, the influence of this attack can be greatly mitigated or even eliminated.
 1.1. Tough but Bright Intrusion Detection System
@@
 It is noted in Figure 1(a) that the number of CVE [^1] Common Vulnerabilities and Exposures (CVE) is a security project for security information sharing and vulnerability management.
-CVE is a publicly accessible vulnerability recording database where each vulnerability has a common name and a unique identifier.
+CVE is a publicly accessible database where each vulnerability has a common name and a unique identifier.
 records has presented an accelerating uptrend, especially in 2016, which suffered a sharp rise.
@@
 Borrowing the strong generalizability from DL techniques, DL-IDS detection can be extended to zero-day intrusions that are almost impossible to detect with the traditional DL-IDS.
-Some studies (Yang et al., 2023b; Wang et al., 2022a; Zengy et al., 2022) demonstrate this speculation.
+Some studies (Yang et al., 2023; Wang et al., 2022; Zengy et al., 2022) demonstrate this speculation.
 In their experiments, DL-IDS are all reported with an achievement of over 90% detection accuracy while the traditional DL-IDS sometimes only have around 50% detection accuracy.
@@
 Unfortunately, none of the related surveys in the last decade have systematically investigated DL-IDS.
-On one hand, some related surveys may only focus on a few parts of DL-IDS, such as log parsers (S and Nasreen, 2021; Lupton et al., 2021; Zhang et al., 2023), datasets (Stojanović et al., 2020) and attack modeling (Stojanović et al., 2020; Alshamrani et al., 2019).
+On one hand, some related surveys may only focus on a few parts of DL-IDS, such as log parsers (S and Nasreen, 2021; Lupton et al., 2021; Zhang et al., 2023b), datasets (Stojanović et al., 2020), attack modeling (Stojanović et al., 2020; Alshamrani et al., 2019), and specific DL technique type (Bilot et al., 2023).
 On the other hand, while several surveys (He et al., 2020; Inam et al., 2023; Lv et al., 2022; Li et al., 2021; Zipperle et al., 2022; Nassar et al., 2019; Martins et al., 2022; Bridges et al., 2019; Nasir et al., 2022; Khraisat et al., 2019; Liao et al., 2013) involve some DL-based approaches, they did not review DL-IDS from the perspective of DL particularly.
 Partial Investigation for DL-IDS
-The surveys (Zhang et al., 2023; S and Nasreen, 2021; Lupton et al., 2021; Stojanović et al., 2020; Alshamrani et al., 2019) are the typical example papers describing only a few parts of DL-IDS.
-Among them, Adel (Alshamrani et al., 2019) mainly studied various techniques and solutions that were tailored to APT attacks, as well as discussed where to make the APT detection framework smart.
+The surveys (Zhang et al., 2023b; S and Nasreen, 2021; Lupton et al., 2021; Stojanović et al., 2020; Alshamrani et al., 2019) are the typical example papers describing only a few parts of DL-IDS.
+Among them, Adel et al.
+(Alshamrani et al., 2019) mainly studied various techniques and solutions that were tailored to APT attacks, as well as discussed where to make the APT detection framework smart.
 Scott et al.
@@
 Zhang et al.
-(Zhang et al., 2023) created an exhaustive taxonomy of system log parsers and empirically analyzed the critical performance and operational features of 17 open-source log parsers.
+(Zhang et al., 2023b) created an exhaustive taxonomy of system log parsers and empirically analyzed the critical performance and operational features of 17 open-source log parsers.
+Tristan et al.
+(Bilot et al., 2023) focused on the applications of graph neural networks (GNNs) to IDS.
 For DL-IDS, all the above surveys are obviously insufficient to advance research understanding and provide theoretical suggestions.
 Different Perspectives from DL-IDS
-Another type of existing surveys engaged DL-IDS but studied them from the other perspectives (He et al., 2020; Inam et al., 2023; Lv et al., 2022; Li et al., 2021; Zipperle et al., 2022; Nassar et al., 2019; Martins et al., 2022; Bridges et al., 2019; Nasir et al., 2022; Khraisat et al., 2019; Liao et al., 2013; Ahmad et al., 2021).
+Another type of existing surveys involved DL-IDS but studied them from the other perspectives (He et al., 2020; Inam et al., 2023; Lv et al., 2022; Li et al., 2021; Zipperle et al., 2022; Nassar et al., 2019; Martins et al., 2022; Bridges et al., 2019; Nasir et al., 2022; Khraisat et al., 2019; Liao et al., 2013; Ahmad et al., 2021).
 Specifically, the surveys (Khraisat et al., 2019; Liao et al., 2013) aim to give an elaborate image of IDS and comprehensively explain methods from signature checking to anomaly detection algorithms.
@@
 Scope of Our Survey
-Our work distinguishes from the related surveys by providing a comprehensive literature review of DL-IDS.
-From the perspective of DL, our survey elaborates a common workflow of DL-IDS and introduces the corresponding taxonomies of all modules within this workflow.
+Our work distinguishes the related surveys by providing a comprehensive literature review of DL-IDS.
+From the perspective of DL, our survey elaborates on a common workflow of DL-IDS and introduces the corresponding taxonomies of all modules within this workflow.
 Moreover, our survey discusses the possible challenges and research visions for DL-IDS, which include many DL-related issues that have not yet been studied by the existing surveys.
@@
 In summary, this survey makes the following contributions:
-  • Realizing that IDS has made significant progress with the aid of DL, we present a thorough survey for DL-IDS, formalizing its definition and clarifying its type location.
+  • Realizing that IDS has made significant progress with the aid of DL over the last decade, we present a thorough survey for DL-IDS, formalizing its definition and clarifying its location among other types of IDS.
   • We outline the common workflow for DL-IDS, consisting of the data management stage and intrusion detection stage.
-Further, we systematically illustrate the research advances in all modules of this workflow and innovatively taxonomize the papers based on DL techniques.
+We further systematically illustrate the research advances in all modules of this workflow and innovatively taxonomize the papers based on DL techniques.
   • From the perspective of DL, we discuss the potential challenges and future directions for DL-IDS, especially highlighting the ones unique to DL-IDS for accommodating current researchers.
@@
 Section 3 describes the background knowledge about DL-IDS.
-Section 4 and Section 5 elaborates the recent research trends on data management and intrusion detection, respectively.
+Section 4 and Section 5 elaborates the recent research trends on data management stage and intrusion detection stage, respectively.
 Section 6 illustrates the benchmark datasets and their feature dimensions.
@@
 Figure 2. Source distribution of references.
+Figure 3. Types of IDS.
 To start our literature review, we selected several popular literature databases, including Web of Science (Analytics, 1997), IEEE Xplore (IEEE, 2000), and Scopus (Elsevier, 2021), as the search engine.
@@
 Then, we employed Connected Papers (Papers, 2020), a visual tool that assists researchers in finding relevant academic papers, to ensure that we did not overlook the typical related literature.
-Since the found literature is numerous and rather generalized for the DL-IDS scope, we carefully checked their topics and prioritized only academic papers that are significantly related to DL.
-Finally, all these papers were filtered based on the impact factors of their published journals or academic conferences, leaving us a total of 42 papers.
-We identified a few venues that have published many significant papers in the field of DL-IDS.
-They can be broadly divided into three categories: network security, software and system, and database.
-We report the distribution of these papers with their published years in Figure 2.
+Since the found literature is numerous and rather generalized for the DL-IDS scope, we carefully checked their topics and prioritized only academic papers that are related to DL.
+Finally, all these papers were filtered based on the impact factors of their published journals or academic conferences, leaving us a total of 113 papers.
+We identified a few venues that have published many significant papers in the field of DL-IDS, such as Security, S&P, CCS, NDSS, TIFS, TDSC, ICSE, ASE, ESEC/FSE, TSE, EuroSys, NSDI, OSDI, SOSP, ATC, ICDE, TKDE, SIGMOD, KDD, WWW, and IJCAI.
+We broadly divide them into four categories: security, software, system, and data.
+The distribution of these papers with their published years is reported in Figure 3.
 3. Background
@@
 3.1.2. Types of IDS
-Figure 3. Types of IDS.
-Generally, IDS can be further categorized into various types based on their data source.
-Well-known types include NIDS (Meng et al., 2020; Min et al., 2021; Trivedi et al., 2023; Qazi et al., 2023), HIDS (Xie et al., 2018; Wang et al., 2018; Han et al., 2020; Wang et al., 2020; Shin and Kim, 2020; Wei et al., 2021a; Fält et al., 2021; Anjum et al., 2022; Wu et al., 2022; Zengy et al., 2022; Wang et al., 2022a; Yang et al., 2023b; Jia et al., 2024; Rehman et al., 2024; Kapoor et al., 2021; Liao et al., 2023), and Provenance-based IDS (PIDS) (Xie et al., 2018; Han et al., 2020; Wang et al., 2020; Wei et al., 2021a; Anjum et al., 2022; Wu et al., 2022; Zengy et al., 2022; Wang et al., 2022a; Yang et al., 2023b; Jia et al., 2024; Rehman et al., 2024; Kapoor et al., 2021; Liao et al., 2023).
-In Figure 3, we present the taxonomy of these IDS types and their corresponding data sources, along with the location of DL-IDS within the taxonomy.
+Generally, IDS can be further categorized into various types based on their data sources (Zipperle et al., 2022).
+Well-known types include NIDS, HIDS, and Provenance-based IDS (PIDS).
+Figure 3 depicts IDS types, their data sources, and the location of DL-IDS within those IDS types.
 Definition 3.4.
@@
 Benefiting from the global perception of the whole computer system, NIDS does well in large-scale multi-host intrusions such as Distributed Denial-of-Service (DDoS) attacks.
-However, NIDS performs poorly in intra-host intrusions and is unable to analyze intrusions in the form of encrypted network traffic.
+However, NIDS performs poorly in intra-host intrusions and is difficult to analyze intrusions in the form of encrypted network traffic.
 Definition 3.5.
@@
 PIDS is a subtype of HIDS, particularly referring to HIDS that utilizes data provenance as its data source.
-Due to analysis in the intact trail of events, PIDS is proven to be effective in coping with advanced attacks (Zipperle et al., 2022).
-By performing causality analysis on data provenance, PIDS can significantly reduce false alarms compared with traditional IDS.
-However, data provenance can be very expensive to obtain, requiring complicated technical tools for monitoring operating systems, network protocols, and applications.
+Due to analysis in the intact trail of events, PIDS is proven effective in coping with advanced attacks (Zipperle et al., 2022).
+By performing causality analysis on data provenance, PIDS can significantly reduce false alarms.
+Yet, data provenance is very expensive to obtain, requiring complicated technical tools for monitoring operating systems, network protocols, and applications.
 Definition 3.8.
@@
 Taking advantage of the generalizability of DL techniques, DL-IDS is allowed to handle zero-day attacks precisely and thus become extremely interested in the cybersecurity community recently.
-In this work, we comprehensively review DL-IDS, detailing its entire workflow and describing its visions and challenges.
 Figure 4. Common workflow of DL-IDS.
 3.2. Common Workflow
-Figure 4 illustrates the common workflow of DL-IDS.
-It usually consists of 7 steps: raw data, collection, storage, parsing, summarization, detection, and investigation.
-The brief introductions of these steps are as follows:
+Figure 4 depicts the common workflow of DL-IDS. It usually consists of 7 steps: raw data, collection, storage, parsing, summarization, detection, and investigation, which are explained as follows:
   • Raw Data is unprocessed data for uncovering attack details or benign system behaviors. The raw data analyzed by cyber experts commonly include network traffic and audit logs.
   • Collection indicates data collection tools for different systems, such as cloud and cross-platforms, which gather valuable raw data to describe important system behavior scenarios.
-  • Storage involves managing large amounts of collected log data, with storage engines and blockchain methods offering solutions for reliable and efficient performance.
+  • Storage involves storage and search engines to manage large amounts of collected log data. Log data is labeled with indexes for efficient retrieval.
   • Parsing is the act of analyzing the stored logs and other useful data. It extracts and organizes the underlying information within the data for subsequent processing.
@@
 Attack investigation can be directly conducted without detection of intrusions (Alsaheel et al., 2021).
-Basically, the above steps of DL-IDS can be divided into two stages: data management stage and intrusion detection stage.
-In the data preprocessing stage, security analysts collect crucial data to describe the states of computer systems via various data collection tools.
-This data is then efficiently and reliably stored and managed with the aid of ELK and blockchain methods.
-Lastly, log parsers are applied to organize the data into structured ones in which only key information are reserved.
-In the intrusion detection stage, data is utilized at the semantic level.
-Based on its semantics, redundant and missing values of data are merged or removed to significantly reduce the computing costs.
-Subsequently, DL-based detection methods are developed to detect anomalies within data to identify the presence of intrusions.
-In the end, by conducting a causality analysis of the anomalies, intrusion scenarios can be recovered as detection reports that are provided to security analysts.
-Our survey will be structured according to these two stages, illustrating the data management stage in Section 4 and the intrusion forensics stage in Section 5.
+This survey is organized by the common workflow.
 4. Data Management
@@
 The first step of DL-IDS is to collect useful data from raw data.
-Raw data indicates records that document events, activities, and operations that occur within a system, application, or network (a.k.a., logs), represented by network traffic or audit logs.
-Generally, network traffic refers to data transmitted over a network, governed by standardized network protocols.
-And audit logs mainly refer to chronological record data generated by systems or applications.
-By collecting useful logs, DL-IDS is allowed to monitor the health condition and operational status of information systems (Zhang et al., 2023).
+Raw data indicates records that document events, activities, and operations that occur within a system, application, or network (a.k.a., logs), represented by audit logs or application logs within hosts, or network traffic between hosts.
+By collecting useful logs, DL-IDS is allowed to monitor the health condition and operational status of information systems (Zhang et al., 2023b).
 Common attributes of logs include timestamp, event type, subject, object, description, etc.
-However, on different platforms, logs possess different formats and organizational structures (Zhang et al., 2023; Zipperle et al., 2022; Bridges et al., 2019; Li et al., 2021).
+On different platforms, logs possess different formats and organizational structures (Zhang et al., 2023b; Zipperle et al., 2022; Bridges et al., 2019; Li et al., 2021).
 To counter this, researchers have created various log collection tools specialized for various systems.
@@
 saBPF (Lim et al., 2021) An extension tool of eBPF for containers in cloud computing
+ISDC (Mirnajafizadeh et al., 2024) Eliminating overheads on in-network resources
 Cross platform DTrace (Gregg and Mauro, 2011) Real-time tracing framework that supports many platforms
@@
 It consists of three components: providers, controllers, and consumers.
-ETW instrument applications to provide kernel event logging and allows developers to start and stop event tracing sessions momentarily.
+ETW instruments applications to provide kernel event logging and allows developers to start and stop event tracing sessions momentarily.
 Panorama (Yin et al., 2007) exploits hardware-level and OS-aware dynamic taint tracking to collect logs.
@@
 4.1.2. Linux Platform Tools
-auditd (Grubb, 2008) is a native collection tools supported by the Linux kernel, which is responsible for writing audit logs to disk and monitoring a variety of auditable events such as system calls, file accesses, and modifications.
+auditd (Grubb, 2008) is a native collection tool supported by the Linux kernel, which is responsible for writing audit logs to disk and monitoring a variety of auditable events such as system calls, file accesses, and modifications.
 sysdig (Kili, [n.
@@
 Furthermore, it provides a mechanism to adapt the captured data provenance to applications and can be integrated across distributed systems.
-Tracee (Tracee, 2022) takes advantage of the extended Berkeley Packet Filter (eBPF) framework to observe systems efficiently. It uses eBPF to tap into systems and expose that information as events.
+Tracee (Tracee, 2022) takes advantage of the extended Berkeley Packet Filter (eBPF) framework to observe systems efficiently.
+It uses eBPF to tap into systems and expose that information as events.
 DataTracker (Stamatogiannakis et al., 2015) is an open-source data provenance collection tool using dynamic instrumentation.
@@
 4.1.3. Cloud Platform Tools
-Cloud-native scenarios introduce different challenges compared with Windows or Linux platforms.
-For example, there are many different types of components such as containers, microservices, and Kubernetes (K8S) clusters, each of which generates its own logs with widely varying formats and contents.
-Additionally, components are basically characterized by dynamic expansion and contraction, making it hard to capture the whole data provenance.
-To address the above challenges, Chen et al.
+Although some collection tools in Windows and Linux platforms such as auditd (Grubb, 2008), sysdig (Kili, [n.
+d.]), and Tracee (Tracee, 2022) can be applied in cloud computing environment, cloud-native scenarios introduce different challenges compared with Windows or Linux platforms.
+First, there are many different types of components such as containers, microservices, and Kubernetes (K8S) clusters in cloud platforms, each of which generates its own logs with varying formats and contents.
+Additionally, components are basically characterized by dynamic expansion and contraction, making it hard to capture complete log data.
+To address them, Chen et al.
 (Chen et al., 2023) design a cloud log collection architecture on the basis of K8S, which is a central platform based on cloud-native technology.
 Josef et al.
-(Horalek et al., 2022) propose a log collection and analysis tool operated as Software as a Service (SaaS) in the cloud environment in Kubernetes technology, aiming to provide comprehensive logs across all microservices.
+(Horalek et al., 2022) propose a log collection and analysis tool operated as Software as a Service (SaaS) in the cloud environment in K8S technology, aiming to provide comprehensive logs across all microservices.
 saBPF (Lim et al., 2021) is an extension tool of eBPF, aiming to deploy fully-configurable, high-fidelity, system-level audit mechanisms at the granularity of containers.
-saBPF is further developed with proof-of-concept intrusion detection system and access control mechanism to demonstrate its practicability.
-Note that some collection tools in Windows and Linux platforms such as auditd (Grubb, 2008), sysdig (Kili, [n.
-d.]), and Tracee (Tracee, 2022) can also be applied in cloud computing environment.
-To distinguish, we only introduce collection tools tailored to the cloud platform here.
-Cross-platform Tools
+saBPF is further developed with proof-of-concept IDS and access control mechanism to demonstrate its practicability.
+ISDC (Mirnajafizadeh et al., 2024) is designed to eliminate the bottleneck between network infrastructure (where data is generated) and security application servers (where data is consumed), which prioritizes specific flows to effectively optimize resource consumption.
+4.1.4. Cross-platform Tools
 To effectively detect intrusions, an intuitive idea is to incorporate log data from various platforms to obtain a global view of the running system.
-However, it is impossible to merge these data directly due to the difference in their log formats.
 DTrace (Gregg and Mauro, 2011) is a real-time dynamic tracing framework for troubleshooting kernel and application problems on production systems.
@@
 Support for Provenance Auditing in Distributed Environments (SPADE) (Gehani and Tariq, 2012) develops a novel provenance kernel that mediates between the producers and consumers of provenance information, and handles the persistent storage of records.
-Generally, it supports heterogeneous aggregating for system-level data provenance for data analysis across multiple platforms.
+It supports heterogeneous aggregating for system-level data provenance for data analysis across multiple platforms.
 4.2. Log Storage
-After collecting logs from raw data, the subsequent step is to store these logs (Zhang et al., 2023; Ding et al., 2023a).
-Data storage involves integration, management, maintenance, and retrieval of the data, mainly focusing on metrics such as reliability and efficiency.
-The storage cannot tamper with the completeness of data, enabling developers to precisely reconstruct history scenarios.
-Additionally, when utilizing the collected data, developers should pursue low latency and high throughput for efficient data queries.
-Here, we will introduce two essential components for the data storage process: log storage systems and compression algorithms for these systems.
+The subsequent step of log collection is to store these logs (Ding et al., 2023a; Altinisik et al., 2023).
+We will introduce two essential components for data storage: log storage systems and compression algorithms for these systems.
 4.2.1. Log Storage Systems
@@
 ELK is a powerful log management solution consisting of three open-source software components: Elasticsearch (Elastic, 2010), Logstash (Elastic, 2009), and Kibana (Elastic, 2013).
-Elasticsearch is the leading distributed, RESTful search and analytics data engine designed with speed and scalability.
-In Elasticsearch, everything is indexed with finite state transducers and thus it supports rapid searches across vast repositories.
-Meanwhile, Elasticsearch implements a powerful Domain Specific Language, making it convenient and flexible to retrieve data.
-Logstash is a server-side data preprocessing pipeline to collect and integrate data from multiple sources.
-It performs filtration and transformation on its data and then sends the data to a favorable “stash” (usually Elasticsearch for ELK).
-Kibana is a data analytics and visualization platform at both speed and scale.
-It provides straightforward user interfaces to create a variety of charts and dashboards to visualize the trends, distribution, or other information of the data stored in Elasticsearch.
+Elasticsearch (Elastic, 2010) is the leading distributed, RESTful search and analytics data engine designed with speed and scalability.
+Logstash (Elastic, 2009) is a server-side data preprocessing pipeline to collect and integrate data from multiple sources.
+Kibana (Elastic, 2013) is a data analytics and visualization platform at both speed and scale.
 ELK is powerful enough to be applied in enterprise scenarios, however, its performance comes at a price.
 ELK sacrifices ease of configuration and installation, and may simultaneously introduce severe runtime overhead for its hosts.
-In contrast, Loki is a lightweight logging system with low resource overhead developed by Grafana Labs.
+In contrast, Loki (Bautista et al., 2022) is a lightweight logging system with low resource overhead developed by Grafana Labs.
 It is designed with simple operations and efficient storage.
@@
 Integrating these two tools enables Loki to construct a complete monitoring and log analysis platform for information systems.
-With the development of cloud computing, log storage systems are usually deployed on the cloud and reinforced by blockchain methods.
-Cao et al.
-(Cao et al., 2021) propose a cloud-native log database named LogStore.
-It uses highly-scalable, cost-effective cloud object storage to overcome the bandwidth limitations that arise when writing large amounts of logs.
-In their work, an optimized column index structure called LogBlock has been designed to support full-text search queries.
-Huang (Huang, 2019) proposes a log storage framework that uses the Ethereum blockchain to store the hash of log files and employs smart contracts to create indices for log files.
-Xu et al.
-(Xu et al., 2023b) propose a log data storage scheme that combines blockchain networks and IPFS technology.
-IPFS is used to store a large amount of file log data and the blockchain system is combined to achieve decentralized and secure log storage.
-Pourmajidi et al.
-(Pourmajidi and Miranskyy, 2018) propose Logchain, a blockchain-based logging system.
-It encrypts and seals logs and adds them to a hierarchical ledger to prevent log tampering, providing an immutable platform for log storage.
 4.2.2. Log Compression Algorithms
@@
 For example, it is measured that a browser can produce about 10 GB of log data each day (Ding et al., 2023a).
-Such oversize data thus should be compressed before storage.
-Typically, log compression algorithms can be categorized into two types: general-purpose algorithms and those specifically adapted to log data.
+Such oversize data should be compressed before storage.
+Log compression algorithms can be categorized into two types: general-purpose algorithms and those specifically adapted to log data.
 General Compression Algorithms
@@
 Type Well-acknoledged compression algorithm
-Dictionary-based LZ77 in gzip (Foundation, 1992), LZMA in 7zip_lzma (Pavlov, 2001), LZSS in quickLZ (QuickLZ, 2006)
-Sorting-based BWT in bzip2 (Seward, 1996), ST in szip (Schindler, 1999)
-Statistical-based PPMD in 7zip_ppmd, DMC in ocamyd (Schwellinger, 2008)
-General compression algorithms refer to algorithms to reduce the size of data while preserving the information of the data.
-It is worth noting that general compression algorithms are also capable of compressing data other than logs, namely, log data is only one type of data they can compress.
+Dictionary-based LZ77 in gzip (Foundation, 1992), LZMA in 7zip_lzma (Pavlov, 2001), and LZSS in quickLZ (QuickLZ, 2006)
+Sorting-based BWT in bzip2 (Seward, 1996) andST in szip (Schindler, 1999)
+Statistical-based PPMD in 7zip_ppmd and DMC in ocamyd (Schwellinger, 2008)
+General compression algorithms refer to algorithms to reduce the size of data (e.g., log data) by handling token-level or byte-level duplicates in the data.
 General compression algorithms can be classified into three categories based on their principles (Yao et al., 2020):
@@
 Table 2 presents representative algorithms of the above three types.
-Due to the indeterminacy of statistical techniques, this type of compression algorithm may introduce losses in compression.
+Due to the indeterminacy of statistical techniques, statistical-based compression algorithms may introduce losses in compression.
 Yet the other two types of algorithms are generally lossless.
-By evaluating the compression performance on 9 log files and 2 natural language files, a study shows that some general compression algorithms can achieve high compression ratios for log data, and log data is even easier to compress than natural language data (Yao et al., 2020).
+By validating 9 log files and 2 natural language files, a study (Yao et al., 2020) shows that some general compression algorithms can achieve high compression ratios for log data and log data is even easier to compress than natural language data.
 Tailored Compression Algorithms
-Different from natural language data, log data usually has specific structures and formal expressions.
-Therefore, compression algorithms should be adjusted for log data to improve the compression performance further.
+Different from natural language data, log data usually has specific structures and formal expressions that help further compression.
 Yao et al.
@@
 (Liu et al., 2019) propose Logzip, which employs clustering algorithms to iteratively extract templates from raw logs and then obtain coherent intermediate representations for compressing logs.
-Facing the massive data streams generated by cloud services, Lin et al.
-(Lin et al., 2015) propose Cowic, which utilizes the characteristics of logs for column compression and can independently analyze one or several columns in log entries.
-Wei et al.
-(Wei et al., 2021b) propose a parser-based log compression tool, LogReducer, which shows a high compression ratio and fast compression speed in large-scale log scenarios such as the Alibaba Cloud environment.
-They further propose LogGrep (Wei et al., 2023) for fine-grained organization of log data using static and runtime patterns.
 Rodrigues et al.
@@
 4.3. Log Parsing
-To analyze log data, DL-IDS should first parse the data into structured and standardized formats for unified processing.
-This is because log data often originates from multiple different devices such as terminals, sensors, and network devices, which will easily lead to differences in their expressions.
-The data parsing focuses on application logs and is usually executed by data classification and template extraction.
-Data classification is to classify log data into several groups through data structures such as vector representations of the data.
-Each group constitutes a template, which is used to extract features from log data and construct the structured logs.
-Basically, the existing log parsers mainly have three categories: clustering-based parsers, pattern-based parsers, and heuristic-based parsers.
-Figure 5 depicts the taxonomy of log parsers.
+Log data often originates from multiple different devices such as terminals, sensors, and network devices.
+To analyze it, log parsers are employed to format them into structured and unified ones.
+Log parsing is usually executed by data classification and template extraction.
+Data classification is to classify log data into several groups.
+Each group constitutes a template for extracting features from log data and constructing the structured logs.
+As shown in Figure 5, the existing log parsers can be taxonomized into 3 categories: clustering-based, pattern-based, and heuristic-based parsers.
 […]
@@
 4.3.1. Clustering-based Parsing.
-Clustering-based parsers classify data using clustering algorithms.
+Clustering-based parsers classify data using clustering algorithms for log parsing.
 Xiao et al.
 (Xiao et al., 2023) propose LPV, which employs a hierarchical clustering algorithm to incrementally group logs based on Euclidean distance.
-Pokharel et al.
-(Pokharel et al., 2023) convert log messages into bi-grams and directly cluster them using the DBSCAN algorithm.
 Hamooni et al.
@@
 LogCluster (Lin et al., 2016) reduces the number of logs that need to be manually checked and improves the accuracy of problem identification through log clustering and the use of knowledge bases.
-LogOHC (Yang et al., 2019) achieves efficient log template extraction by combining text vectorization and online hierarchical clustering technology.
-It has high accuracy and real-time performance and is particularly suitable for processing multi-source heterogeneous log data.
 METING (Coustié et al., 2020) provides a robust and efficient log parsing method through frequent n-gram mining and flexible log grouping strategy, which can effectively process various types of log data.
-Kimura et al.
-(Kimura et al., 2019) propose an active fault detection system based on log generation patterns.
-Active fault detection of network logs was achieved by automatically extracting log templates, constructing log feature vectors, aggregating features, and using machine learning classifiers.
 4.3.2. Frequency-based Parsing.
@@
 (Sedki et al., 2022) propose the log parsing tool ULP, which combines string matching and local frequency analysis to efficiently parse large log files.
-This method first uses text processing methods to group log events, and then conducts frequency analysis on the log instances in each group to identify fixed and variable content in the logs.
 Dai et al.
@@
 (Dai et al., 2023) further proposed an entropy-based log parser PILAR, which balances parsing accuracy and efficiency.
-Zhou et al.
-(Zhou and Su, 2023) propose the log parser Polo, which extracts log templates from logs using a prefix forest composed of a ternary search tree.
 Xu et al.
@@
 Prefix-Graph (Chu et al., 2021) is an efficient, adaptive, and universal log parsing method that can stably extract log templates without relying on domain knowledge and manual parameter tuning.
-FT-tree (Zhang et al., 2017) dynamically builds a template tree by identifying frequently appearing word combinations in system logs, supporting incremental learning.
-It is used to extract templates from switch system logs in data center networks.
 4.3.3. Heuristic-based Parsing.
@@
 (He et al., 2017) propose the online log parsing method Drain, which employs a depth-fixed parsing tree to group the original logs and encodes them using specially designed parsing rules.
-Each new log is matched layer by layer and assigned to the group with the highest similarity at the leaf node.
 Le et al.
@@
 It achieves constant parsing time and can efficiently parse raw log messages in a streaming manner.
-MoLFI (Messaoudi et al., 2018) transforms the log message format recognition problem into a multi-objective optimization problem and uses the NSGA-II algorithm to automatically generate high-quality log message templates.
 5. Intrusion Detection
-The intrusion detection stage uncovers intrusions using the semantic-level information.
+The intrusion detection stage uncovers intrusions relying on the semantic-level information.
 This section classifies and summarizes the mainstream graph summarization (Section 5.1), attack detection (Section 5.2), and attack investigation (Section 5.3).
 5.1. Graph Summarization
-It is illustrated that stealthy malware will inevitably interact with the underlying OS and be captured by provenance monitoring systems (Wang et al., 2020), which is the reason why PIDS (a form of DL-IDS) work and flourish recently.
-Log data generated from provenance monitoring system is referred to as data provenance as mentioned.
-Offering advantages in high precision, however, DL-IDS using data provenance sacrifices memory performance.
-They are required to process a vast volume of log data, since data provenance records all trails of events from their creations to their current states at a very fine-grained level, even some of which is trivial.
-Therefore, unlike network traffic and application logs used by the other forms of DL-IDS, data provenance is fine-grained, detailed, and rich in semantics.
-The aforementioned log storage systems (Section 4.2) and log compression algorithms (Section 4.2.2) are insufficient in improving the space efficiency of data provenance due to the absence of semantic information.
-They compress log data at the token level or the byte level, aiming to find duplicates that can be merged or re-represented.
-To this end, graph summarization is widely investigated to incorporate semantic-level information to further reduce the size of log data.
+It is illustrated that stealthy malware will inevitably interact with the underlying OS and be captured by provenance monitoring systems (Wang et al., 2020), which is the reason why PIDS (a form of DL-IDS) has worked and flourished recently.
+Log data generated from provenance monitoring systems is referred to as data provenance as mentioned.
+Offering advantages in high precision, data provenance sacrifices memory performance to record all trails of events from their creations to their current states, even some of which are trivial.
+Unlike network traffic and application logs, data provenance is fine-grained, detailed, and rich in semantics.
+As a result, the token-level or byte-level log storage systems (Section 4.2.1) and log compression algorithms (Section 4.2.2) are insufficient to handle the memory efficiency of data provenance due to the absence of semantic-level information.
+To this end, graph summarization is investigated to further reduce the size of log data semantically.
 In graph summarization, data provenance is transformed into a provenance graph, of which the causal relations are utilized to build the semantic understanding of system activities.
-Referring to the definition of data provenance (Definition 3.7), the definition of provenance graph can be defined as follows:
+Referring to the definition of data provenance (Definition 3.7), provenance graph is defined as follows:
 Definition 5.1.
@@
 It is a directed acyclic graph $G=<V,E>$ where nodes $V$ are system entities and edges $E$ are system events.
-Based on provenance graphs, graph summarization approaches are allowed to reduce the size of log data by directly removing irrelevant events, aggregating similar events, gathering similar execution entities, etc.
-Due to the exploitation of specific causal relations (or semantic information), graph summarization approaches are capable of performing erasion confidently and straightforwardly.
-This categorizes them as a type of lossy reduction, whereas the aforementioned log storage systems and log compression algorithms are usually lossless.
+Provenance graphs allow graph summarization approaches to reduce the size of log data by confidently removing irrelevant events, aggregating similar events, gathering similar execution entities, etc.
+This categorizes them as a type of lossy reduction, yet the aforementioned log storage and compression are usually lossless (except for statistical-based log compression).
 We note that some surveys (e.g., (Zipperle et al., 2022; Inam et al., 2023)) may interchangeably use graph summarization and log compression to identify the approaches that reduce the size of log data.
-In this work, we explicitly distinguish them and refer to the lossless reduction as compression and the opposite one as summarization, respectively.
+In this work, we explicitly distinguish them and refer to the lossless reduction as compression and the opposite one as summarization.
+Table 3 presents the overview of graph summarization approaches.
+We classify them into two categories: offline graph summarization and online graph summarization.
 Table 3. Overview of graph summarization approaches.
@@
 Winnower (Hassan et al., 2018) 2018 No Summarization None
-KCAL (Ma et al., 2018) 2018 No Summarization Instrumentation
 GS + SS (Zhu et al., 2021) 2021 FD + SD None
@@
 AudiTrim (Sun et al., 2024) 2024 CPR + GS + F-DPR None
-Table 3 presents representative data summarization approaches. In terms of their execution modes, we classify them into two categories: offline graph summarization and online graph summarization.
 5.1.1. Offline Graph Summarization
@@
 They argue that provenance graphs have similar organizational structure and characteristics to web graphs, such as locality, similarity, and consecutiveness.
-On this basis, existing web graph compression algorithms, enhanced with name-identified reference list and node-crossing gaps, are applied in provenance graphs.
 BEEP (Lee et al., 2013a) is developed based on the fact that a long-running execution can be partitioned into individual units.
 BEEP reverse engineers application binaries and instructions to perform selective logging for unit boundaries and unit dependencies.
-LogGC (Lee et al., 2013b) is an audit log system equipped with Garbage Collecting capabilities.
-During the system’s execution, LogGC can be invoked at any time to obtain the current audit logs and then generate audit logs that have been summarized.
+LogGC (Lee et al., 2013b) is a summarized audit log system that can be invoked at any time during the system execution.
 Xu et al.
-(Xu et al., 2016) initially put forward an aggregation algorithm PCR that preserves event dependencies during the data reduction process, ensuring high-quality forensic analysis.
-Subsequently, they propose a summarization algorithm PCAR and utilized domain knowledge to conduct further data summarization.
+(Xu et al., 2016) propose an aggregation algorithm PCR that preserves event dependencies during log data reduction.
+They further propose an algorithm named PCAR that utilizes domain knowledge to conduct graph summarization.
 Hossain et al.
-(Hossain et al., 2018) propose two dependency-preserving summarization approaches, FD and SD.
+(Hossain et al., 2018) propose two dependency-preserving graph summarization approaches, FD and SD.
 FD is allowed to keep backward and forward forensic analysis results.
 SD preserves the results of common forensic analysis, which runs backward to find the entry points of intrusions and then runs forward from these points to unveil their impacts.
-LogApprox (Michael et al., 2020) targets the most space-intensive events found in logs, namely file I/O activity, which can account for up to 90% of the log content.
-Once a regular expression for a given process is learned, LogApprox matches and eliminates new events that match the regular expression.
+LogApprox (Michael et al., 2020) aims to summarize the most space-intensive events found in logs, namely file I/O activity, which can account for up to 90% of the log content.
 5.1.2. Online Graph Summarization
-Online graph summarization performs real-time summarization for continually coming provenance graphs, rather than dealing with a static graph.
+Online graph summarization performs real-time summarization for continually coming provenance graphs, rather than dealing with a static provenance graph.
 ProTracer (Ma et al., 2016) alternates between system event logging and unit-level taint propagation.
 It has a lightweight kernel module and user space daemon for concurrent, out-of-order event processing.
-By combining logging and tainting, and building/evaluating a prototype, it boosts provenance tracing.
 NodeMerge (Tang et al., 2018) is a template-based graph summarization system for online event storage.
@@
 For tasks replicated across nodes in distributed applications, it can define a model over audit logs to concisely summarize the behaviors of multiple nodes, thus eliminating the necessity of transmitting redundant audit records to the central monitoring node.
-Ma et al.
-(Ma et al., 2018) propose an online audit log compression system based on kernel cache to achieve high-performance audit logging.
-It features a multi-layer cache scheme, which is distributed among various kernel data structures and utilizes the cache to detect and compress redundant events.
 The approach proposed by Zhu et al.
@@
 FAuST consists of modular parsers that parse different audit log formats to create a unified in-memory provenance graph representation.
-AudiTrim (Sun et al., 2024) is an online data summarization approach based on event dependency analysis, which can significantly reduce the amount of log data while ensuring the accuracy of intrusion detection.
+AudiTrim (Sun et al., 2024) is an efficient graph summarization approach that reduces log sizes without impacting user experiences, which allows adaptable deployment on different operating systems.
 5.2. Attack Detection
@@
 The objective of attack detection is to accurately identify malicious system events in log data while minimizing false alarms of normal system behaviors.
-Based on the types of log data used in DL-IDS, we categorize the attack detection approaches into audit log-based detectors, application log-based detectors, network traffic-based detectors, and cross-log-based detectors.
-Table 4. Overview of attack detection approaches.
-Data Type Approach Release Time Base Model Use of Graph Detection Style Detection Granularity
-Audit Pagoda (Xie et al., 2018) 2018 - Online Edge
-Wang et al. (Wang et al., 2018) 2018 LSTM - - Log Sequence
-Unicorn (Han et al., 2020) 2020 - Online Node
-Shin et al. (Shin and Kim, 2020) 2020 SVM, LR, KNN - - Sequence
-DeepHunter (Wei et al., 2021a) 2021 GNN - Node
-Fält et al. (Fält et al., 2021) 2021 Transformer - Offline Log Sequence
-ANUBIS (Anjum et al., 2022) 2022 BNN - Event Traces
-Paradise (Wu et al., 2022) 2022 - Online Node
-ShadeWatcher (Zengy et al., 2022) 2022 GNN Offline Node
-threaTrace (Wang et al., 2022a) 2022 GraphSAGE Online Node
-ProGrapher (Yang et al., 2023b) 2023 - - Subgraph
-MAGIC (Jia et al., 2024) 2024 GAT On/Offline Node, Subgraph
-Flash (Rehman et al., 2024) 2024 GNN Online Node
-Application DeepLog (Du et al., 2017) 2017 LSTM - Online Log Sequence
-LogC (Yin et al., 2020) 2020 LSTM - Online Log Sequence
-Yu et al. (Yu et al., 2021a) 2021 Bi-LSTM - - Log Sequence
-LogST (Zhang et al., 2022) 2022 Sentence-BERT - - Log Sequence
-LogBD (Liu et al., 2023) 2023 BERT - - Log Sequence
-MDFULog (Li et al., 2023a) 2023 BERT - - Log Sequence
-Traffic ACO-SVM (Meng et al., 2020) 2020 SVM - - Sequence
-Min et al. (Min et al., 2021) 2021 MemAE - - Network Traffic
-SigML++ (Trivedi et al., 2023) 2023 ANN - - Single Log
-HDLNIDS (Qazi et al., 2023) 2023 CRNN - - Network Traffic
-Audit and Application LogAnomaly (Meng et al., 2019) 2019 LSTM - Online Log Sequence
-PROV-GEM (Kapoor et al., 2021) 2021 GCN - Node
-LogBERT (Guo et al., 2021) 2021 BERT - - Log Sequence
-NeuralLog (Le and Zhang, 2021) 2021 BERT - Online Log Sequence
-GAN-EDC (Duan et al., 2021) 2021 GAN - - Single Log
-LogLS (Chen et al., 2022) 2022 Dual LSTM - Offline Log Sequence
-LogUAD (Wang et al., 2022b) 2022 Word2Vec - Offline Log Sequence
-Adanomaly (Qi et al., 2022) 2022 BiGAN - - Single Log
-LogBASA (Liao et al., 2023) 2023 BERT - - Log Sequence
-LAnoBERT (Lee et al., 2023) 2023 BERT - Online Log Sequence
-LogDAPT (Zheng et al., 2023) 2023 BERT - - Log Sequence
-ASGNet (Yang et al., 2023a) 2023 RoBERTa - - Log Sequence
-TCN-Log2Vec (Jiang et al., 2023) 2023 BERT - - Log Sequence
-The overview of attack detection approaches is presented in Table 4 and its corresponding taxonomy is depicted in Figure 6.
-We note that recent years have also published many research papers for attack detection (Li et al., 2024a, 2023c; Xu et al., 2022; Hassan et al., 2020; Milajerdi et al., 2019; Yu et al., 2021b; Chen et al., 2024; El Khairi et al., 2024; Wang et al., 2023).
-Yet these papers do not belong to DL-IDS, which are thus excluded hereafter in our survey.
+Based on the types of log data, we categorize the attack detection approaches into audit log-based, application log-based, network traffic-based, and hybrid log-based detectors.
+Table 4. Overview and taxonomy of attack detection approaches.
+Data Type Taxonomy Approach Release Time Base Model Detection Style Detection Granularity
+Traditional StreamSpot (Manzoor et al., 2016) 2018 K-Medoids Online Subgraph
+Learning Unicorn (Han et al., 2020) 2020 K-Medoids Online Node, Subgraph
+DistDet (Dong et al., 2023) 2023 HST Online Subgraph
+Audit ShadeWatcher (Zengy et al., 2022) 2022 TransR Offline Node
+Log Graph threaTrace (Wang et al., 2022) 2022 GraphSAGE Online Node
+Neural ProGrapher (Yang et al., 2023) 2023 graph2vec Online Subgraph
+Network MAGIC (Jia et al., 2024) 2024 GAT Online Node, Subgraph
+Flash (Rehman et al., 2024) 2024 GraphSAGE Online Node
+R-caid (Goyal et al., 2024) 2024 GNN Offline Node
+Argus (Xu et al., 2024b) 2024 MPNN, GRU - Node
+Traditional Wei et al. (Xu et al., 2009) 2009 PCA, TF-IDF - Log Entry
+Learning Bodik et al. (Bodik et al., 2010) 2010 Logistic Regression Online Log Entry
+DeepLog (Du et al., 2017) 2017 LSTM Online Log Entry
+LogRobust (Zhang et al., 2019) 2019 Attention LSTM - Log Entry
+Application Sequence LogAnomaly (Meng et al., 2019) 2019 template2vec, LSTM Online Log Entry
+Log Neural LogC (Yin et al., 2020) 2020 LSTM Online Log Entry
+Network NeuralLog (Le and Zhang, 2021) 2021 BERT - Log Entry
+PLELog (Yang et al., 2021) 2021 Attention GRU Online Log Entry
+SpikeLog (Qi et al., 2023) 2023 DSNN - Log Entry
+LogCraft (Zhang et al., 2024) 2024 Meta Learning - Log Entry
+Traffic Whisper (Fu et al., 2021) 2021 K-Means - Host
+Traditional SigML++ (Trivedi et al., 2023) 2023 ANN - Encrypted Log
+Learning OADSD (Zhang et al., 2023a) 2023 Isolation Forest Online Packet
+LtRFT (Tang et al., 2023) 2023 LambdaMART Offline Packet
+Kitsune (Mirsky et al., 2018) 2018 AutoEncoder Online Packet
+MT-FlowFormer (Zhao et al., 2022a) 2022 Transformer - Flow
+I2RNN (Song et al., 2022) 2022 I2RNN - Packet
+Graph and ERNN (Zhao et al., 2022b) 2022 ERNN - Flow
+Sequence Euler (King and Huang, 2023) 2023 GNN, RNN - Flow
+Neural pVoxel (Fu et al., 2023) 2023 - - Packet, Flow
+Network NetVigil (Hsieh et al., 2024) 2024 E-GraphSage - Flow
+Exosphere (Fu et al., 2024) 2024 CNN - Packet
+DFNet (Zhao et al., 2024a) 2024 DFNet - Packet
+RFH-HELAD (Zhong et al., 2024) 2024 RPGAN, Deep kNN - Packet
+Hybrid Hybrid OWAD (Han et al., 2023) 2024 Autoencoder Online Hybrid
+FG-CIBGC (Ni et al., 2025) 2025 DisenGCN, ICL - Hybrid
+The overview and taxonomy of attack detection approaches are presented in Table 4.
+We note that recent years have also published many other academic papers for attack detection (Xie et al., 2018; Wu et al., 2022; Li et al., 2023b; Milajerdi et al., 2019a; Yu et al., 2021; Chen et al., 2024; El Khairi et al., 2024; Wang et al., 2023; Hassan et al., 2020a).
+Yet these papers are slightly related to DL-IDS, which are thus excluded in our survey for conciseness.
 5.2.1. Audit Log-based Detectors
-Audit logs are collected from hosts and provide atomic information to construct data provenance.
-In terms of learning techniques, detectors based on audit logs can be classified as traditional learning, graph neural network, and sequence neural network.
+Audit logs are collected from hosts and thus detectors based on them are basically referred to as HIDS.
+Audit logs provide fine-grained information through provenance graphs to depict system behaviors.
+Depending on the learning techniques, audit log-based detectors can be further classified as traditional learning and graph neural network.
 Traditional Learning
-Traditional learning-based detectors refer to those that utilize fundamental learning techniques.
-Shin et al.
-(Shin and Kim, 2020) propose to employ diverse machine learning algorithms to categorize benign and malicious logs.
-Pagoda (Xie et al., 2018) takes into account the degree of abnormality of individual provenance paths and the entire provenance graph.
-It can quickly uncover intrusions if a serious compromise has been identified on paths.
-The detection rate of Pagoda is further improved by considering the behavior representation in the whole provenance graph.
+Traditional learning-based detectors refer to those that utilize naive machine learning techniques.
+StreamSpot (Manzoor et al., 2016) is a clustering-based anomaly detection that tackles challenges in heterogeneity and streaming nature.
 Unicorn (Han et al., 2020) is a real-time intrusion detector that efficiently constructs a streaming histogram to represent the history of system executions.
-The counting results of the histogram are updated immediately if new edges are generated.
-By exploring the neighborhood relationships of provenance graphs, Unicorn discovers the causal relationships of system entities in the context.
-Similarly, ANUBIS (Anjum et al., 2022) also employs provenance graphs to capture causal relationships to achieve high detection performance.
-The core predictive capability of ANUBIS is a Bayesian Neural Network, which can inform users about how confident it is in its predictions.
-Paradise (Wu et al., 2022) introduces a novel extraction strategy at the system log level to prune and extract process feature vectors, which are then stored in an efficient in-memory database.
-Paradise is independent of types of operating systems or provenance collection frameworks, and is capable of independently calculating provenance-based dependencies at the detection stage.
-[…]
-Figure 6. Taxonomy of attack detection approaches.
+The counting results within the histogram are updated immediately if new edges (or events) occur.
+DistDet (Dong et al., 2023) is a distributed detection system that builds host models in the client side, filters false alarms based on their semantics, and derives global models to complement the host models.
 Graph Neural Network
-Graph Neural Network (GNN) is demonstrated to do well in graph-structured DL-IDS (Wei et al., 2021a; Yang et al., 2023b; Jia et al., 2024; Rehman et al., 2024; Zengy et al., 2022).
-DeepHunter (Wei et al., 2021a) is a GNN-based graph matching approach that is designed with two novel networks: attribute embedding networks that could incorporate Indicators of Compromise (IoC) information and graph embedding networks that could capture the relationships between IOCs.
-threaTrace (Wang et al., 2022a) emerges as an online approach dedicated to detecting and tracing host-level threats.
-Its model is custom-tailored on the basis of GraphSAGE (Hamilton et al., 2017), an inductive graph neural network framework, aiming to learn the role of every benign entity within the data source graph.
-MAGIC (Jia et al., 2024) leverages Graph Attention Network (GAT) (Velickovic et al., 2017) as a crucial constituent of its graph representation module.
-It employs masked graph representation learning to mimic benign system events and executes multi-granularity detection under diverse supervision levels.
-MAGIC can be applied in both online detection and offline detection scenarios.
-MAGIC is able to adapt to concept drift and has minimal computational overhead, which makes it applicable to real-world online APT detection.
-ProGrapher (Yang et al., 2023b) is a graph embedding-based anomaly detection system that extracts features and detects anomalies by utilizing information in relational networks.
-ShadeWatcher (Zengy et al., 2022) is a recommendation-guided detector.
-It models high-order connectivity via item-side information, borrowing the recommendation concepts of user-item interactions and applying them to the security concepts of system entity interactions.
-Sequence Neural Network
-All logs can be formatted as sequential data.
-Hence, sequence neural networks such as Recurrent Neural Network (RNN) and Transformer are employed in some work (Wang et al., 2018; Fält et al., 2021).
-The approach of Wang et al.
-(Wang et al., 2018) employs Long Short-Term Memory (LSTM), a variant of RNN, for anomaly detection.
-In their work, LSTM is compared with other machine learning algorithms such as Gradient Boosting Decision Tree and Naive Bayes.
-The results indicate that LSTM can attain the optimal anomaly detection performance when utilizing two feature extraction methods.
-Fält et al.
-(Fält et al., 2021) present a log anomaly detection method that is entirely trained based on auxiliary system data and is unsupervised with respect to the target system.
-It utilizes transformers as its base model, whose performance results reveal great potential of the self-attention mechanism in the field of DL-IDS.
+GNN is demonstrated to do well in processing provenance graphs (Yang et al., 2023; Jia et al., 2024; Rehman et al., 2024; Zengy et al., 2022; Wang et al., 2022).
+ProGrapher (Yang et al., 2023) extracts temporal-ordered provenance graph snapshots from the ingested logs, and applies whole graph embedding and sequence-based learning to capture rich structural properties of them.
+The key GNN technique leveraged by ProGrapher is graph2vec.
+ShadeWatcher (Zengy et al., 2022) is a recommendation-guided intrusion detector using provenance graphs.
+It borrows the recommendation concepts of user-item interactions into security concepts of system entity interactions and analyzes cyber threats in an automated and adaptive manner.
+threaTrace (Wang et al., 2022) emerges as an online approach dedicated to detecting host-based threats at the node level.
+Its GNN model is a tailored GraphSAGE (Hamilton et al., 2017) for learning rich contextual information in provenance graphs.
+MAGIC (Jia et al., 2024) leverages Graph Attention Network (GAT) (Velickovic et al., 2017) as its graph representation module.
+MAGIC employs masked graph representation learning to incorporate the capability of pre-training.
+It can adapt to concept drift with minimal computational overhead, making it applicable to real-world online APT detection.
+Flash (Rehman et al., 2024) is a comprehensive and scalable approach on data provenance graphs to overcome the limitations in accuracy, practicality, and scalability.
+Flash incorporates a novel adaptation of a GNN-based contextual encoder to encode both local and global graph structures into node embeddings efficiently.
+R-caid (Goyal et al., 2024) first incorporates root cause analysis into PIDS.
+Before training GNNs, R-caid links nodes to their root causes to build a new graph, intending to prevent it from mimicry and evasion attacks.
+Argus (Xu et al., 2024b) finds the performance of the prior IDS is questionable on large scale.
+It thus devises a form of discrete temporal graph and uses encoder-decoder unsupervised learning to detect different types of attacks.
 5.2.2. Application Log-based Detectors
 Application logs are generated from the installed binaries.
-Generally, application logs are in the form of natural language text, as the generated templates of them are pre-defined by their developers.
-Hence, many Natural Language Processing (NLP) techniques are introduced in this type of DL-IDS.
-Recurrent Neural Network
-As one of the most classic methods, RNN is used commonly in DL-IDS.
+Generally, application logs are in the form of natural language text, namely sequence data.
+It is thus common to introduce sequence-based DL techniques into application log-based DL-IDS.
+Traditional Learning
+For traditional learning, Wei et al.
+(Xu et al., 2009) propose a general methodology to mine rich semantic information in console logs to detect large-scale system problems.
+Bodik et al.
+(Bodik et al., 2010) leverage a logistic regression model on a new and efficient representation of a datacenter’s state called fingerprint to detect previously seen performance crises in that datacenter.
+Sequence Neural Network
+Due to the similarity between application logs and natural language texts, sequence neural networks such as Recurrent Neural Network (Hochreiter and Schmidhuber, 1997) and Transformer (Vaswani et al., 2017; Devlin et al., 2019) are widely employed.
 DeepLog (Du et al., 2017) employs LSTM to model system logs as natural language sequences.
-As a result, it is able to automatically learn log patterns from normal operations and detect anomalies when there is a deviation between log patterns and the trained model.
-LogC (Yin et al., 2020) comprises two stages: transforming log messages into log template sequences and component sequences.
-These two sequences are then input into a combined LSTM model for the detection of abnormal logs.
-LogC only requires normal log sequences to train the combined model.
-Both DeepLog and LogC are capable of conducting online detection.
-Yu et al.
-(Yu et al., 2021a) propose a novel framework for the automatic detection of log anomalies, which leverages a bidirectional LSTM (Bi-LSTM) model based on an attention mechanism.
-In contrast to existing log anomaly detection methods, this approach not only pays attention to the sequential and quantitative information of log sequences but also considers the hidden semantic information in logs, thus enhancing the efficiency of anomaly detection.
-Transformer Architecture
-Transformer (Vaswani et al., 2017) is an effective model for sequence data, which exploits a multi-head self-attention mechanism to realize a large model that can be parallelly computed.
-The most common Transformer-based model in DL-IDS is BERT (Devlin et al., 2019), which takes advantage of encoding sequence data and pre-training model parameters.
-LogBD (Liu et al., 2023) utilizes BERT to acquire the semantic information of logs and constructs distance-based anomaly detection via a domain adaptation approach.
-It employs a Temporal Convolutional Network (TCN) to extract the common characteristics of different system logs and map them into the same hypersphere space.
-MDFULog (Li et al., 2023a) introduces a BERT-based semantic feature extraction model to preserve the semantics of log messages and map them to log vectors.
-This effectively eliminates the work randomness and noise injection caused by log template updates and enhances the robustness of anomaly detection.
-LogST (Zhang et al., 2022) obtains the semantic representation of log events through the SBERT model, taking into account the semantics of each word and the word order relationship, thus facilitating a better understanding of the context of log sequences.
+It is able to automatically learn benign log patterns and detect anomalies when there is a deviation between log patterns and the trained model.
+LogRobust (Zhang et al., 2019) finds previous methods do not work well under the close-world assumption and utilizes an attention-based LSTM model to handle unstable log events and sequences.
+LogAnomaly (Meng et al., 2019) identifies previous studies tend to cause false alarms by using indexes rather than semantics of log templates.
+Empowered by a novel, simple yet effective method termed template2vec, LogAnomaly is proven to successfully detect both sequential and quantitive log anomalies simultaneously.
+LogC (Yin et al., 2020) is a new log-based anomaly detection approach with component-aware analysis.
+It feeds both log template sequences and component sequences to train a combined LSTM model for detecting anomalous logs.
+NeuralLog (Le and Zhang, 2021) targets the performance caused by log parsing errors such as out-of-vocabulary words and semantic misunderstandings and employ BERT to perform neural representation.
+PLELog (Yang et al., 2021) is a semi-supervised anomaly detection approach that can get rid of time-consuming manual labeling and incorporate the knowledge on historical anomalies.
+SpikeLog (Qi et al., 2023) adopts a weakly supervised approach to train an anomaly score model, with the objective of handling a more reasonable premise scenario where a large number of logs are unlabeled.
+LogCraft (Zhang et al., 2024) is an end-to-end unsupervised log anomaly detection framework based on automated machine learning, which mitigates the cost of understanding datasets and makes multiple attempts for building algorithms.
 5.2.3. Network Traffic-based Detectors
-Network traffic comes from communications between hosts across a computer network. It is ruled by network protocols such as Transmission Control Protocol (TCP) and the User Datagram Protocol (UDP).
+Network traffic comes from communications between hosts across a computer network.
+It is ruled by network protocols such as Transmission Control Protocol (TCP) and the User Datagram Protocol (UDP) and can be utilized for intrusion detection.
+Basically, network traffic-based detectors are termed NIDS.
 Traditional Learning
-ACO-SVM (Meng et al., 2020) is a network intrusion detection model that optimizes the parameters of SVM based on the ant colony algorithm.
-In this approach, the parameters of the SVM are regarded as the paths traversed by ants.
-Through collecting network status information, performing feature extraction and processing, and utilizing the pheromone update and node transfer of the ant colony algorithm to achieve path crawling, the optimal parameter combination of ACO-SVM is found.
-Min et al.
-(Min et al., 2021) propose an approach for network anomaly detection employing a memory-augmented deep autoencoder.
-This model learns the prototype patterns of normal inputs via the memory module and reconstructs the inputs of abnormal samples that are close to normal samples during training.
-Thus, even the latent vectors of abnormal inputs are transformed through the memory module and aggregated and reconstructed from similar normal samples learned from the memory that is learned only with normal samples, thereby resolving the issue of over-generalization.
-SigML++ (Trivedi et al., 2023) is a supervised anomaly detection approach featuring probabilistic polynomial approximation.
-It employs Fully Homomorphic Encryption and Artificial Neural Networks to approximate the sigmoid activation function, enabling the execution of anomaly detection without decrypting logs.
-Recurrent Neural Network
-HDLNIDS (Qazi et al., 2023) employs a Convolutional Recurrent Neural Network (CRNN) for the detection and classification of malicious network traffic, with the objective of enhancing the detection efficiency and accuracy of existing intrusion detection systems.
-It conducts local feature extraction via CNN and then utilizes the RNN layer to capture the temporal dependency of features, thereby enabling a more comprehensive identification of network traffic patterns.
-5.2.4. Cross Log-based Detectors
-Audit logs contain detailed system-level underlying information, network traffic logs depict interactions between systems, and application logs possess high-level semantic information.
-Employing various kinds of logs in DL-IDS can enhance the interpretability of logs and obtain a more comprehensive understanding of intrusions.
-Traditional Learning
-LogUAD (Wang et al., 2022b) is a method for offline detection.
-It does not require a log parsing procedure.
-LogUAD directly takes the original log message as input to avert the noise brought in by parsing.
-It uses Word2Vec to generate word vectors and combines Term Frequency-Inverse Document Frequency (TF-IDF) to generate weighted log sequence feature vectors for handling the evolution of log statements.
-Ultimately, it makes use of a computationally efficient unsupervised clustering algorithm to detect anomalies.
-Recurrent Neural Network
-LogAnomaly (Meng et al., 2019) constitutes a system log anomaly detection framework.
-The central notion of this framework is to model the log stream as a natural language sequence.
-LogAnomaly employs a pattern of offline learning and online detection.
-In the offline learning segment, LogAnomaly utilizes the FT-Tree to extract templates from historical logs and matches the historical logs with these templates.
-Subsequently, each template is transformed into a template vector sequence, and sequential and quantitative features are extracted via the LSTM model.
-In the online detection portion, LogAnomaly initially determines whether real-time logs can be matched with existing templates.
-If affirmative, it is converted into a template vector; otherwise, based on the similarity among template vectors, it will be approximated to an existing template vector.
-LogLS (Chen et al., 2022) is a system log anomaly detection approach based on symmetrically structured double-long short-term memory.
-LogLS is an optimization of the DeepLog method referred to in Section 5.2.2.
-It is designed to tackle the gradient issue of LSTM in long sequence prediction and accomplishes the prediction of unobserved logs by providing a feedback mechanism.
-Transformer Architecture
-Both NeuralLog (Le and Zhang, 2021) and LAnoBERT (Lee et al., 2023) are free from the need for a log parser.
-They are able to learn directly from the raw log data, thereby avoiding the issue of potentially losing crucial information during the standardization process.
-They both make use of BERT to encode the semantic significance of log messages and are capable of conducting online detection.
-The work presented in the paper LogDAPT (Zheng et al., 2023) is predominantly founded on NeuralLog.
-It is noted that NeuralLog necessitates a substantial amount of labeled data to yield satisfactory outcomes.
-In contrast, LogDAPT tackles the issue of insufficient labeled data by applying the second stage of in-domain pre-training, namely DAPT, on the BERT base model.
-Not only LogBERT (Guo et al., 2021) but also LogBASA (Liao et al., 2023) employ the log parser Drain to parse log events.
-Subsequently, they make use of the pre-trained BERT model to extract the semantic information of the events.
-In the TCN-Log2Vec (Jiang et al., 2023) framework, the log parsing component is optimized on the basis of Drain.
-Meanwhile, the BERT model is employed to precisely extract semantics from log templates.
-Additionally, TF-IDF is utilized to capture the context of log templates, and window-based techniques are adopted to obtain log sequence information.
-The ASGNet (Yang et al., 2023a) model is composed of three components: log statistics information representation (V-Net), log deep semantic representation (S-Net), and an adaptive semantic threshold mechanism (G-Net).
-V-Net employs an unsupervised variational autoencoder (VAE) to learn the global representation of each statistical feature vector.
-S-Net extracts the latent semantic representation of text input via a pre-trained RoBERTa model.
-G-Net aligns the information from the two sources and then adjusts the information flow.
-Generative Adversarial Network
-GAN-EDC (Duan et al., 2021) and Adanomaly (Qi et al., 2022) utilize technologies related to Generative Adversarial Network (GAN).
-GAN-EDC employs an LSTM-based encoder-decoder framework as the generator.
-Log keywords are taken as the input of the encoder, and the decoder outputs the generated log template.
-The discriminator utilizes CNN to identify the disparities between the generated log template and the real log template.
-Adanomaly is a log anomaly detection method based on Bidirectional GAN (BiGAN) and the stacking approach.
-It uses a CNN-based BiGAN model to extract features.
-Subsequently, it trains multiple classifiers by partitioning the extracted features into multiple subsets and integrates the classifiers using the stacking method.
-Graph Neural Network
-As data provenance graphs are complex datasets, typically represented as heterogeneous graphs and multiplex networks, PROV-GEM (Kapoor et al., 2021) employs a GCN-based approach.
-It is specially designed for attribute heterogeneous multiplex networks and is able to handle large networks and the heterogeneity of data provenance graphs.
+Given the fact that network traffic is usually encrypted for secure communications, feature engineering-guided machine learning is widely applied in NIDS.
+Whisper (Fu et al., 2021) pays attention to both high accuracy and high throughput by utilizing frequency domain features.
+SigML++ (Trivedi et al., 2023) is an extension of SigML for supervised anomaly detection approach.
+SigML++ employs Fully Homomorphic Encryption and Artificial Neural Network (ANN) for detection, resulting in execution without decrypting the logs.
+OADSD (Zhang et al., 2023a) achieves task independently and has the ability of adapting to the environment over SD-WAN by using On-demand Evolving Isolation Forest.
+LtRFT (Tang et al., 2023) innovatively introduces Learning-To-Rank scheme for mitigating the low-rate DDoS attacks targeted at flow tables.
+Graph and Sequence Neural Network
+In network traffic, packets consist of various contents and their flows can be represented as graphs.
+As a result, both graph neural network and sequence neural network are adopted in NIDS.
+Kitsune (Mirsky et al., 2018) is a plug and play NIDS that is allowed to detect attacks efficiently on the local network without supervision.
+It alleviates the problem that network gateways and router devices simply do not have the memory or processing power.
+MT-FlowFormer (Zhao et al., 2022a) is a semi-supervised framework to mitigate the lack of a mechanism for modeling correlations between flows and the requirement of a large volume of manually labeled data.
+I2RNN (Song et al., 2022) is an incremental and interpretable RNN for encrypted traffic classification, which can be efficiently adapted for incremental traffic types.
+ERNN (Zhao et al., 2022b) represents error-resilient RNN, which is a robust and end-to-end RNN model specially designed against network-induced phenomena.
+Euler (King and Huang, 2023) accelerates the most memory-intensive part, message-passing stage within GNN, with several concurrently-executed replicated GNNs.
+pVoxel (Fu et al., 2023) is an unsupervised method that proposes to leverage point cloud analysis to reduce false positives for the previous NIDS such as Whisper and Kitsune without requiring any prior knowledge on the alarms.
+NetVigil (Hsieh et al., 2024) is specially designed for east-west traffic within data center networks.
+It utilizes E-GraphSage and contrastive learning techniques to strengthen its resilience.
+Exosphere (Fu et al., 2024) detects flooding attacks by analyzing packet length patterns, without investigating any information in encrypted packets.
+DFNet (Zhao et al., 2024a) is a DDoS prevention paradigm denoted by preference-driven and in-network enforced shaping.
+RFH-HELAD (Zhong et al., 2024) consists of a $K$ classification model based on a deep neural network and a $K+1$ classification combining GAN and Deep kNN for detecting anomalies in network traffic.
+5.2.4. Hybrid Log-based Detectors
+Based on the above discussions, a natural idea is to combine various types of log data for improving detection capability.
+OWAD (Han et al., 2023) is a general framework to detect, explain, and adapt to normality shifts in practice.
+OWAD is validated to be effective in various detection granularity, covering provenance graphs, application logs, and network packets.
+FG-CIBGC (Ni et al., 2025) mines syncretic semantics in multi-source logs including audit logs, application logs, and network traffic using LLM under in-context learning, which generates behavior graphs for comprehensive analysis.
 5.3. Attack Investigation
-Table 5. Overview of attack investigation approaches.
-Approach Release Time Audit Log Applica- tion Log Use of Graph Base Model Starting Node Investigation Granularity
-ProvDetector (Wang et al., 2020) 2020 - Doc2Vec - Path
-ATLAS (Alsaheel et al., 2021) 2021 - LSTM Sequence
-Liu et al. (Liu et al., 2022a) 2022 - Struc2Vec Provenance Graph
-LogTracer (Niu et al., 2022) 2022 - DeepLog Path
-ConLBS (Li et al., 2023b) 2023 - Transformer Provenance Graph
-AirTag (Ding et al., 2023b) 2023 - BERT - Sequence
-Karios (Cheng et al., 2024) 2023 GNN - Summary Graph
-CAPTAIN (Ramaki et al., 2023) 2023 - - Provenance Graph
-Except for identifying individual compromised nodes, IDS are supposed to unveil the full story of intrusions (a.k.a., attack scenarios).
-This process is referred to as attack investigation, which can be done by directly detecting attack scenarios (Wang et al., 2020), or analyzing the causal relations between compromised nodes (Alsaheel et al., 2021; Ding et al., 2023b).
-The attack scenarios are defined with scenario graphs as follows:
+Except for identifying individual intrusive nodes, IDS are supposed to detect the full story of intrusions (a.k.a., attack scenario graphs).
+This process is referred to as attack investigation, which can be done by directly detecting attack scenario graphs (Wang et al., 2020), or analyzing the causal relations between compromised nodes progressively to construct attack scenario graphs (Alsaheel et al., 2021; Ding et al., 2023b; Xu et al., 2022; Jiang et al., 2025).
+The attack scenario graphs are defined with scenario graphs as follows:
 Definition 5.2.
@@
 Definition 5.3.
-(Attack Scenario). Attack scenario is a form of scenario graph, where its nodes of interest are compromised nodes.
-In the past, attack investigation is conducted by forward analysis and backward analysis.
+(Attack Scenario Graph). Attack scenario graph is a scenario graph where its nodes of interest are compromised nodes.
+In the past, attack investigation is conducted by forward analysis and backward analysis (Hossain et al., 2017).
 Forward analysis discovers the influence that nodes of interest will cause and backward analysis traces back how nodes of interest are generated.
-Benefiting from DL techniques, both forward and backward analysis can be achieved by learning patterns and predicting possible attack scenarios.
+Benefiting from DL techniques, both forward and backward analysis can be achieved by learning patterns of attack scenario graphs.
 Table 5 summarizes the overview of attack investigation approaches.
+Similar to Section 5.2, we exclude papers (Fang et al., 2022; Gao et al., 2018; Ma et al., 2017; Li et al., 2024a; Milajerdi et al., 2019b; Hassan et al., 2020b; Yang et al., 2020; Kwon et al., 2018; Alhanahnah et al., 2022; Ji et al., 2017; Zhu et al., 2023b; Hossain et al., 2017; Wang et al., 2023) slightly relevant to DL for conciseness.
+Table 5. Overview of attack investigation approaches.
+Taxonomy Approach Release Time Audit Log Applica- tion Log Base Model Starting Node Investigation Granularity
+Traditional ProvDetector (Wang et al., 2020) 2020 doc2vec Path
+Sequence ATLAS (Alsaheel et al., 2021) 2021 LSTM Graph
+Neural LogTracer (Niu et al., 2022) 2022 DeepLog Path
+Network ConLBS (Li et al., 2023a) 2023 Transformer Graph
+AirTag (Ding et al., 2023b) 2023 BERT Graph
+Graph Liu et al. (Liu et al., 2022a) 2022 struc2vec Graph
+Neural Karios (Cheng et al., 2024) 2023 GNN Graph
+Network TREC (Lv et al., 2024) 2024 GNN Graph
+Orthrus (Jiang et al., 2025) 2025 UniMP Path
 Traditional Learning
+Unlike detecting intrusive nodes, attack scenario graphs are complicated and thus are hard to handle by traditional learning methods.
 ProvDetector (Wang et al., 2020) utilizes doc2vec to learn the embedding representation of paths in the provenance graph.
-This method enables the mapping of different components in the path (e.g., processes, files, and network connections) to a multidimensional numerical vector space.
-CAPTAIN (Ramaki et al., 2023) is a community-based APT analyzer in IT networks to detect Intrusion Kill Chain (IKC).
-CAPTAIN consists of a total of 12 different activities in two stages to detect possible IKCs of the APT attacks.
-CAPTAIN uses graph-based algorithms to model relationships among logged events.
-It detects event communities and summarizes them with centrality measures.
-By considering the nature of included events, it identifies malicious communities related to IKC stages.
-This ultimately leads to the reconstruction of the IKC of APT attacks.
-Recurrent Neural Network
-ATLAS (Alsaheel et al., 2021) is a framework that constructs end-to-end attack stories from readily available audit logs.
-It employs a novel combination of causal analysis, natural language processing, and machine learning techniques to build a sequence-based model.
-In ATLAS, LSTM enables the model to automatically learn to distinguish patterns in attack and non-attack sequences.
-This model establishes key patterns of attack and non-attack behaviors from causal graphs.
-LogTracer (Niu et al., 2022) is an efficient log-based anomaly tracing method.
-In attack path extraction, it weights edges of the initial provenance graph with an outlier algorithm, filters edges by threshold, and restructures the scenario graph.
-It finds the longest path as an attack path by assigning anomaly degrees to edges.
-It also uses an incremental graph processing strategy and the DeepLog model to process coarse-grained log information to assist the tracing process.
-Transformer Architecture
-AirTag (Ding et al., 2023b) employs unsupervised learning to train DL models directly from log texts rather than relying on labeled causal graphs.
-It uses the BERT model for unsupervised learning and retrains the BERT model based on unlabeled log data.
-ConLBS (Li et al., 2023b) combines a contrastive learning framework and multilayer Transformer network for behavior sequence classification.
-It constructs behavior sequences from audit logs, uses a lemmatization strategy to map semantics to the attack pattern layer, and can perform unsupervised representation learning on unlabeled sequences.
+Then a density-based detection is deployed to detect abnormal causal paths in the provenance graph.
+Sequence Neural Network
+Log data is in the form of natural language text or is allowed to be transformed into sequences of events, which facilitates the introduction of sequence neural networks.
+ATLAS (Alsaheel et al., 2021) is a framework to construct end-to-end attack stories from readily available audit logs, which employs a novel combination of causal analysis and natural language processing.
+ATLAS exploits LSTM to automatically learn the pattern difference between attack and non-attack sequences.
+LogTracer (Niu et al., 2022) is an efficient anomaly tracing framework that combines data provenance and system log detection together.
+An outlier function with an abnormal decay rate is introduced to improve the accuracy.
+ConLBS (Li et al., 2023a) combines a contrastive learning framework and multilayer Transformer network for behavior sequence classification.
+AirTag (Ding et al., 2023b) employs unsupervised learning to train BERT directly from log texts rather than relying on provenance graphs.
+AirTag constructs attack scenario graphs by integrating the detected victim nodes.
 Graph Neural Network
+To capture causal relations within graphs, GNN is commonly adopted.
 Liu et al.
 (Liu et al., 2022a) propose an automated attack detection and investigation method via learning the context semantics of the provenance graph.
-The provenance graph analyzes temporal and causal dependencies of system events.
-The pre-trained Struc2Vec model is used to learn representations of unique nodes in a general set for node embedding.
+The provenance graph analyzed by struc2vec captures temporal and causal dependencies of system events.
 Kairos (Cheng et al., 2024) is a practical intrusion detection and investigation tool based on whole-system provenance.
-Kairos utilizes GNN to analyze system execution history and detect and reconstruct complex APTs.
-It employs a graph neural network-based encoder-decoder architecture to learn the temporal evolution of provenance graph structure changes and quantify the abnormal degree of each system event.
-[…]
-Figure 7. Taxonomy of attack investigation.
+Kairos utilizes GNN to analyze system execution history, so that detects and reconstructs complex APTs.
+It employs a GNN-based encoder-decoder architecture to learn the temporal evolution of provenance graph structure changes and quantify the abnormal degree of each system event.
+TREC (Lv et al., 2024) abstracts APT attack investigation problem as a tactics / techniques recognition problem.
+TREC trains its model in a few-shot learning manner by adopting a Siamese neural network.
+Orthurus (Jiang et al., 2025) identifies Quality of Attribution as the key factor contributing to whether or not the industry adopts IDS.
+It first detects malicious hosts using a GNN encoder and then reconstructs the attack path through dependency analysis.
 6. Benchmark Datasets
@@
 6.1. Dimensions of Datasets
-To illustrate the quality of datasets, it is general to use the following dimensions:
+To illustrate the quality of DL-IDS datasets, it is general to use the following dimensions:
   • Benign Scenarios: Benign data should cover benign behaviors and system activities to the greatest extent, enabling DL-IDS to learn patterns of benign behaviors to differentiate malicious behaviors.
@@
 CERT Dataset (Glasser and Lindauer, 2013; Lindauer, 2020) 2020 87 GB - Yes .csv W
-Loghub (Zhu et al., 2023) 2020 77 GB - - .txt W, L, M, S
+LogChunks (Brandt et al., 2020) 2020 24.1 MB - Yes .txt -
+Loghub (Zhu et al., 2023a) 2020 77 GB - - .txt W, L, M, S
 ProvSec (Shrestha et al., 2023) 2023 - 11 Yes .json L
@@
 CERT dataset learned important lessons about the benefits and limitations of synthetic data in the cybersecurity domain and carefully discussed models of realism for synthetic data.
-Loghub dataset (Zhu et al., 2023) is a large collection of system log datasets, providing 19 real-world log data from various software systems, including distributed systems, supercomputers, operating systems, mobile systems, server applications, and standalone software.
+LogChunks (Brandt et al., 2020) is an application log dataset for build log analysis, containing 797 annotated Travis CI build logs from 80 GitHub repositories and 29 programming languages.
+These logs are from mature and popular projects, collected through repository, build, and log sampling.
+Each log in the dataset has manually labeled text blocks of build failure reasons, search keywords, and structural categories, and cross-validated with the original developers with an accuracy of 94.4%.
+Loghub dataset (Zhu et al., 2023a) is a large collection of system log datasets, providing 19 real-world log data from various software systems, including distributed systems, supercomputers, operating systems, mobile systems, server applications, and standalone software.
 The objective of Loghub is to fill the significant gap between intelligent automated log analysis techniques and successful deployments in the industry.
@@
 This leads to unreliable descriptions of system behaviors that may mislead DL-IDS.
-For example, in some cases of the DARPA TC dataset, the PROCESS object and its source fail to properly resolve conflicts, resulting in possible incorrect entries in the actor_id field.
-And the acuity_level value of the FLOW object is 0, while the value range for this field in other objects is from 1 to 5.
+For example, in some cases of the DARPA TC dataset, the PROCESS object and its source fail to properly resolve conflicts, resulting in possible incorrect transformation.
+Besides, the acuity_level value of the FLOW object is 0, while the value range for this field in other objects is from 1 to 5.
 Another example could be the LogChunks (Brandt et al., 2020) dataset.
@@
 Meanwhile, high-quality ground-truth labels are hard to acquire, which is impeded by the contradiction between fine-grained manual labeling and automated label generation.
-On one hand, for unknown intrusions such as zero-day attacks, security analysts usually have no trouble recovering attack scenarios, while it is very labor-intensive to correspond each attack scenario to certain log entries.
-The DAPRA TC dataset (DARPA, 2022) is a typical example of this phenomenon.
+On one hand, for unknown intrusions such as zero-day attacks, it is very labor-intensive for security analysts to correspond each attack scenario to certain log entries, although coarse-grained attack scenarios may have been acquired.
+The DAPRA TC dataset (DARPA, 2022) is a typical example for this.
 It only provides a ground truth report for attack scenarios, which does not correspond to any specific log entries.
-Although a few researchers (Wang et al., 2022a) provide the third-party ground-truth labels that are manually identified by themselves, we empirically find some ambiguities between their ground-truth labels and the official attack scenario report.
+Although a few researchers (Wang et al., 2022) provide the third-party ground-truth labels that are manually identified by themselves, we empirically find some ambiguities between their ground-truth labels and the official attack scenario report.
 These ambiguities have an obviously negative effect on DL-IDS, and to some extent, they may even cause the accumulation of errors.
 On the other hand, the development of automated labeling tools is in an awkward position.
-The automation requires prior knowledge of intrusions to design auto-labeling techniques such as instrumentation.
-In other words, the automation can only handle intrusions that it can understand and detect, making it relatively useless to develop such auto-labeling tools.
-In addition, there are no unified evaluation metrics for DL-IDS on these datasets (Cheng et al., 2024).
-For example, precision, recall, F1 score are usually exploited in most of work (Wang et al., 2020; Alsaheel et al., 2021; Rehman et al., 2024; Jia et al., 2024).
-Yet some papers (Ding et al., 2023b) propose to use True Positive Rate (TPR) and False Positive Rate (FPR) as evaluation metrics.
-All the comparison experiments between DL-IDS are thus hard to guarantee that the obtained performance is significant and practical.
-The potential of datasets will be weakened under this non-unified performance evaluation.
+The log data is generated based on its given prior knowledge of intrusions (Cheng et al., 2025), whereas the challenge of DL-IDS is to detect zero-day intrusions.
+This tends the development of such automated tools to be somewhat pointless.
+In addition, there are no unified and effective evaluation metrics for DL-IDS (Cheng et al., 2024), which further weakens the potential of datasets.
+For example, precision, recall, F1 score are usually exploited in most studies (Wang et al., 2020; Alsaheel et al., 2021; Rehman et al., 2024; Jia et al., 2024), while some papers (Ding et al., 2023b) propose to use True Positive Rate (TPR) and False Positive Rate (FPR) as evaluation metrics.
+This makes the comparison experiments usually unfair and hard to tell if the validation is convincing.
+We also note that in many cases where the percentage of negatives (or malicious log entries) is low, sacrificing FPR can always significantly increase TPR.
+For example, sacrificing 1,000 false positives for one true positive might only increase FPR by 0.05%, but would increase TPR by 5%.
 7.1.2. Insufficient Amount of Data
 Although log data is generated very quickly (e.g., eBay generates 1.2 PB log data per day by 2018 (Samuel, 2018)), DL-IDS is still facing challenges in insufficient amounts of data.
-The reasons can be three-fold:
-First, log data has an extremely large number of trivial events, which are proven to be not effective and usually removed by graph summarization (Zengy et al., 2022; Yang et al., 2023b).
+Discounting the above data quality issues such as inaccuracies, the reasons are three-fold:
+First, log data has an extremely large number of trivial events, which are proven ineffective and usually removed by graph summarization (Zengy et al., 2022; Yang et al., 2023).
 For example, data provenance provides fine-grained information about memory-related events, such as data-to-memory mapping and protection of certain memory addresses.
 These memory-related events basically do not involve attacks, and unfortunately, are always orthogonal to the existing DL-IDS.
-However, to ensure the completeness of data provenance and to capture very infrequent but inevitable memory attacks, these memory-related events are still recorded in benchmark datasets.
-Therefore, the usable part of each dataset is rather smalle for DL-IDS, which can be reflected by the high summarization ratio achieved by graph summarization approaches (e.g., 70% (Xu et al., 2016)).
+However, to ensure the completeness requirement of data provenance and to capture very infrequent but inevitable memory attacks, these memory-related events are still recorded in benchmark datasets.
+As a result, the usable part of each dataset is rather small for DL-IDS, which can be reflected by the high summarization ratio achieved by graph summarization approaches (e.g., 70% (Xu et al., 2016)).
 The second reason for an insufficient amount of data is the limited dataset representativeness.
@@
 Hence, training DL-IDS on these non-representative datasets could be a disaster for the computer systems that they protect.
-Table 7. Summarization of computating resources.
-Type Approach Device Quantity
-GPU DeepHunter (Wei et al., 2021a) Tesla P4 2
-SHADEWATCHER (Zengy et al., 2022) Tesla V100 1
-PROGRAPHER (Yang et al., 2023b) GeForce GTX TITAN X 1
-FLASH (Rehman et al., 2024) GeForce RTX 2080 1
-LogBD (Liu et al., 2023) GeForce RTX 2080 1
-HDLNIDS (Qazi et al., 2023) GeForce GTX 2080Ti 1
-NeuralLog (Le and Zhang, 2021) Tesla K40c 1
-LogUAD (Wang et al., 2022b) Tesla V100 1
-TCN-Log2Vec (Jiang et al., 2023) GeForce GTX 1660s 1
-LogTracer (Niu et al., 2022) GeForce GTX 3080Ti 1
-AIRTAG (Ding et al., 2023b) GeForce RTX 6000 1
-MAGIC (Jia et al., 2024) GPU Unknown
-CPU Only UNICORN (Han et al., 2020) vCPU 8
-THREATRACE (Wang et al., 2022a) vCPU 16
-Pagoda (Xie et al., 2018) Physical CPU 2
-PROVDETECTOR (Wang et al., 2020) Physical CPU 1
-SigML++ (Trivedi et al., 2023) Physical CPU 1
-LogAnomaly (Meng et al., 2019) Physical CPU 1
-GAN-EDC (Duan et al., 2021) Physical CPU 1
-LogBASA (Liao et al., 2023) Physical CPU 1
-KAIROS (Cheng et al., 2024) Physical CPU 1
-CAPTAIN (Ramaki et al., 2023) Physical CPU 1
-Paradise (Wu et al., 2022) Physical CPU Unknown
 Finally, the accessibility of datasets further exacerbates the insufficient data problem.
@@
 7.1.3. Potential Heavy Computation Requirements
-Similar to DL techniques, DL-IDS also requires a large amount of computing resources to improve their performance.
+Similar to other DL techniques, DL-IDS also requires a potentially large amount of computing resources to improve their performance.
 According to (Roberts et al., 2024), the generalizability of neural models is proportional to the investment of computing resources.
-To illustrate, we report the computing resources of some well-known DL-IDS in Table 7.
-It shows that existing DL-IDS are all relatively lightweight, as manifested by only one NVIDIA Tesla V100 GPU or only one NVIDIA GeForce RTX 6000 GPU.
-However, the requirements for computing resources are related to the number of log data.
-When the challenge of insufficient data is mitigated and more data is available, more computing resources are inevitably required.
-Meanwhile, we will illustrate in Section 7.2 that there are plenty of powerful techniques that have not been introduced in DL-IDS, which similarly will bring in computation requirements.
-Additionally, several DL-IDS (Han et al., 2020; Cheng et al., 2024; Wu et al., 2022; Trivedi et al., 2023; Duan et al., 2021; Meng et al., 2019) have been specifically engineered for real-time intrusion detection.
-For example, the experiments in Unicorn (Han et al., 2020) demonstrate that the processing time of Unicorn is linear to its workloads with only one core.
-Yet the computational units, either CPU or GPU, all have their performance upper bound.
-Unicorn should be implemented in parallel if its workload is too high.
+Supposing that the challenge of insufficient data is mitigated and a large volume of log data is available, more computing resources are inevitably required.
+Besides, we will illustrate in Section 7.2 that there are plenty of powerful techniques that have not been introduced in DL-IDS, which will also bring in computation requirements.
+Unfortunately, acceleration methods like parallel computation and efficient retrieval have not been fully scheduled by the cybersecurity community.
+An example is that the computation time of Unicorn equipped with one core is proven linear to its workloads (Han et al., 2020).
+It is clear that the efficiency of Unicorn, which is not implemented in parallel, will reach the bottleneck as this core does.
 7.1.4. Future Directions
 To conclude, the challenges for DL-IDS in fundamental resources consist of data quality, data volume, and computational overhead.
-Apart from unintentional errors and non-technical issues in fundamental resources, the research questions that urgently need to be addressed include the contradiction between unaffordable manual labeling and non-generalizable auto-labeling techniques, non-unified benchmark datasets and performance metrics, as well as potential heavy computational overheads.
+Apart from unintentional errors and non-technical issues in fundamental resources, the research questions that urgently need to be addressed include the contradiction between unaffordable manual labeling and non-generalizable auto-labeling techniques, non-unified benchmark datasets and evaluation metrics, as well as potential heavy computational overheads.
 Therefore, we summarize the future directions as follows:
@@
 T5 (Raffel et al., 2020), BERT (Devlin et al., 2019), GPT (Radford, 2018), GPT-4 (Achiam et al., 2023), LaMDA (Thoppilan et al., 2022), and LLaMA (Touvron et al., 2023) are notable examples.
-With the development of pre-training techniques, LLMs have been adopted in many fields such as finance (Zhao et al., 2024), education (Neumann et al., 2024), medicine (Peng et al., 2023), and even other domains of cybersecurity (Cummins et al., 2025; Hu et al., 2024; Gu, 2023).
-Nevertheless, the adoption of LLMs in DL-IDS is stagnant, as shown in Figure 8.
+With the development of pre-training techniques, LLMs have been adopted in many fields such as finance (Zhao et al., 2024b), education (Neumann et al., 2024), medicine (Peng et al., 2023), and even other domains of cybersecurity (Cummins et al., 2025; Hu et al., 2024; Gu, 2023).
+In contrast, the adoption of LLMs in DL-IDS is stagnant, as shown in Figure 6.
 We can observe that LLMs developed at full speed beginning in 2019.
@@
 In the following, we will illustrate the identified issues and then point out the future directions.
-Figure 8. Interactions between DL models and DL-IDS. While DL models proposed before 2019 have already leveraged in DL-IDS, the emerging LLMs since 2020 remains underdeveloped in this domain.
+Figure 6.
+Interactions between DL models and DL-IDS.
+While DL models proposed before 2019 have already leveraged in DL-IDS, the emerging LLMs (or pre-training theories and thechniques) since 2020 remains underdeveloped in this domain.
 7.2.1. Trade-off between Reliability and Generalizability
@@
 LLMs are developed on the basis of statistical language modeling (Rosenfeld, 2000; Jozefowicz et al., 2016), which is not insufficiently studied for log data.
-The statistical modeling of natural language can be traced back to the early 1950s, when Shannon pioneered the technique of predicting the next element of natural language text (Shannon, 1948) and discussed the n-gram model for English (Shannon, 1951).
+The statistical modeling of natural language can be traced back to the early 1950s when Shannon pioneered the technique of predicting the next element of natural language text (Shannon, 1948) and discussed the n-gram model for English (Shannon, 1951).
 After that, as machine learning came into view of the NLP research communities, language modeling flourished, and many models such as TreeBank (Marcus et al., 1993), word2vec (Mikolov et al., 2013a, b) and LSTM (Hochreiter and Schmidhuber, 1997) were proposed.
@@
 Unfortunately, there are almost no research efforts on statistical modeling of log data, resulting in pre-training techniques of DL-IDS remaining underdeveloped.
-By contrast, as illustrated in Table 8, other data types whose statistical modeling studies started earlier have already done well in pre-training techniques.
-For example, Hindle et al.
+By contrast, as illustrated in Table 7, the statistical modeling studies of other types of data have already started.
+Hindle et al.
 (Hindle et al., 2016) demonstrate that the source code is very repetitive and predictable, and, in fact, even more so than natural language.
-Driven by such statistical modeling conclusion, DL-based source code applications (Xu et al., 2023c, 2024; Li et al., 2024c, 2022; Svyatkovskiy et al., 2020; Feng et al., 2020; Gu et al., 2025) such as code generation and code clone detection make great progress, many of which are proven to be greatly strengthened by pre-training techniques.
+Driven by such statistical modeling conclusion, DL-based source code applications (Xu et al., 2023b, 2024a; Li et al., 2024c, 2022; Svyatkovskiy et al., 2020; Feng et al., 2020; Gu et al., 2025) such as code generation and code clone detection flourish, many of which have already becomes common applications in LLMs.
 Similar cases can be found for speech data, whose applications are like text to speech (Ren et al., 2019; Guo et al., 2024; Park et al., 2025) and speech recognition (Baevski et al., 2021).
-Log data is also created by humans, similar to text, speech, and source code.
-It is generated according to developer-defined log templates, with a form of natural language (e.g., application logs) and formal language (e.g., data provenance in CDM format).
-Given the fact that natural language (e.g., text and speech) and formal language (e.g., source code) both exhibit positive features in statistical modeling, log data urgently demands similar achievements to facilitate its pre-training research.
+We argue that log data is also created by humans, similar to text, speech, and source code.
+It is generated according to developer-defined log templates, with a form of both natural language (e.g., application logs) and formal language (e.g., data provenance in CDM format).
+Given the fact that natural language (e.g., text and speech) and formal language (e.g., source code) both exhibit positive performance in pre-training, log data urgently demands statistical modeling achievements to facilitate its pre-training research.
 Although several works (Michael et al., 2020; Inam et al., 2023) have discussed the features of log data, they are orthogonal to the explainable combination of DL and IDS.
@@
 It is very common that the length of one single log entry is the same as that of one paragraph in natural language texts.
-This happens to be the shortcomings of LLMs - not being trustworthy in generated contents and not being able to handle long text.
+These challenges happen to be the shortcomings of LLMs - not being able to handle long text and not being trustworthy in generated contents.
 7.2.3. Future Directions
@@
 We summarize the future directions as follows:
-Future Directions Investigating how and where to introduce LLMs into DL-IDS, with the objective of balancing the generalizability provided by LLMs and the reliability required by DL-IDS.
+Future Directions Investigating how and where to introduce LLMs into DL-IDS like (Ni et al., 2025), with the objective of balancing the generalizability provided by LLMs and the reliability required by DL-IDS.
 Exploring fundamental statistical modeling theories for log data.
@@
 DL-IDS possess abilities that the traditional IDS lack, or are difficult to realize, such as generalizability for zero-day attacks and modeling ability for complicated downstream tasks.
-We will elaborate the possible new-style applications and discuss the challenges in and introduced by them.
+We will elaborate on the possible new-style applications and discuss the challenges in and introduced by them.
 7.3.1. Limited Forward and Backward Tracing Scope
@@
 Generalized intrusion is the malicious attempts against a computer, a network, or the corresponding security facilities, whose attributes encompass not only itself but also its underlying root causes and the relevant control measures.
-In this way, the detection of DL-IDS has been extended to the broadly defined intrusions, including their attributes such as root causes and control measures.
-When executing backward tracing analysis, DL-IDS are not only required to detect the starting symptom nodes, but also required to find the root causes of these nodes (or vulnerabilities in source codes).
-In the forward tracing analysis, except for detecting the symptom nodes affected by intrusions, DL-IDS should perform an in-depth analysis to discover the potential symptom nodes and provide control measures for the intrusions.
+In this way, the detection of DL-IDS has been extended to the broadly defined intrusions, including their attributes of both root causes and control measures.
+When executing backward tracing analysis, DL-IDS are not only required to detect the starting symptom nodes of intrusions, but also required to find the root causes of these symptom nodes (i.e., vulnerabilities in source codes).
+In the forward tracing analysis, except for detecting the symptom nodes affected by intrusions, DL-IDS should perform an in-depth analysis to discover the potentially compromised nodes and provide control measures for handling intrusions.
 Thankfully, several pioneering works have studied similar problems (Chen et al., 2024; Manso et al., 2019).
 In AiVl (Chen et al., 2024), algorithms to bridge log entries and program models are developed using dynamic-static program analysis.
-Root causes of code vulnerabilities are thus allowed to derive from intrusions.
+Root causes for the exploited vulnerabilities are capable of directly deriving from intrusion detection.
 Pedro et al.
-(Manso et al., 2019) investigate detection and mitigation for DDoS attacks, which can be regarded as control measures against specific intrusions.
-We note that these research attempts are all based on heuristics and can only achieve limited efficacy, either in generating root causes using pre-defined rules, or in developing control measures for specific intrusions.
-The main obstacle is the lack of robust modeling and reasoning capabilities, which gives rise to a substantial need for the introduction of advanced techniques like DL.
+(Manso et al., 2019) investigate detection and mitigation methods for DDoS attacks, aiming to control them immediately.
+We note that these research attempts are all based on heuristics, either using pre-defined rules to generate root causes, or developing control measures for specific intrusions.
+Thus, there is a substantial need to introduce advanced DL techniques to this problem.
 7.3.2. Concerns about Data-driven Adversarial Attacks
-To clarify the detection performance, DL-IDS commonly idealize the parts other than detection models in their threat model, of which the most commonly idealized part is data.
+To validate the detection performance, DL-IDS commonly idealize the experimental data in their threat model.
 Such idealization, however, leaves DL-IDS with weaknesses that could be exploited by invaders.
-One usual assumption for data is related to data poisoning attacks, as (Rehman et al., 2024; Hassan et al., 2019; Jia et al., 2024; Han et al., 2020) did.
-In this assumption, no attacks are considered to compromise the security of the log collection systems, that is, log data utilized in DL-IDS is absolutely harmless.
-But apparently, as attacks become more stealthy and complicated, it is impossible to satisfy such an assumption to some extent.
-Besides, the harmful data could emerge through non-attacks.
-Security analysts may fail to label log entries correctly, resulting in an effect similar to the data poison attacks.
-Our manual checking empirically identifies that the labeling in threaTrace (Wang et al., 2022a) is not fully consistent with the official ground truth documentation.
-In many cases, by emulating benign system behaviors, the attackers are able to plant backdoors in the training data for DL-IDS, putting the computer systems at invading risk.
-The robustness of DL-IDS is also challenged by evasion attacks.
-To evade the detection, the malicious behaviors usually mimic the benign ones (a.k.a., mimicry attack), making them hard to be detected.
+For example, a common assumption is that no attacks are considered to compromise the security of the log collection systems (Rehman et al., 2024; Hassan et al., 2019; Jia et al., 2024; Han et al., 2020), namely log data utilized in DL-IDS is absolutely harmless.
+But as attacks become more stealthy and complicated, it is impossible to satisfy such an assumption apparently.
+When DL-IDS encounter intentional data poisoning attacks, prediction backdoors could be easily planted as persistent vulnerabilities.
+The robustness of DL-IDS is also challenged by data-driven evasion attacks.
+To evade the detection, the malicious behaviors usually mimic the benign ones (a.k.a., mimicry attacks), making them hard to be detected.
 By early 2002, David et al.
-(Wagner and Soto, 2002) had indicated the danger of mimicry attacks on HIDS.
+(Wagner and Soto, 2002) have indicated the danger of mimicry attacks on HIDS.
 Recently, researchers have started to investigate mimicry attacks on DL-IDS (Mukherjee et al., 2023; Goyal et al., 2023) and their studies all present effective evasion of detection.
-While recognized to be powerful, DL-IDS are plagued by the issue that even a trivial perturbation in the data can result in a successful evasion attack (Chakraborty et al., 2018).
+From a study (Chakraborty et al., 2018), DL-IDS can be even plagued by a trivial perturbation in log data.
 Aware of this issue, R-caid (Goyal et al., 2024) proposes to embed root causes into the detection model for countering adversarial attacks.
@@
 7.3.3. Underexplored Promising Scenarios
-DL-IDS show excellent performance in the protection of computer and network systems recently.
-Unfortunately, there are still many promising scenarios for DL-IDS that have not been explored sufficiently yet.
+While DL-IDS show excellent performance in the protection of computer and network systems recently, there are still many promising scenarios for DL-IDS that have not been explored sufficiently.
 Mobile edge computing (MEC) (Mao et al., 2017; Abbas et al., 2017) is a typical scenario.

### v3

@@
 Intrusion detection systems, deep learning, survey
-copyright: nonejournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 1journalnumber: 1article: 1publicationmonth: 4ccs: Security and privacy Intrusion detection systemsccs: Computing methodologies Machine learningccs: General and reference Surveys and overviews
+copyright: nonejournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 1journalnumber: 1article: 1publicationmonth: 10ccs: Security and privacy Intrusion detection systemsccs: Computing methodologies Machine learningccs: General and reference Surveys and overviews
 1. Introduction
@@
 Then, we employed Connected Papers (Papers, 2020), a visual tool that assists researchers in finding relevant academic papers, to ensure that we did not overlook the typical related literature.
-Since the found literature is numerous and rather generalized for the DL-IDS scope, we carefully checked their topics and prioritized only academic papers that are related to DL.
-Finally, all these papers were filtered based on the impact factors of their published journals or academic conferences, leaving us a total of 113 papers.
-We identified a few venues that have published many significant papers in the field of DL-IDS, such as Security, S&P, CCS, NDSS, TIFS, TDSC, ICSE, ASE, ESEC/FSE, TSE, EuroSys, NSDI, OSDI, SOSP, ATC, ICDE, TKDE, SIGMOD, KDD, WWW, and IJCAI.
-We broadly divide them into four categories: security, software, system, and data.
+Since the found literature is numerous and rather generalized for the DL-IDS scope, we carefully checked their topics and prioritized only academic papers that are highly related.
+Finally, all these papers were filtered based on the impact factors of their published journals or academic conferences, leaving us a total of 131 papers.
+We identified a few venues that have published many significant papers in the field of DL-IDS, such as Usenix Security, S&P, CCS, NDSS, TIFS, TDSC, ICSE, ASE, ESEC/FSE, TSE, OSDI, NSDI, EuroSys, SOSP, ATC, ICML, KDD, WWW, TKDE, ICDE, and SCIS.
+We broadly divide them into five categories: security, software, system, data, and interdiscipline.
 The distribution of these papers with their published years is reported in Figure 3.
@@
 Raw data indicates records that document events, activities, and operations that occur within a system, application, or network (a.k.a., logs), represented by audit logs or application logs within hosts, or network traffic between hosts.
-By collecting useful logs, DL-IDS is allowed to monitor the health condition and operational status of information systems (Zhang et al., 2023b).
+By collecting useful logs, DL-IDS is allowed to monitor the health condition and operational status of information systems (Zhang et al., 2023b; Ma et al., 2018).
 Common attributes of logs include timestamp, event type, subject, object, description, etc.
@@
 Tailored Compression Algorithms
-Different from natural language data, log data usually has specific structures and formal expressions that help further compression.
-Yao et al.
-(Yao et al., 2021) propose LogBlock, which obtains small log blocks before compression and then uses a generic compressor to compress logs.
-Liu et al.
-(Liu et al., 2019) propose Logzip, which employs clustering algorithms to iteratively extract templates from raw logs and then obtain coherent intermediate representations for compressing logs.
-Rodrigues et al.
-(Rodrigues et al., 2021) propose the lossless compression tool CLP, aiming to quickly retrieve log data while meeting compression requirements.
-CLP proposes to combine domain-specific compression and search with a generic lightweight compression algorithm.
-Li et al.
-(Li et al., 2024d) conduct empirical research on log data and propose LogShrink to overcome their observed limitations by leveraging the commonality and variability of log data.
-LogBlock (Yao et al., 2021) is designed to help existing jobs perform better.
-It reduces duplicate logs by preprocessing log headers and rearranging log contents, thereby improving the compression ratio of log files.
-LogReducer (Yu et al., 2023) is a framework that combines log hotspot identification and online dynamic log filtering.
-Its non-intrusive design significantly reduces log storage and runtime overhead.
-$\mu$Slope (Wang et al., 2024) is a compression and search method for semi-structured log data.
-It achieves efficient storage and query performance through data segmentation, pattern extraction, and index-free design.
-Denum (Yu et al., 2024) significantly improves log compression rates by optimizing the compression of digital tokens in logs.
-It is an efficient log compression tool suitable for scenarios where you need to save storage space or transmission bandwidth.
+[…]
 4.3. Log Parsing
@@
 LogApprox (Michael et al., 2020) 2020 GC + CPR + DPR None
+TeRed (Li et al., 2025) 2025 LogGC + CPR + PCAR + F-DPR + NodeMerge None
 Online ProTracer (Ma et al., 2016) 2016 BEEP + No Summarization Instrumentation
@@
 LogApprox (Michael et al., 2020) aims to summarize the most space-intensive events found in logs, namely file I/O activity, which can account for up to 90% of the log content.
+TeRed (Li et al., 2025) employs unit tests to learn the system’s normal behavior patterns for reducing provenance graphs, allowing it not to impact attack detection and investigation.
 5.1.2. Online Graph Summarization
@@
 Based on the types of log data, we categorize the attack detection approaches into audit log-based, application log-based, network traffic-based, and hybrid log-based detectors.
+The overview and taxonomy of attack detection approaches are presented in Table 4.
+We note that recent years have also published many other academic papers for attack detection (Xie et al., 2018; Wu et al., 2022; Li et al., 2023b; Milajerdi et al., 2019a; Yu et al., 2021; Chen et al., 2024; El Khairi et al., 2024; Wang et al., 2023; Hassan et al., 2020a).
+Yet these papers are slightly related to DL-IDS, which are thus excluded in our survey for conciseness.
 Table 4. Overview and taxonomy of attack detection approaches.
@@
 DistDet (Dong et al., 2023) 2023 HST Online Subgraph
+Velox (Bilot et al., 2025) 2025 FCN Online Node
 Audit ShadeWatcher (Zengy et al., 2022) 2022 TransR Offline Node
-Log Graph threaTrace (Wang et al., 2022) 2022 GraphSAGE Online Node
-Neural ProGrapher (Yang et al., 2023) 2023 graph2vec Online Subgraph
-Network MAGIC (Jia et al., 2024) 2024 GAT Online Node, Subgraph
-Flash (Rehman et al., 2024) 2024 GraphSAGE Online Node
+Log threaTrace (Wang et al., 2022) 2022 GraphSAGE Online Node
+Graph ProGrapher (Yang et al., 2023) 2023 graph2vec Online Subgraph
+Neural MAGIC (Jia et al., 2024) 2024 GAT Online Node, Subgraph
+Network Flash (Rehman et al., 2024) 2024 GraphSAGE Online Node
 R-caid (Goyal et al., 2024) 2024 GNN Offline Node
 Argus (Xu et al., 2024b) 2024 MPNN, GRU - Node
+TAPAS (Zhang et al., 2025) 2025 LSTM-GRU Online Task
 Traditional Wei et al. (Xu et al., 2009) 2009 PCA, TF-IDF - Log Entry
 Learning Bodik et al. (Bodik et al., 2010) 2010 Logistic Regression Online Log Entry
+AMOD (Dong et al., 2018) 2018 SVM HYBRID Online Log Entry
 DeepLog (Du et al., 2017) 2017 LSTM Online Log Entry
 LogRobust (Zhang et al., 2019) 2019 Attention LSTM - Log Entry
-Application Sequence LogAnomaly (Meng et al., 2019) 2019 template2vec, LSTM Online Log Entry
-Log Neural LogC (Yin et al., 2020) 2020 LSTM Online Log Entry
-Network NeuralLog (Le and Zhang, 2021) 2021 BERT - Log Entry
-PLELog (Yang et al., 2021) 2021 Attention GRU Online Log Entry
+Application LogAnomaly (Meng et al., 2019) 2019 template2vec, LSTM Online Log Entry
+Log Sequence LogC (Yin et al., 2020) 2020 LSTM Online Log Entry
+Neural NeuralLog (Le and Zhang, 2021) 2021 BERT - Log Entry
+Network PLELog (Yang et al., 2021) 2021 Attention GRU Online Log Entry
 SpikeLog (Qi et al., 2023) 2023 DSNN - Log Entry
 LogCraft (Zhang et al., 2024) 2024 Meta Learning - Log Entry
-Traffic Whisper (Fu et al., 2021) 2021 K-Means - Host
-Traditional SigML++ (Trivedi et al., 2023) 2023 ANN - Encrypted Log
-Learning OADSD (Zhang et al., 2023a) 2023 Isolation Forest Online Packet
+Tweezers(Cui et al., 2024) 2024 GATv2, BERTweet Online Log Entry
+LogSer (Chai et al., 2024) 2024 BERT Online Log Entry
+LogDLR(Zhou et al., 2025) 2025 Transformer, SBERT Online Log Entry
+Traffic NetPro (Li et al., 2017) 2017 Merkle Hash Tree Online Route
+Log CATH (Guo et al., 2019) 2019 Cusp Model Online Flow
+Traditional Whisper (Fu et al., 2021) 2021 K-Means - Host
+Learning SigML++ (Trivedi et al., 2023) 2023 ANN - Encrypted Log
+OADSD (Zhang et al., 2023a) 2023 Isolation Forest Online Packet
 LtRFT (Tang et al., 2023) 2023 LambdaMART Offline Packet
+AGC (Wu et al., 2025c) 2025 Clustering - Packet
 Kitsune (Mirsky et al., 2018) 2018 AutoEncoder Online Packet
@@
 I2RNN (Song et al., 2022) 2022 I2RNN - Packet
-Graph and ERNN (Zhao et al., 2022b) 2022 ERNN - Flow
-Sequence Euler (King and Huang, 2023) 2023 GNN, RNN - Flow
-Neural pVoxel (Fu et al., 2023) 2023 - - Packet, Flow
-Network NetVigil (Hsieh et al., 2024) 2024 E-GraphSage - Flow
-Exosphere (Fu et al., 2024) 2024 CNN - Packet
-DFNet (Zhao et al., 2024a) 2024 DFNet - Packet
+ERNN (Zhao et al., 2022b) 2022 ERNN - Flow
+Euler (King and Huang, 2023) 2023 GNN, RNN - Flow
+Graph and pVoxel (Fu et al., 2023) 2023 - - Packet, Flow
+Sequence NetVigil (Hsieh et al., 2024) 2024 E-GraphSage - Flow
+Neural Exosphere (Fu et al., 2024) 2024 CNN - Packet
+Network DFNet (Zhao et al., 2024b) 2024 DFNet - Packet
 RFH-HELAD (Zhong et al., 2024) 2024 RPGAN, Deep kNN - Packet
+ReTrial (Zhao et al., 2024a) 2024 Bayesian Inference Online Flow
+HEN (Wei et al., 2024) 2024 AE-LSTM - Packet, Flow
+TCG-IDS (Wu et al., 2025b) 2025 TGN Online Flow
+A-NIDS(Zha et al., 2025) 2025 Stacked CTGAN Online Flow
+GTAE-IDS(Ghadermazi et al., 2025) 2025 Graph Transformer Online Packet, Flow
 Hybrid Hybrid OWAD (Han et al., 2023) 2024 Autoencoder Online Hybrid
 FG-CIBGC (Ni et al., 2025) 2025 DisenGCN, ICL - Hybrid
-The overview and taxonomy of attack detection approaches are presented in Table 4.
-We note that recent years have also published many other academic papers for attack detection (Xie et al., 2018; Wu et al., 2022; Li et al., 2023b; Milajerdi et al., 2019a; Yu et al., 2021; Chen et al., 2024; El Khairi et al., 2024; Wang et al., 2023; Hassan et al., 2020a).
-Yet these papers are slightly related to DL-IDS, which are thus excluded in our survey for conciseness.
 5.2.1. Audit Log-based Detectors
@@
 DistDet (Dong et al., 2023) is a distributed detection system that builds host models in the client side, filters false alarms based on their semantics, and derives global models to complement the host models.
+Velox (Bilot et al., 2025) derives from Orthrus and replaces the complex TGN-based encoder with a simple fully-connected network (FCN), leading to a lightweight and efficient neural network.
 Graph Neural Network
@@
 It thus devises a form of discrete temporal graph and uses encoder-decoder unsupervised learning to detect different types of attacks.
+TAPAS (Zhang et al., 2025) leverages a stacked LSTM-GRU model and a task-guided segmentation algorithm to reduce the spatiotemporal dimensions of APT detection, achieving efficient, low-cost, and accurate detection.
+In addition to the aforementioned detectors, recent researchers have developed numerous useful tools for better understanding audit logs, such as data visualization analysis tool (Liu et al., 2025a) and counterfactual-driven attack explanation generator (Wu et al., 2025a).
 5.2.2. Application Log-based Detectors
@@
 (Bodik et al., 2010) leverage a logistic regression model on a new and efficient representation of a datacenter’s state called fingerprint to detect previously seen performance crises in that datacenter.
+AMOD (Dong et al., 2018) uses the SVM HYBRID strategy to filter query annotations from web request logs and then update the stacked generalization detection model to efficiently detect web code injection attacks and obtain malicious queries to update the web application firewall (WAF) library.
 Sequence Neural Network
@@
 LogCraft (Zhang et al., 2024) is an end-to-end unsupervised log anomaly detection framework based on automated machine learning, which mitigates the cost of understanding datasets and makes multiple attempts for building algorithms.
+Tweezers (Cui et al., 2024) uses a large language model to identify entities and build a relationship graph, and generates embeddings through graph attention network optimization to achieve security incident detection.
+LogSer (Chai et al., 2024) parses logs by preprocessing parameters, splitting logs, tree parsing, and template merging.
+It then inputs relevant embeddings into BERT training to detect anomalies, generate reports, and perform incremental updates.
+LogDLR (Zhou et al., 2025) uses SBERT embeddings and a Transformer autoencoder with domain adversarial training to learn domain-invariant features, detecting anomalies via reconstruction error.
 5.2.3. Network Traffic-based Detectors
@@
 Given the fact that network traffic is usually encrypted for secure communications, feature engineering-guided machine learning is widely applied in NIDS.
+NetPro (Li et al., 2017) employs traceability reasoning with Merkle Hash Trees and digital signatures to detect direct and indirect MANET routing attacks while preserving node privacy, and outputs a traceability graph to identify malicious nodes and behaviors.
+CATH (Guo et al., 2019) is a catastrophe-theory-based approach for DoS detection in software-defined networks (SDNs), which leverages the selection, normalization, and fusion of statistical flow attributes to model network states.
 Whisper (Fu et al., 2021) pays attention to both high accuracy and high throughput by utilizing frequency domain features.
@@
 LtRFT (Tang et al., 2023) innovatively introduces Learning-To-Rank scheme for mitigating the low-rate DDoS attacks targeted at flow tables.
+AGC (Wu et al., 2025c) maps the original data into the embedding space through embedding learning to obtain more representative anchor points, thus achieving fine-grained classification of low-quality label data.
 Graph and Sequence Neural Network
@@
 Exosphere (Fu et al., 2024) detects flooding attacks by analyzing packet length patterns, without investigating any information in encrypted packets.
-DFNet (Zhao et al., 2024a) is a DDoS prevention paradigm denoted by preference-driven and in-network enforced shaping.
+DFNet (Zhao et al., 2024b) is a DDoS prevention paradigm denoted by preference-driven and in-network enforced shaping.
 RFH-HELAD (Zhong et al., 2024) consists of a $K$ classification model based on a deep neural network and a $K+1$ classification combining GAN and Deep kNN for detecting anomalies in network traffic.
+ReTrial (Zhao et al., 2024a) employs an improved graph attention network with Bayesian and EM algorithms to iteratively correct misleading links, enabling robust detection of encrypted malicious traffic.
+HEN (Wei et al., 2024) uses SMOTE to enhance data, trains LightGBM, generates explanations via SHAP, trains AE-LSTM to reconstruct SHAP values, sets a threshold from training errors, and marks test traffic with excess errors as attacks for intrusion detection.
+TCG-IDS (Wu et al., 2025b) is the first self-supervised temporal contrastive GNN for network intrusion detection, capturing spatiotemporal traffic dependencies with high accuracy and low false alarms.
+A-NIDS (Zha et al., 2025) uses a shallow fully connected network for real-time detection and a Stacked CTGAN generator to address catastrophic forgetting and old data storage costs.
+GTAE-IDS (Ghadermazi et al., 2025) uses a graph autoencoder with a Transformer encoder and DNN decoder to learn benign traffic, enabling label-free, near-real-time intrusion detection and new attack identification.
 5.2.4. Hybrid Log-based Detectors
@@
 Benefiting from DL techniques, both forward and backward analysis can be achieved by learning patterns of attack scenario graphs.
+Furthermore, visual analytics techniques have been widely used to assist security analysts in understanding the causal chain of intrusions (Zhang et al., 2017; Zhao et al., 2013).
 Table 5 summarizes the overview of attack investigation approaches.
@@
 Traditional ProvDetector (Wang et al., 2020) 2020 doc2vec Path
+Learning BehaviorBaseline (Zhu et al., 2025) 2025 FastText Path
 Sequence ATLAS (Alsaheel et al., 2021) 2021 LSTM Graph
@@
 AirTag (Ding et al., 2023b) 2023 BERT Graph
-Graph Liu et al. (Liu et al., 2022a) 2022 struc2vec Graph
-Neural Karios (Cheng et al., 2024) 2023 GNN Graph
-Network TREC (Lv et al., 2024) 2024 GNN Graph
-Orthrus (Jiang et al., 2025) 2025 UniMP Path
+Liu et al. (Liu et al., 2022a) 2022 struc2vec Graph
+Graph Karios (Cheng et al., 2024) 2023 GNN Graph
+Neural TREC (Lv et al., 2024) 2024 GNN Graph
+Network Orthrus (Jiang et al., 2025) 2025 UniMP Path
+Slot (Qiao et al., 2025) 2025 GNN Graph
+FeCoGraph (Mao et al., 2025) 2025 GCN Graph
 Traditional Learning
@@
 Then a density-based detection is deployed to detect abnormal causal paths in the provenance graph.
+BehaviorBaseline (Zhu et al., 2025) presents a novel learning-based anomaly detection method for large-scale provenance graphs.
+It incorporates dynamic graph processing with adaptive encoding and a tag-propagation framework for real-time detction.
 Sequence Neural Network
@@
 It first detects malicious hosts using a GNN encoder and then reconstructs the attack path through dependency analysis.
+Slot (Qiao et al., 2025), based on provenance graphs and graph reinforcement learning, uncovers hidden relationships among system behaviors, dynamically adapts to new activities and attack strategies, resists adversarial attacks, and automatically constructs attack chains.
+FeCoGraph (Mao et al., 2025) directly processes traffic embedding through line graphs to adapt to various GNNs, covering more attack scenarios while protecting data privacy.
 6. Benchmark Datasets
@@
 Due to copyright concerns, some replicates (Jia et al., 2024; Ding et al., 2023b) merely provide post-processed log data without their processing source codes.
+  • Operating Systems: The operating system determines the generalizability of the dataset.
+The more operating systems a dataset covers and the more common they are, the more comprehensively it can evaluate PIDS performance.
 6.2. Public Datasets
@@
 Loghub (Zhu et al., 2023a) 2020 77 GB - - .txt W, L, M, S
+ATLAS (Alsaheel et al., 2021) 2021 0.5 GB 10 Yes .txt W
+ATLASv2 (Riddle et al., 2023) 2023 12 10 Yes .txt W
 ProvSec (Shrestha et al., 2023) 2023 - 11 Yes .json L
+AutoLabel (Peng et al., 2025) 2025 136 GB 29 Yes .json L
 LANL Dataset (Kent, 2015) is collected within the internal computer network of Los Alamos National Laboratory’s corporate.
@@
 For the usage scenarios of Loghub, about 35% are anomaly detection, 13% are log analysis, and 8% are security.
+ATLAS dataset (Alsaheel et al., 2021) implements 10 attacks based on their detailed reports on real-world APT campaigns and generates audit logs in a controlled testbed environment.
+Among the ten attacks, four are from single host and the rest six are from multiple hosts.
+All attacks were developed and executed on Windows 7 32-bit virtual machines and took an hour to complete, along with a 24-hour-window audit logs for benign system behaviors.
+ATLASv2 dataset (Riddle et al., 2023) enriches the ATLAS dataset with higher quality background noise and additional logging vantage points.
+In this dataset, two researchers use the victim machines as their primary work stations throughout the course of engagement, instead of depending on automated scripts to generate activity.
+System logging, in contrast, cover a five-day period, where the first four days simulate normal work days and the fifth day begins with benign activity then trasitions into execution of the corresponding attack.
 ProvSec dataset (Shrestha et al., 2023) is created for system provenance forensic analysis.
@@
 In ProvSec, 11 realistic attack scenarios with real software vulnerabilities and exploits are used and an algorithm to improve the data quality in the system provenance forensics analysis is presented.
+AutoLabel dataset (Peng et al., 2025) automates fine-grained log labeling by reducing the labeling problem to obtaining an accurate attack subgraph in a provenance graph.
+Its experiments consist of 29 scenarios, including 25 real CVE vulnerabilities across 12 widely-used applications (spanning 5 programming languages) plus a Sandworm threat simulation by MITRE CTID.
 7. Challenges and Future Directions
@@
 This limited number of attack scenarios suggests that existing datasets are almost impossible to represent the diversified attack methods, as the number of CVE records has already been over 280,000 (Corporation, 2025).
-Furthermore, the existing datasets such as DAPRA TC (DARPA, 2022) are collected in a specific experimental environment and may not cover other types of normal system behaviors.
+Furthermore, the existing datasets such as DAPRA TC E3 (DARPA, 2022) are collected in a specific experimental environment and may not cover other types of normal system behaviors, and are proven that a significant amount of synthetic data exists (Liu et al., 2025a).
+DARPA TC E5 (DARPA, 2022) is unusable for most experiments due to the sparse and error-filled documentation.
 Unicorn SC (Han et al., 2020) is generated by an idealized simulation of supply chain scenarios, which means many real-world features are prone to be ignored in this dataset.
@@
 Due to privacy and copyright issues, some datasets may be proprietary or difficult to obtain (Wang et al., 2020, 2023).
-For example, ProvDetector (Wang et al., 2020) conducted a three-month system evaluation in an enterprise environment with 306 hosts and collected benign provenance data of 23 target programs.
+Moreover, ProvDetector (Wang et al., 2020) conducted a three-month system evaluation in an enterprise environment with 306 hosts and collected benign provenance data of 23 target programs.
 Yet this dataset has not been made public, rendering it unavailable to improve other DL-IDS and almost all the assessment settings related to ProvDetector are susceptible to inequity.
@@
 T5 (Raffel et al., 2020), BERT (Devlin et al., 2019), GPT (Radford, 2018), GPT-4 (Achiam et al., 2023), LaMDA (Thoppilan et al., 2022), and LLaMA (Touvron et al., 2023) are notable examples.
-With the development of pre-training techniques, LLMs have been adopted in many fields such as finance (Zhao et al., 2024b), education (Neumann et al., 2024), medicine (Peng et al., 2023), and even other domains of cybersecurity (Cummins et al., 2025; Hu et al., 2024; Gu, 2023).
+With the development of pre-training techniques, LLMs have been adopted in many fields such as finance (Zhao et al., 2024c), education (Neumann et al., 2024), medicine (Peng et al., 2023), and even other domains of cybersecurity (Cummins et al., 2025; Hu et al., 2024; Gu, 2023).
 In contrast, the adoption of LLMs in DL-IDS is stagnant, as shown in Figure 6.
@@
 Nowadays, some researchers (Ali, 2024; G.
-Lira et al., 2024; Li et al., 2024b) have started to explore the applications of LLMs on DL-IDS.
+Lira et al., 2024; Li et al., 2024b; Mukherjee and Kantarcioglu, 2025) have started to explore the applications of LLMs on DL-IDS.
 Yet the theories and techniques of such combination remain challenges.
@@
 The governing concern for the employment of LLMs in DL-IDS is reliability (or explainability).
-Although offering generalizability, LLMs have long been denounced to have issues with hallucinations (Martino et al., 2023; Yao et al., 2023), privacy (Yao et al., 2024) and overreliance (Kim et al., 2024).
+Although offering generalizability, LLMs have long been denounced to have issues with hallucinations (Martino et al., 2023; Yao et al., 2023), privacy (Yao et al., 2024; Yang et al., 2025; He et al., 2025), overreliance (Kim et al., 2024), and backdoor threats (Liu et al., 2025b).
 These unexplainable and uncontrollable features are an absolute disaster for DL-IDS.
@@
 (Hindle et al., 2016) demonstrate that the source code is very repetitive and predictable, and, in fact, even more so than natural language.
-Driven by such statistical modeling conclusion, DL-based source code applications (Xu et al., 2023b, 2024a; Li et al., 2024c, 2022; Svyatkovskiy et al., 2020; Feng et al., 2020; Gu et al., 2025) such as code generation and code clone detection flourish, many of which have already becomes common applications in LLMs.
+Driven by such statistical modeling conclusion, DL-based source code applications (Xu et al., 2023b, 2024a; Li et al., 2024c, 2022a; Svyatkovskiy et al., 2020; Feng et al., 2020; Gu et al., 2025) such as code generation and code clone detection flourish, many of which have already becomes common applications in LLMs.
 Similar cases can be found for speech data, whose applications are like text to speech (Ren et al., 2019; Guo et al., 2024; Park et al., 2025) and speech recognition (Baevski et al., 2021).
@@
 (Manso et al., 2019) investigate detection and mitigation methods for DDoS attacks, aiming to control them immediately.
+Additionally, semi-automated adaptive network defense (SAND) (Chen et al., 2022) leverages SDN to dynamically generate and deploy defense rules.
 We note that these research attempts are all based on heuristics, either using pre-defined rules to generate root causes, or developing control measures for specific intrusions.
@@
 (Wagner and Soto, 2002) have indicated the danger of mimicry attacks on HIDS.
-Recently, researchers have started to investigate mimicry attacks on DL-IDS (Mukherjee et al., 2023; Goyal et al., 2023) and their studies all present effective evasion of detection.
+Recently, researchers have started to investigate mimicry attacks on DL-IDS (Mukherjee et al., 2023; Goyal et al., 2023; Liu et al., 2022c) and their studies all present effective evasion of detection.
 From a study (Chakraborty et al., 2018), DL-IDS can be even plagued by a trivial perturbation in log data.
@@
 While DL-IDS show excellent performance in the protection of computer and network systems recently, there are still many promising scenarios for DL-IDS that have not been explored sufficiently.
-Mobile edge computing (MEC) (Mao et al., 2017; Abbas et al., 2017) is a typical scenario.
+Mobile edge computing (MEC) (Mao et al., 2017; Abbas et al., 2017; Li et al., 2022b) is a typical scenario.
 In the MEC environment, mobile computing, network control, and storage are pushed at the network edges so as to enable computation-intensive tasks at the resource-limited devices.